### v3

# skipabove=1em,skipbelow=1em

Abstract.
Figure 1. Overall framework of ingestion health sensing and intervention
Ingestive behavior plays a critical role in health, yet many existing interventions remain limited to static guidance or manual self-tracking. With the increasing integration of sensors and perceptual computing, recent systems have begun to support closed-loop interventions that dynamically sense user behavior and provide feedback during or around ingestion episodes. In this survey, we review 136 studies that leverage sensor-enabled or interaction-mediated approaches to influence eating behavior. We propose a behavioral closed-loop paradigm comprising three core components: target behaviors, sensing modalities, and feedback strategies. A taxonomy of sensing and intervention modalities is presented, organized along human- and environment-based dimensions. Our analysis also examines evaluation methods and design trends across different modality-behavior pairings. This review reveals prevailing patterns and critical gaps, offering design insights for future adaptive and context-aware ingestion health interventions.
ingestion health, behavior intervention, behavioral sensing, perceptual computing, interactive systems, closed-loop intervention
copyright: acmlicensedjournal: IMWUTjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: General and reference Surveys and overviews
1. Introduction
Ingestion health has been shown to significantly influence quality of life, longevity, and overall well-being (Jovanov et al., 2014; Mesas et al., 2012). Given its critical importance, it is essential to develop effective strategies that integrate both sensing and intervention for ingestive behaviors. Traditional approaches to ingestion monitoring and intervention, such as food diaries (Thompson and Subar, 2017), self-report questionnaires (Shim et al., 2014), and periodic counseling sessions (Greaves et al., 2011), are heavily dependent on user self-discipline and engagement. Moreover, these methods are typically static, offering one-time or episodic interventions without continuous feedback mechanisms (Raber et al., 2021; Takemoto et al., 2018). As a result, they often fail to dynamically support individuals in developing sustainable and healthy ingestive behaviors over time.
With the advancement of sensing technologies, new opportunities have emerged for low-burden and continuous sensing of ingestive behaviors. These technologies not only support the development of systems capable of capturing rich behavioral data with minimal user effort, but also enable more personalized and adaptive feedback and intervention strategies (Bell et al., 2020). Consequently, there is a growing body of research that leverages sensing and interactive technologies to bridge the gap between data collection and real-time intervention in ingestion health. This shift represents a critical evolution from static, user-driven methods toward context-aware, behavior-informed support systems.
Building on these advances, research in behavioral sensing and intervention has expanded considerably in recent years, alongside growing interest in ingestion health. Several recent reviews have addressed related domains, which can be broadly categorized into three research streams: (1) dietary intake and nutrition assessment, (2) wearable ingestion detection technologies, and (3) behavior change interventions.
Dietary Intake Nutrition Assessment
Traditional dietary assessment methods face challenges in both accuracy and user burden, prompting researchers to explore sensor-based technologies as a more objective alternative. Skinner et al. (2020) proposed a framework that integrates physical, chemical, and biosensors to objectively capture what, how, and when individuals eat. Wang et al. (2022) reviewed real-time sensor systems for nutritional support, while emphasizing deployment challenges. While these advances improve the reliability of intake measurement objectivity, they remain largely confined to assessment purposes and lack extensions toward adaptive behavioral interventions in ingestion health.
Wearable Ingestion Detection Technologies
Wearable-based ingestion detection has been extensively reviewed from different perspectives. Bell et al. (2020) surveyed automatic intake detection studies in free-living settings, highlighting sensor modalities and validation metrics. He et al. (2020) categorized detection systems by sensor type, while Hiraguchi et al. (2023) provided a comprehensive survey covering both eating and drinking behaviors. Beyond general ingestion detection, specialized methods for chewing detection have been explored. Wei et al. (2022) and Mortazavi and Gutierrez-Osuna (2023) reviewed wearable-based chewing sensing techniques and signal processing approaches. Despite these efforts, existing studies and reviews consistently emphasize detection performance, without advancing toward behavioral feedback or adaptive intervention frameworks.
Behavior Change Intervention
Previous studies have reviewed a variety of digital intervention strategies aimed at promoting healthy eating behaviors. Zhang et al. (2020) proposed a generative design framework for technology-mediated eating interventions, while Pan et al. (2021) surveyed digital tools, including mobile apps and wearable devices, designed to support healthy eating in occupational settings. However, these reviews primarily focus on organizing design considerations and intervention strategies, often relying on traditional education or standalone gamification without incorporating sensing technologies or perceptual computing. As a result, they typically lack integration with behavioral sensing or interactive feedback mechanisms.
In summary, while many existing reviews emphasize isolated stages such as intake detection, dietary assessment, or single-instance feedback, few focus on the complete behavioral closed-loop process from sensing to intervention.
This fragmentation highlights the urgent need for a systematic review that frames ingestion health interventions within a behavioral closed-loop paradigm, grounded in perceptual computing and spanning the full pipeline from sensing to feedback and intervention.
To address these gaps, we present a structured review guided by the following research questions (RQs), aimed at evaluating how ingestion health interventions are conceptualized and implemented within behavioral closed-loop systems:
  • RQ1: What behaviors are targeted in the existing literature on ingestion health intervention?
  • RQ2: What are the structural characteristics and design space of perceptual computing-based paradigms for sensing and intervening in ingestive behavior?
  • RQ3: How effective are current intervention paradigms, and what outcomes have they achieved?
  • RQ4: What are the gaps in the implementation of behavioral closed-loop systems in current ingestion health interventions, and what are the recommendations for them?
Contribution Statement:
  • We conduct a systematic review of ingestion health interventions that employ behavioral sensing and interactive systems, and develop a taxonomy of key components across different stages of the behavioral closed-loop paradigm.
  • We propose a behavioral closed-loop framework tailored to ingestive behavior interventions, and review the evaluation of effectiveness and user acceptance in current literature.
  • We identify critical gaps between current ingestion health interventions and real-world behavioral complexities, and provide design insights for better bridging behavioral targets, sensing, reasoning, and intervention stages toward effective closed-loop management.
2. Method
Drawing inspiration from prior review methodologies (Zhang et al., 2020; Jansen et al., 2022; Motahar and Wiese, 2022), we conducted a Systematic Literature Review (SLR) to address our research questions. Specifically, our review process followed the guidelines of the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) proposed by Page et al. (2021).
2.1. Paper Retrieval
Given that most relevant publications place greater emphasis on the feedback or intervention aspect, we also include studies that do not cover the full pipeline from sensing to intervention. After careful discussion, we decided to include all publications in which sensor-based interactive devices or perceptual computing methods are involved at any stage within the behavioral closed-loop process, ranging from sensing to feedback or intervention.
To ensure consistency in inclusion and classification throughout in this paper, we first examined three core terms and finalized their definitions as follows:
  • Ingestive Behavior: The process by which an individual ingests nutrition from food or liquid intake, excluding the intake of medications or medical nutrition supplements, and comprising appetitive behaviors, consummatory behaviors, and post-ingestive physiological responses;
  • Health Intervention: A system-generated feedback mechanism, whether active or passive, that aims to alter the user’s behavior with the goal of improving health from a medical perspective;
  • Perceptual Computing: Technologies that enable machines to perceive, compute, and reason with data from real-world sensory inputs.
Specifically, Benoit and Tracy (2008) defined ingestive behavior as ”a complex process controlled by multiple physiological and psychological mechanisms”, and divided it to two general categories, including ”behaviors involved in seeking, locating, and acquiring foods” and ”behaviors associated with the act of ingestion itself, including chewing and swallowing”. We refer to these two categories as ”appetitive behaviors” and ”consummatory behaviors”, and additionally added ”post-ingestive behaviors” to cover physiological feedback mechanisms and related behaviors after ingestive behavior. Moreover, Mendel and Wu (2010) described perceptual computing as a sensor-based method “associated with machinery for computing and reasoning with perceptions and data,” while Poole (2013) noted that “health interventions leverage technology to deliver timely support, monitoring, and guidance to influence health behaviors and outcomes.” These definitions align closely with the conceptual scope of our study. Accordingly, we adopted and refined the above terms to form the foundation of our inclusion criteria and analytical framework.
Following the PRISMA principle and the search strategy adopted in related work (Hiraguchi et al., 2023), we began our process by identifying keywords terms closely related to core definitions. Moreover, to refine the keyword strategy for literature retrieval, we compiled a seed set of approximately 50 papers, including some reviewed studies that clearly met our inclusion criteria, and top-ranked search results from digital libraries (Röddiger et al., 2022). Subsequently, we applied a TF-IDF (Salton et al., 1975) and N-gram (Cavnar et al., 1994) analysis on the titles, abstracts, and keywords of this seed set to identify additional high-frequency or semantically relevant terms. These were used to augment our original keyword list, ensuring broader coverage of technically and behaviorally relevant studies. To illustrate the outcome of this refinement process, we present the top five high-frequency terms extracted via TF-IDF and Bi-gram analysis (excluding semantically uninformative combinations such as function words or prepositional phrases) as follows:
{mdframed}
[backgroundcolor=gray!15] TF-IDF: eating, behavior, feedback, food, intake
Bi-gram: behavior(al) change, eating rate, eating habits, eating behavior, to improve
We selected three primary databases for our literature retrieval: the ACM Digital Library, IEEE Xplore, and Google Scholar. These sources are considered to comprehensively cover publications closely related to sensing technologies, interactive devices, and the Human–Computer Interaction (HCI) domain. Given the variation in keyword syntax and thesaurus constraints across platforms, we tailored the search strings and scoping targets for each database accordingly. The final keyword formulations and corresponding search results for each source are summarized below. All searches were conducted on April 4, 2025, without restricting the publication time interval.
{mdframed}
[backgroundcolor=gray!15] Keywords (combined with AND):
  • eating-related: eat*, ”food”, diet*, ”intake”
  • intervention-related: interven*, ”feedback”, ”behavior(al) change”, promot*, adjust*, design*, improv*, alter*
  • technology-related: ”wearable”, ”sensor(s)”, ”device(s)”, ”interactive”
ACM Digital Library
  • query target: Title, Keywords, Abstract
  • filters: No keynote, tutorial, invited talk, panel
  • inquiry result: 679
IEEE Explore
  • query target: Document Title (eating), Abstract (eating, intervention), All Metadata (technology)
  • filters: Conferences, Journals
  • inquiry result: 911
Google Scholar
  • filters: No books
  • inquiry result: 100
2.2. Screening
To further narrow down the set of retrieved articles to those aligned with the defined scope of our review, two authors conducted a structured two-stage screening process in accordance with PRISMA guidelines. During the initial screening, which involved title and abstract review followed by full-text assessment, each article was evaluated against a predefined set of inclusion and exclusion criteria. Articles were labeled as “include”, “exclude”, or “uncertain” and any disagreements were resolved through discussion until consensus was reached. The selection criteria used for this process are outlined below and have been adapted to follow PRISMA recommendations.
(1) The paper was not accessible in full text or was not written in English;
(2) The paper was a duplicate or an extended version of a previously included record;
(3) The study focused on behaviors outside our scope or not pertaining to ingestive behavior;
(4) The study did not include any behavioral intervention in the process (e.g., detection-only systems);
(5) The system did not involve any sensor, interactive device, or perceptual computing capability;
(6) The study targeted clinical or therapeutic contexts (e.g., medication ingestion, artificial feeding);
(7) The intervention goal was not health-related (e.g., flavor or entertainment enhancement);
(8) The paper was not a primary research article (e.g., reviews, position papers, dataset descriptions, fictional designs).
Two authors reviewed 1690 relevant publications (679 ACM-DL, 911 IEEE-X, 100 Google Scholar) based on the selecting criteria and confirmed 98 papers (50 ACM-DL, 40 IEEE-X, 8 Google Scholar) for our downstream analysis. Then, we added 9 papers from our previous bibliography and conducted backward-chaining to fill in some papers published in other databases or neglected from keywords. An initial set of papers from ACM and IEEE collections was screened by examining related work sections, and the selections were independently reviewed by another author. 29 additional papers were selected via backward-chain, and among this 8 papers are from IEEE and 7 papers are from ACM, whose missions are due to keywords. Moreover, 8 papers are from Springer, 3 papers are from NIH and other 3 papers are from ResearchGate.
Figure 2. Flow chart of SLR process based on the PRISMA guidelines
Our overall multi-staged screening and selection process is depicted in Figure 2. The final set of review paper set includes 136 papers overall. Annual publication count is shown in Figure 3, illustrating a growing research interest in ingestion health in recent years. We identified 93 distinct publication venues across all included studies, with full details provided in LABEL:tab:venues_list. Among them, the following venues featured two or more included papers: ACM Conference on Human Factors in Computing Systems (CHI, n=16), ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT)/UbiComp (N=16), ACM Interaction Design and Children Conference (IDC, n=2), ACM Designing Interactive Systems Conference (DIS, n=4), IEEE Sensors (n=2), IEEE International Conference on Robot and Human Interactive Communication (RO-MAN, n=2), ACM/IEEE International Conference on Human-Robot Interaction (HRI, n=2), PervasiveHealth (n=2), SIGGRAPH Asia (n=2), International Conference on Tangible, Embedded, and Embodied Interaction (TEI, n=2), Innovations in Power and Advanced Computing Technologies (i-PACT, n=2), International Conference on Computing Communication and Networking Technologies (ICCCNT, n=2), World Congress on Medical Physics and Biomedical Engineering (n=2).
Figure 3. Total number of papers published overall (split to process behavior only, context factor only, and both included)
To facilitate systematic analysis, we created a Google Sheets table with 57 columns. The column design and the taxonomy of core features (corresponding to Chapter 4 and Chapter 5) were iteratively developed. Two authors first defined preliminary categories during the filtering stage, followed by collaborative refinement of the taxonomy through a full review of the papers. Afterward, papers were split among authors for detailed annotation, with further refinements made throughout the reading process.
3. Result Overview
Following the review methodology and the research questions outlined above, this section presents an overview derived from the selected studies, including a classification based on Just-in-Time (JIT), and the primary behavioral closed-loop framework that synthesizes the key stages from sensing to intervention in ingestion health.
3.1. Primary Taxonomy based on JIT
Given the critical role of timing and context-awareness in optimizing the effectiveness of behavioral interventions, as highlighted in previous reviews (Hsu et al., 2025), we classified the included studies based on whether their intervention strategies exhibit JIT characteristics. With a primary overview based on JIT, this classification offers a high-level understanding of the integration mechanisms across existing ingestive behavior interventions, encompassing both sensing and feedback/intervention components.
JIT interventions are defined as approaches that deliver support dynamically at behaviorally or contextually opportune moments, relying on real-time assessments of users’ behavioral, physiological, or environmental states (Nahum-Shani et al., 2018; Sarker et al., 2014). Determining the appropriate timing for intervention typically involves identifying periods when users are vulnerable to adverse health outcomes and are physically, cognitively, and socially available to receive support (Choi et al., 2019; Sarker et al., 2014).
As shown in Figure 4(a), early researches in this domain primarily focused on non-JIT interventions, exhibiting a steady but slow growth over the years. Beginning from 2007, studies incorporating JIT intervention strategies began to emerge, and by 2011, the cumulative number of JIT publications surpassed that of non-JIT publications. A marked acceleration in the growth of JIT-related work can be observed around 2015–2016, which can be attributed to the rapid advancement of wearable sensing technologies and the increasing availability of real-time behavioral and contextual data. The distribution of JIT versus non-JIT studies is further illustrated in Figure 4(b), where JIT interventions account for a substantial majority (73.7%) of the included literature. Together, these findings provide a high-level understanding of the temporal trends and overall characteristics of research in the field of ingestion health interventions.
(a) Cumulative publication count
(b) JIT - No JIT Pie
Figure 4. Cumulative publication count of ingestion health research (a) and publications pie (b).
3.2. Behavioral Closed-loop Framework
After providing an overview of the presence of JIT characteristics in the reviewed literature, we shift our focus to systematically examining behavior sensing and intervention paradigms, structured through a behavioral closed-loop framework. This framework serves as the analytical foundation for addressing the review questions outlined earlier, providing a coherent structure for understanding the current landscape of ingestive behavior interventions throughout the remainder of the paper.
In the broader field of human-computer interaction (HCI), ”closed-loop” architectures are defined as systems in which signals from the user’s brain and body are continuously collected, analyzed, and categorized as inputs to an adaptive controller, which then triggers appropriate responses at the interface (Fairclough, 2015). Such systems facilitate a continuous, bidirectional dialogue between the user and the technology. Furthermore, interaction loops in HCI can be conceptualized as exchanges between mutually interacting agents, where both the user and the system perceive and act upon each other’s states over time (Murray-Smith et al., 2024). In the context of ingestive behavior interventions discussed in this paper, we define ”behavioral closed-loop” as dynamic systems or approaches modeled as exchanges between interacting agents. In these systems, user ingestive behavior is monitored, interpreted, and utilized to deliver adaptive feedback or interventions, with subsequent behavioral changes tracked to iteratively refine future measures. This framework emphasizes the dynamic, responsive, and iterative nature of behavioral closed-loop interventions, capturing the essential stages involved in intervening from perception to ingestive behavior.
Specifically, the stages in behavioral closed-loop system include identifying target behaviors, sensing behaviorally relevant data, interpreting the observed phenomena through reasoning processes, developing feedback or intervention strategies, and monitoring subsequent behavioral adjustments to close the loop. This structure guides the subsequent analysis and provides a detailed understanding of how perceptual computing techniques are applied across different stages of intervention design. An overview of the primary behavioral closed-loop paradigm and its staged evolution is illustrated in Figure 5. While existing studies often focus on specific stages rather than achieving a fully closed-loop design, we structure the subsequent analysis and synthesis based on the ideal of a complete behavioral closed-loop framework.
Figure 5. Behavior closed-loop overview
4. Taxonomy of Target Factors and Behaviors
Daily ingestive behavior interventions can play an important role in promoting healthier eating and drinking habits. The targets of these interventions reflect the key behaviors believed to be critical for promoting healthy eating and improving overall health outcomes. Understanding the specific factors and behaviors targeted in the existing intervention literature is crucial, as they not only highlight important aspects of ingestive behaviors relevant to health, but also provide insight into the intervention strategies designed to address them. By categorizing and analyzing these targeted behaviors, we can better understand how current interventions aim to modify eating habits, assess the effectiveness of different approaches, and identify gaps in existing methods.
This chapter aims to explore the targeted behaviors in current ingestive behavior intervention studies, categorized into contextual factors (external to the human body) and process behaviors (individual actions during the ingestion process). These two broad categories encompass a wide range of interventions designed to optimize various target factors or behaviors for ingestion and promote healthier ingestion habits.
Figure 3 presents trends in the number of intervention studies over the years which can be categorized into three groups: contextual factors only, process behaviors only, and studies targeting both. Both categories have garnered significant attention from researchers, with a greater number of interventions focusing on contextual factors.
Figure 6. Taxonomy of target factors and behaviors with the number of studies involving each target.
Building on this, Figure 6 displays the taxonomy of these target factors and behaviors, providing a structured overview of the specific intervention targets within each category and the number of studies associated with each. This taxonomy organizes the diverse intervention targets into a hierarchical structure, clarifying the relationship between broader categories and more specific target behaviors. A summary of these categories and the corresponding publications included in this review is shown in Table 6.
4.1. Contextual Factors
Contextual factors refer to external influences that affect ingestive behaviors, which are usually not limited to the duration of the ingestion process itself. These factors, which include food choices, intake timing, the environment in which eating takes place, and activity around ingestion process, can shape an individual’s dietary patterns and habits. Since these influences often operate outside the immediate context of the meal, they play a crucial role in determining overall ingestive behaviors and are frequently targeted in interventions aimed at promoting healthier eating habits.
4.1.1. Food Choices
Food choices play a fundamental role in shaping dietary patterns and are directly linked to the types and quality of nutrients consumed, significantly impacting long-term health outcomes. Poor food choices are strongly associated with the development of chronic diseases, such as obesity, diabetes, cardiovascular disease, and certain cancers, while healthier selections promote better metabolic health and disease prevention.
Given the central role of food choices in overall dietary health, many current ingestive behavior interventions prioritize food choices as a primary target. In fact, among the reviewed literature, interventions targeting food choices constitute the largest proportion. These interventions aim to improve dietary quality by modifying nutrient intake, promoting healthier eating habits, and restricting the consumption of unhealthy foods.
Within these interventions, food choice-related targets are commonly categorized into several specific areas: Nutrition Structure Control (aiming for balanced nutrition and dietary variety), Calorie Control (regulating total energy intake), Macronutrient and Micronutrient Control (focusing on specific nutrients such as carbohydrates, fats, sugars, and salt), Limiting Specific Food Categories (e.g., reducing intake of snacks and sugary beverages), and Unspecified Food Choices (general recommendations for healthier eating without specific nutrient guidelines). Among these categories, interventions focusing on nutrition balance and calorie control are particularly prevalent, reflecting their perceived critical role in promoting healthy eating and preventing diet-related chronic diseases.
Nutrition Structure Control: This target behavior focuses on ensuring balanced nutrient intake and promoting dietary variety.
Nutrition structure control interventions focus on fostering a balanced intake of various nutrients by encouraging dietary diversity. Rather than emphasizing individual food items or nutrients, these interventions encourage individuals to consume a wide range of food groups, including fruits, vegetables, whole grains, proteins, and dairy. A significant portion of these interventions specifically target children, with the goal of instilling balanced eating habits early in life. Many studies aim to prevent picky ingestive behaviors and encourage the development of healthy, varied eating patterns from a young age. By promoting a broader range of food intake, these interventions help address nutrient deficiencies and support long-term health maintenance.
Calorie Target Control: This target behavior aims to regulate energy intake through portion size estimation or meal planning.
Calorie control interventions primarily aim to manage overall energy intake to prevent overweight, obesity, and related chronic diseases such as diabetes and cardiovascular conditions. These interventions guide individuals in selecting lower-calorie, nutrient-dense foods, such as fruits, vegetables, and lean proteins. Strategies often include reducing portion sizes and replacing high-calorie foods with healthier alternatives. In the reviewed literature, calorie control is frequently employed as a strategy for weight management and reducing obesity risk, underscoring its central role in ingestion health interventions.
Macronutrient and Micronutrient Control: This category includes interventions targeting the regulation of specific nutrient intakes, such as carbohydrates, fats, sugars, and salt. Many interventions focus on adjusting these components to address health issues like diabetes, hypertension, and cardiovascular diseases. For example, regulating carbohydrate intake is often emphasized for individuals at risk of or diagnosed with diabetes, while reducing fat (particularly saturated fat) and sodium intake is associated with cardiovascular risk reduction.
Current intervention targets include control over specific types of nutrients, such as:
  • Carbohydrate Intake — Regulating or reducing carbohydrate consumption can improve glycemic control and reduce the risk of chronic diseases, particularly among individuals with or at risk of diabetes (Reynolds et al., 2019). Current interventions mainly target individuals with diabetes (Arefeen et al., 2024).
  • Sugar Intake — Excessive sugar consumption is associated with obesity, diabetes, and dental problems (Moynihan, 2016). Current interventions associate sugar intake with glucose level (Liang, 2024).
  • Fat Intake — Reducing total fat intake, especially saturated and trans fats, is essential for mitigating the risk of cardiovascular diseases and obesity, conditions strongly linked to excessive consumption of these unhealthy fats (Astrup et al., 2020). Current interventions often regulate fat intake to prevent obesity (Vandelanotte et al., 2005; Haerens et al., 2007; Burke et al., 2011; Chukwu, 2011).
  • Salt Intake — Limiting sodium intake helps prevent hypertension and cardiovascular risks (Jaques et al., 2021). Current interventions focus on controlling salt intake to promote healthier eating habits in patients with metabolic syndrome, such as those with cardiovascular disease (Eyles et al., 2017; Kim et al., 2016c).
Limiting Specific Food Categories: Some interventions specifically target the reduction or elimination of certain food categories considered unhealthy, such as snacks high in sugar or fat, and sugary beverages like soft drinks and alcohol. By limiting the intake of these foods, the aim is to reduce empty calorie consumption, lower the risk of obesity, and improve overall diet quality. This approach often complements broader nutrition and calorie-focused strategies by removing key sources of dietary excess.
Current research focuses on reducing intake of:
  • Snacks — Interventions discourage high-calorie, low-nutrient snack consumption to reduce unnecessary caloric intake (Bi et al., 2016; Liu et al., 2024; Maramis et al., 2014; Bedri et al., 2022; Robinson et al., 2020).
  • Unhealthy Drinks — Interventions mainly focus on alcohol intake (Matsui et al., 2018; Robinson et al., 2021) or sugar-sweetened soft drink intake control (Faltaous et al., 2021; Haerens et al., 2007; Ritschel et al., 2018; Fuchs et al., 2019).
Unspecific Food Choices: This refers to general recommendations that do not focus on specific nutrients or food categories.
A smaller subset of interventions addresses food choices in a more general way, providing broad guidance to ”eat healthier” without specifying particular nutrients or food categories. These interventions typically focus on encouraging awareness about healthy eating principles and promoting better food decision-making habits, sometimes through educational tools (Prasetyo et al., 2020; Fernandez et al., 2013), reminders (Abisha and Rajalakshmy, 2017; Luo et al., 2021), or behavioral nudges (Huang et al., 2018; Terzioğlu et al., 2023; Eigen et al., 2018).
4.1.2. Intake Timing
Intake timing refers to the timing factors of ingestion, which significantly influence metabolic processes, health, and satiety. Studies have consistently shown that meal timing and frequency are closely linked to overall metabolic health, including insulin sensitivity (Alkhulaifi and Darkoh, 2022). Unhealthy eating patterns, such as irregular meal timing or late-night eating, have been associated with a range of adverse health outcomes. Disruptions to the natural timing of meals can contribute to the development of chronic diseases, including obesity, cardiovascular disease, and diabetes (Yoshida et al., 2018; Rask-Madsen and Kahn, 2012). Additionally, meal frequency — how often a person eats throughout the day — has been found to influence metabolic health. Appropriate meal frequency may promote weight management and improve insulin sensitivity and blood glucose levels (Holmbäck et al., 2010; Carlson et al., 2007). Along with maintaining regular meal times, consistent fluid intake is also essential for supporting bodily functions, such as temperature regulation, nutrient transportation, and waste removal.
In current approaches, target intake timing is typically defined by the regular intervals between meals or water intake for appropriate eating frequency or to prevent dehydration (Xie et al., 2023; Bi et al., 2016; Fortmann et al., 2014; Varun et al., 2023; Lessel et al., 2016; Ravindran et al., 2022), as well as the optimal timing for meals consumption throughout the day to promote regular and healthy eating habits (Liang, 2024; Faltaous et al., 2021; Ye et al., 2016).
4.1.3. Environment
The environment in which an individual eats can significantly impact ingestive behaviors. Both social, cultural, and physical environments contribute to shaping food choices, eating patterns, and the overall eating experience. For example, the availability of healthy food options, social influences, and food cues in the surroundings can either promote or hinder healthy ingestive behaviors. Research has shown that eating location and the surrounding environment can notably affect the intake of unhealthy foods. For instance, higher intakes of unhealthy snacks, such as those rich in sugars, fats, and sodium, are often observed in less structured eating environments like schools and while socializing (Casey et al., 2021).
Current interventions targeting the environment may address behaviors such as eating quickly while standing in the kitchen or not allocating enough time for meals at the dining table, both of which can negatively affect healthy eating. These environmental interventions aim to create supportive eating settings that encourage mindful eating and healthier food choices (Faltaous et al., 2021).
4.1.4. Activity Around Ingestion Process
These interventions involve promoting healthy behaviors surrounding the ingestion process, such as encouraging individuals to engage in physical activity after meals (Alexander et al., 2017).
4.2. Process Behaviors
Process behaviors refer to the specific actions individuals perform during ingestion. Interventions targeting process behaviors typically occur within the duration of an ingestion episode, although they may also span a longer ingestion period. These behaviors, including intake quantity, ingestion pace and oral motor activities, food awareness or emotional state, and safe eating practices, directly impact food intake and digestion. argeting these aspects aims to improve eating habits, prevent overeating, promote mindful consumption, and reduce eating-related safety risks.
4.2.1. Intake Quantity
The quantity of solid food and liquids consumed is crucial for managing energy balance, supporting overall health, and preventing weight-related issues. Adequate regulation of solid food intake helps control energy consumption and maintain a healthy weight, while sufficient liquid intake is essential to prevent dehydration, support physiological functions, and may also influence appetite and caloric intake. Interventions targeting intake quantity typically focus on two key aspects: controlling solid food portion sizes and managing liquid intake to promote healthier eating and drinking behaviors.
Solid Food: In this context, solid food refers to the ingestion of meals or snacks, excluding water and other beverages. By promoting better awareness and control of food quantities, these interventions aim to prevent overeating, reduce the risk of obesity, and support healthy ingestive behaviors, including the prevention of eating disorders (ED) (Rolls et al., 2002; Livingstone and Pourshahidi, 2014).
Interventions targeting solid food portion control often measure intake quantity by assessing changes in food mass before and after a meal (Xie et al., 2023; Chaudhry et al., 2016; Chen et al., 2022; Alshurafa et al., 2015; Scisco et al., 2011; Maramis et al., 2014; Farooq et al., 2017; Moses et al., 2023; Li and Bailenson, 2018).
Liquid: In this category, liquid intake specifically refers to the consumption of water and beverages. Most interventions targeting liquid intake emphasize the importance of maintaining adequate hydration throughout the day. Sufficient fluid intake is crucial for supporting various physiological functions, including temperature regulation, cardiovascular and cognitive health, and preventing conditions such as kidney stones and urinary tract infections. It may also aid in managing energy intake and weight by promoting healthier beverage choices (Popkin et al., 2010; Daniels and Popkin, 2010).
Interventions focusing on fluid intake typically measure water consumption by tracking changes in volume, such as the amount consumed between two time points or the total daily intake (Faltaous et al., 2021; Alshurafa et al., 2015; Kaner et al., 2018; Poddar et al., 2024; Kreutzer et al., 2015; Varun et al., 2023; Bobin et al., 2018; Lessel et al., 2016; Yıldız and Coşkun, 2019; Chiu et al., 2009; Ko et al., 2007; Pankajavalli et al., 2017; Wijanarko et al., 2019; Tommy et al., 2017; Ravindran et al., 2022; Zhou et al., 2021).
4.2.2. Ingestion Pace and Oral Motor Activities
Ingestion pace and oral motor behaviors play a crucial role in regulating energy intake and promoting healthy eating patterns. Interventions targeting these behaviors focus on modifying the speed and manner in which food is consumed to enhance satiety, prevent overeating, and support weight management. Faster eating rates have been consistently associated with higher energy intake, increased risk of overweight and obesity, and poorer appetite regulation. A growing body of evidence supports the effectiveness of slowing down ingestion pace as a practical strategy for preventing obesity and related chronic conditions (Robinson et al., 2014; Slyper, 2021). Current approaches typically target three main aspects: chewing and swallowing behaviors, bite intervals and size, and total meal duration.
Chewing and swallowing behaviors: Chewing and swallowing behaviors are key components influencing the pace of eating and the overall ingestion process. Many interventions target chewing behavior specifically by encouraging individuals to chew more times per bite or to slow down the rate of chewing. By increasing the number of chews per mouthful, individuals can naturally slow their eating pace, enhance satiety signaling, and reduce overall food intake, which are important strategies for managing weight and preventing overeating (Kadomura et al., 2013b; Kleinberger et al., 2023; Koizumi et al., 2011; Bi et al., 2016; Scisco et al., 2011; Maramis et al., 2014; Chen and Yen, 2022; Sugita et al., 2018; Turner-McGrievy et al., 2017; Hori et al., 2023).
Some interventions combine the regulation of chewing and swallowing, aiming to control the full oral processing cycle rather than focusing solely on chewing. These approaches aim to slow ingestion by coordinating the timing between chewing and swallowing, ensuring thorough oral processing before taking the next bite (Kim et al., 2016b; Kamachi et al., 2023).
Additionally, a few interventions focus on modifying oral motor behaviors by adjusting chewing force and modulating the swallowing phase. These interventions promote safer, more mindful eating practices, improving digestion and reducing the risk of eating-related complications (Biyani et al., 2019; Zheng et al., 2025).
Bite Intervals and Size: Bite size and the interval between bites are critical aspects of eating pace regulation and have been targeted in current interventions. This strategy allows more time for satiety signals to develop, helping individuals recognize fullness earlier and thus reducing overall food intake. Most interventions focus on extending the time interval between successive bites, encouraging individuals to pause longer before taking the next bite (Nakamura et al., 2023; de Vries et al., 2020; Lee et al., 2019; Kim et al., 2016a; Kim and Bae, 2018b, a; Mendi et al., 2013; Hermans et al., 2017). In addition to manipulating bite intervals, the amount of food consumed in a single bite is also a key target behavior. By reducing bite size, individuals naturally slow down their eating pace and consume fewer calories during a meal (Chen et al., 2022).
Total Meal Duration: Total meal duration — the overall time taken to complete a meal — is another important target for interventions aimed at regulating eating pace. Most studies monitor and estimate total meal duration by tracking changes in food weight over time, providing an objective measure of ingestion pace. Many interventions focus on prolonging meal duration to slow down eating speed, helping individuals better recognize satiety cues and prevent overeating, which supports healthier dietary behaviors (Moses et al., 2023; Zhang et al., 2019; Kim et al., 2016c; Ioakimidis et al., 2009). Some interventions, particularly those targeting young children, aim to prevent meals from extending excessively long. These interventions help children establish structured, healthy eating routines early in life (Krishna et al., 2025; Lo et al., 2007). Additionally, some interventions are designed to synchronize eating paces in social settings, where regulating total meal duration across a group helps reduce discomfort caused by uneven eating speeds (Mitchell et al., 2015).
4.2.3. Food Awareness or Emotional State
Interventions targeting food awareness or emotional state focus on enhancing individuals’ cognitive and emotional engagement with their eating experiences to promote healthier behaviors. A key approach in this area is encouraging greater attention to food during meals, either through practices like mindful eating, which emphasizes being fully present during eating (Kleinberger et al., 2023; Koizumi et al., 2011; Parra et al., 2023; Wang et al., 2025; Khot et al., 2020), or through interventions designed to help children develop better focus on their food (Kadomura et al., 2013b; Han and Kang, 2017). In addition to promoting awareness during ingestion, some interventions focus on improving individuals’ self-perception and cognitive evaluation of their ingestive behaviors after meals. By increasing reflective thinking about eating habits, these interventions help individuals build healthier long-term dietary patterns and strengthen their self-regulation capacity (Sun et al., 2020; Zhang and Parker, 2020; Tang et al., 2015). Moreover, another line of interventions aims to enhance the emotional quality of the eating experience itself, such as promoting positive emotions through novel experiences and reducing negative emotions like anxiety during eating. A healthy emotional state during ingestion has been linked to better digestion, improved dietary choices, and a lower risk of disordered ingestive behaviors (Mayumi et al., 2023; Carroll et al., 2013; Mitchell et al., 2015; Narumi et al., 2011).
4.2.4. Safe Eating Practices
Safe eating practices aim to ensure that the ingestion process does not pose additional health risks. Current interventions in this area primarily target two aspects: hygiene behaviors and food temperature control. Hygiene-related interventions focus on promoting safe practices during eating to prevent foodborne infections (Wang and Zhang, 2021). Other interventions address food temperature control to prevent oral injuries, such as burns (Singhal et al., 2024).
5. Structures and Design Space for Behavioral Sensing and Intervention Paradigms
This chapter analyzes the structural and modality-level characteristics of ingestion health paradigms, as outlined in Chapter 3.2, with a focus on two essential stages of the behavioral loop: sensing and feedback/intervention. While reasoning processes typically mediate between perception and action, they are often embedded within system pipelines, reflected through how sensing data is interpreted and applied. Therefore, our analysis centers on the sensing and feedback/intervention components, where modality design plays a direct and observable role.
Inspired by Jaimes and Sebe (2007) and Jansen et al. (2022), we adopt modality as the classification standard for both the sensing and intervention components, encompassing interactive devices and perceptual computing methods. Specifically, Jaimes and Sebe (2007) define modality as “mode of communication according to human senses and computer input devices activated by humans or measuring human qualities,” emphasizing that modality serves as a communication channel designed around human perception and computational input-output mechanisms. Furthermore, they introduce a human-based perspective on modality, identifying core modalities corresponding to human senses, including vision, touch, hearing, smell, and taste, and categorizing more complex examples within these sensory categories. Building on this foundation, Jansen et al. (2022) expand the concept of human-based modality by introducing two additional categories—cerebral and cardiac modalities—to accommodate advances in interactive device technologies.
Following insights from Jaimes and Sebe (2007), we recognize that, in the sensing process, certain modalities do not directly correspond to individual human sensory channels but involve interactions mediated through external devices or environmental signals. Thus, we introduce the category of environment-based modalities, capturing input and feedback mechanisms targeting environmental objects independently of human senses. This extension provides a more comprehensive classification for sensing and intervention designs in ingestion health systems, covering modalities arising from device manipulation, environmental interactions, or indirect physiological monitoring.
We adapt and refine this modality classification to suit ingestion health interventions, distinctly addressing both sensing and feedback/intervention stages. The primary classification human-based modality contains visual, auditory, haptic, gustatory, olfactory, motion, and physiological, and environment-based modality contains Visual, Acoustic, Force, Spectral, and Electrical. Almost each modality further includes submodalities considering focus and implications different from sensing and feedback/intervention. Accordingly, slight structural and descriptive variations in modality classifications are introduced to align precisely with each phase of the behavioral closed-loop, as shown in following subsections.
5.1. Sensing
This subsection focuses on the sensing modalities within overall ingestion health intervention paradigms, where behavioral data are primarily acquired through various sensor-based or perceptual computing-driven modalities. The detailed structure and descriptions of each modality, along with its associated submodalities in the sensing process, are summarized in Table 1.
Notably, gustatory and olfactory modalities are absent in the sensing stage across existing systems. We believe that sensing gustatory or olfactory information through perceptual or sensor-based means remains technically challenging, and these modalities are mainly employed in the feedback/intervention phase rather than for direct behavioral sensing.
The following sections present the identified sensing modalities, categorized according to their classifications, with representative approaches within ingestion health intervention paradigms.
Table 1. Descriptions and category of modality terms in sensing process
Modality Submodality Description Paper count
Visual Human-based visual captures human-derived visual cues to infer the user’s attention, emotional state, or physical actions during ingestion-related behaviors 8
[…]
Facial expression Facial expression analyzes dynamic changes in facial muscle movements to detect emotional responses, physiological states, or behavioral reactions associated with eating activities 4
Body posture Body posture captures the orientation, alignment, and movement of the user’s body or head to interpret ingestion-related behaviors via camera, such as eating gestures, drinking actions, or body positioning during meals 3
[…]
Sound Sound captures non-verbal acoustic signals generated during ingestion activities, such as chewing sounds, swallowing noises, or utensil interactions, to infer ingestive behaviors or consumption events 8
[…]
Chewing and Micro-Head Motion Chewing and micro-head motion monitors rhythmic jaw movements and subtle head or neck dynamics associated with eating, drinking, or swallowing 5
[…]
Metabolic Metabolic captures nutritional and biochemical markers such as blood glucose, cholesterol, and body fat rate, typically using biosensors to assess metabolic responses and long-term ingestion-related health status 6
Human-based Autonomic Autonomic captures involuntary signals, including electrodermal activity (EDA) and skin temperature, using biosensors to monitor autonomic activity during ingestion 7
Visual Environment-based visual captures information about food items, environments, or objects through image-based techniques such as classification, segmentation, or object detection 29
[…]
Environment-based Magnetic Magnetic sensing detects the presence, identity, or proximity of ingestion-related objects using techniques such as Radio-Frequency Identification (RFID) or Near-Field Communication (NFC) technologies 5
5.1.1. Human-based
Visual: Visual sensing has been widely adopted in ingestion-related systems to extract behavioral and attentional cues through camera-based observation. We identified three common submodalities within visual sensing: gaze, facial features, and body movement.
Gaze-based sensing tracks eye movements and fixation points to assess users’ attention during food-related interactions. For example, Karkar et al. (2018) integrated an eye tracker into a VR system to analyze children’s visual attention toward nutrition-themed stimuli. Webcam-based gaze estimation has also been used to infer attentional disengagement during meals, which can inform timely interventions (Khot et al., 2020).
Facial-expression-based sensing is primarily used to detect emotional or behavioral responses and confirm user presence. A webcam was employed to extract facial expressions and assess affective states during meals (Mayumi et al., 2023). Moreover, Kim and Bae (2018a) used lip and jaw motion analysis to model bite rhythm. Face detection has also supported social presence recognition in multi-user contexts, helping to determine whether a child is present (Triantafyllidis et al., 2019) or enabling socially responsive camera-based interaction (Robinson et al., 2020).
Body-based visual sensing targets ingestion-related gestures and postures, such as hand movements and upper-body orientation. Xie et al. (2023) tracked caregivers’ feeding gestures to estimate feeding frequency and potential satiety cues in children. Skeleton tracking with depth cameras has been used to monitor upper-body posture during meals (Mayumi et al., 2023). Additionally, egocentric video data were collected for post-hoc analysis of mealtime behaviors and social dynamics (Zhang and Parker, 2020).
Auditory: Auditory sensing in ingestion health systems typically involves capturing either verbal input or eating-related sounds, which can be classified into two main submodalities: speech and non-verbal sound, covering existing approaches.
Speech-based auditory sensing is commonly used for semantic data acquisition or enabling conversational interfaces. Spoken input has been leveraged to extract structured dietary information using linguistic pattern recognition (Prasetyo et al., 2020). Systems such as those proposed by Mayumi et al. (2023) and Robinson et al. (2020) receive users’ queries for food-related facts or recommendations and extract key information from speech signals. Speech has also been employed as an input modality for logging intake events or interacting with nutrition tracking systems (Luo et al., 2021; Arefeen et al., 2022). Parra et al. (2023) used Wizard-of-Oz setups and processed users’ responses for further system-guided prompting.
Sound-based sensing focuses on detecting physiological cues, such as chewing or swallowing, using microphones. Some researchers have placed microphones near the throat (Kalantarian et al., 2014) or collar (Chen and Yen, 2022) to capture oral sound signals during ingestion, which can be analyzed for bite detection (Bi et al., 2016), food classification (Kleinberger et al., 2023), or intake rhythm estimation. Microphones have also been implemented in earpiece form factors to detect rhythmic mastication events, such as chewing sensors (Maramis et al., 2014) and bone conduction microphones (Kamachi et al., 2023), which capture sound with reduced ambient interference.
Additionally, although categorized separately above, some microphones are designed for specific contexts, such as capturing chewing in denture wearers (Koizumi et al., 2011) or jointly recording speech and chewing sounds as contextual data (Zhang and Parker, 2020; Kamachi et al., 2023).
Haptic: Haptic sensing aims to detect physical interactions or internal mechanical activity through pressure, force, or contact-based sensors. A common form is weight sensing, which tracks changes in body weight as a long-term reflection of dietary intake (Ludwig et al., 2011; Bastida et al., 2023; Nedungadi et al., 2018).
A substantial body of work has focused on haptic sensing using piezoelectric or force sensors. Neck-mounted piezoelectric sensors have been embedded into wearable necklaces to capture throat vibrations associated with swallowing (Alshurafa et al., 2015; Kim et al., 2016b; Abisha and Rajalakshmy, 2017). Similar techniques have been used to detect mechanical displacement in the temporalis muscle to infer chewing activity (Farooq et al., 2017). Other studies have explored intraoral pressure detection, such as placing sensors on palatal retainers to measure tongue force (Biyani et al., 2019), or using external force sensors to estimate masticatory force (Zheng et al., 2025). Additionally, pressure-sensitive buttons or interfaces have been applied to allow users to log ingestion-related events (Fortmann et al., 2014; Tang et al., 2015; Cai et al., 2024; Robinson et al., 2020).
Motion: Motion sensing captures gestures and subtle physical activity using inertial or proximity sensors.
Most gesture-oriented systems employ inertial measurement units (IMUs) mounted on utensils (Wang et al., 2025; Khot et al., 2020; Zhang et al., 2019; Li and Bailenson, 2018; Kadomura et al., 2013a, 2014), wristbands (Kim et al., 2016a; Kim and Bae, 2018b; Ye et al., 2016; Turner-McGrievy et al., 2017; Mendi et al., 2013; Kumbhare et al., 2023; Gaikwad et al., 2024; Dragoni et al., 2017; Scisco et al., 2011), or containers (Tommy et al., 2017; Bobin et al., 2018; Chiu et al., 2009; Zhou et al., 2021) to detect hand-to-mouth movement, bite frequency, or drinking actions. IMUs have also been integrated into smart forks (Hermans et al., 2017), smart glasses (Bedri et al., 2022), and smartphones (Tirasirichai et al., 2018). Additionally, some platforms or frameworks use pedometers (Harous et al., 2017) and activity trackers (Alexander et al., 2017; Vazquez et al., 2015; Fuchs et al., 2019) to estimate ingestion or physical activity context. These approaches have made it easier for broader health-monitoring systems to achieve dietary reasoning support (Bastida et al., 2023; Mishra et al., 2024; Maramis et al., 2014; Chukwu, 2011; Badawi et al., 2012a; Jung et al., 2007).
Chewing and micro-head motion sensing focus on jaw, throat, or mouth-level interactions. Biyani et al. (2019) used a 3D accelerometer to track throat motion, and Hori et al. (2023) applied infrared distance sensing to capture subtle skin vibrations. Additionally, electrical conductivity sensors (Kadomura et al., 2013a, b) or static charge sensors (Han and Kang, 2017) were embedded in utensils or cups to infer mouth contact and drinking actions.
Physiological: Physiological sensing provides insight into users’ internal states that influence or contextualize ingestion behavior, including appetite, metabolic status, stress, and cardiovascular conditions.
Autonomic sensing typically involves EMG electrodes to detect muscle activity during chewing or wrist motion (Lee et al., 2019; Sugita et al., 2018; Kim et al., 2016b), as well as thermistors and respiratory sensors to track body temperature and breathing patterns (Park and Min, 2007; Jung et al., 2007). Emotional arousal or stress is commonly inferred through electrodermal activity (EDA) sensors (Carroll et al., 2013; Nedungadi et al., 2018), which may affect eating pace or impulsive behaviors.
Metabolic sensing employs BIA sensors to estimate body fat composition (Chukwu, 2011; Balbin et al., 2020), and incorporates CGM and cholesterol sensors to continuously monitor glucose levels and lipid markers (Liang, 2024; Harous et al., 2017; Arefeen et al., 2024; Nedungadi et al., 2018).
Cardiac sensing captures heart rate, blood pressure, and oxygen saturation using ECG (Badawi et al., 2012b), pulse wave (Varun et al., 2023), or optical PPG sensors (Park and Min, 2007; Ludwig et al., 2011; Jung et al., 2007; Huang et al., 2018; Nedungadi et al., 2018), which are frequently used to contextualize ingestive behavior or guide personalized intervention timing.
Others: Apart from the sensing modalities detailed above, there are several approaches that do not involve sensing or perceptual computing methods. These approaches rely entirely on user-reported data collected through manual input. Common input methods include the manual entry of dietary events or physical status, image selection or annotation, and questionnaire-style logging via mobile or chatbot interfaces. Although these systems may still provide feedback, they fall outside our sensing modality framework due to the lack of automated perceptual input or computational sensing logic. A detailed overview of this manual approach is summarized in Table 7, with approximately 20 publications adopting this method.
5.1.2. Environment-based
Visual: Environment-based visual sensing focuses on object-level perception of food and eating-related artifacts to support dietary monitoring and context acquisition. Most systems utilize food image analysis, applying computer vision to classify food types, estimate portion sizes, or infer nutritional content from captured images (Kanjalkar et al., 2024; Eigen et al., 2018; Cordeiro et al., 2015; YR et al., 2024; Epstein et al., 2016). These methods are commonly implemented through smartphone photography (Mishra et al., 2024; Waltner et al., 2015; Achilleos et al., 2017), utensil-mounted or refrigerator cameras (Nakamura et al., 2023; Kwon et al., 2016; Faltaous et al., 2021), or wearable vision modules (Fuchs et al., 2019).
Some systems use food color (Kadomura et al., 2013a, 2014), utensil-based visual cues (Wang and Zhang, 2021), or visual distortion sensing (Akter et al., 2021; Chiu et al., 2009) to infer consumption context or eating correctness. Other approaches incorporate plate-level dynamics (Ganesh et al., 2014), receipt parsing (Mankoff et al., 2002), puzzle-based interaction (Cai et al., 2024), or packaging and interface-based food input (Arefeen et al., 2022; Hsiao et al., 2011) for supplemental inference and logging. Additionally, visual code recognition is widely used to access structured nutritional information through barcode scanning (Volkova et al., 2016; Eyles et al., 2017), QR code interpretation via mobile interfaces (Fernandez et al., 2013; Vazquez et al., 2015; Alloghani et al., 2016), and edible marker detection in AR-enhanced scenarios (Gutiérrez et al., 2018).
Acoustic: Acoustic sensing often utilizes ultrasonic distance sensors to monitor liquid intake or food-related actions. Most related works involve embedding ultrasonic transducers into smart bottles or storage containers to measure fluid levels and infer hydration behavior (Poddar et al., 2024; Wijanarko et al., 2019; Ravindran et al., 2022; Varun et al., 2023; Tommy et al., 2017). These systems estimate drinking volume based on real-time fluid height or remaining capacity. Other systems apply similar sensing techniques to detect discrete ingestion-related events, such as the removal of a snack from a container using ultrasonic signal disruption (Liu et al., 2024), or monitoring food depletion in a storage cabinet based on internal distance measurements (Faltaous et al., 2021).
Force: Force sensing typically involves the use of load cells or pressure sensors embedded in plates, coasters, or tables to measure changes in food or liquid weight over time. This approach enables passive monitoring of eating dynamics, such as intake amount, bite frequency, or consumption rate. Smart plates equipped with load cells (Mitchell et al., 2015; Kim et al., 2016c; Maramis et al., 2014; Ioakimidis et al., 2009; de Vries et al., 2020) or force-sensitive resistors (FSRs) (Jung et al., 2017) have been used to continuously track food depletion and infer eating pace or satiety.
Beyond intake tracking, similar sensing mechanisms have been extended to estimate nutrient intake across meals (Krishna et al., 2025; Zhao et al., 2021; de Vries et al., 2020; Jung et al., 2017), monitor liquid consumption via drink weight changes (Ritschel et al., 2018; Lessel et al., 2016), or support playful interactions through weight-sensitive input (Lo et al., 2007). Additionally, multi-plate setups have been applied to detect food preferences and portion control behaviors (Joi et al., 2016; Hermans et al., 2017).
Spectral: Spectral sensing leverages temperature, proximity, and light-based sensors to capture physical or chemical properties of food, beverages, and environmental conditions. Common sensing components include infrared proximity detectors, temperature sensors, and light-spectrum analyzers. Specifically, infrared or optical sensors have been used to detect the presence of containers (Ritschel et al., 2018), estimate proximity to food (Bedri et al., 2022), or assess food temperature (Singhal et al., 2024). Temperature sensors embedded in bottles or utensils help monitor liquid warmth (Tommy et al., 2017; Wijanarko et al., 2019; Varun et al., 2023) or environmental conditions affecting food stability (Bastida et al., 2023; Badawi et al., 2012b; Kwon et al., 2016). In more advanced cases, Matsui et al. (2018) applied spectral absorption analysis to estimate the alcohol concentration in beverages and assess drinking speed.
Electrical: Electrical sensing supports ingestion-related monitoring by detecting changes in conductivity, capacitance, or magnetic response, allowing for the inference of liquid consumption, food identity, or environmental conditions. This method enables passive, contact-based tracking of ingestion without requiring user instrumentation. Common sensor types include conductivity probes, humidity sensors, capacitive pads, and RFID/NFC modules.
Specifically, conductivity and humidity sensors have been embedded into bottles or utensils to monitor water levels or detect drinking events (Bobin et al., 2018; Tommy et al., 2017; Kreutzer et al., 2015; Varun et al., 2023), while capacitive level sensing has been used to estimate water volume based on changes in dielectric properties (Kaner et al., 2018). Some systems integrate float switches or conductive tilt detection for recognizing bottle use (Pankajavalli et al., 2017; Ko et al., 2007). Ambient humidity is also used to characterize dining environments and inform food safety assessments (Bastida et al., 2023; Badawi et al., 2012b; Kwon et al., 2016).
On the magnetic side, RFID and NFC technologies enable automatic identification of food items, storage states, or utensil presence (Vazquez et al., 2015; Huang et al., 2010; Jung et al., 2017), and advanced NFC designs even estimate consumed volume through passive milliliter tracking (Yıldız and Coşkun, 2019). Additionally, Han and Kang (2017) used metal contact detection to support real-time eating activity sensing.
5.2. Feedback and Intervention
As the final stage in the ingestion behavior loop, feedback and intervention mechanisms convey information back to users to support behavioral awareness, regulation, or change. In contrast to sensing, which focuses on data acquisition through sensors or perceptual computing, intervention emphasizes human interpretation and experience. Accordingly, the modalities at this stage are inherently human-based, as they operate by stimulating human sensory channels to deliver prompts, guidance, or other affective cues.
To systematically present the design of the feedback and intervention process, we do not assess whether these mechanisms rely on dedicated sensors. Rather, since all systems included in our review engage in sensor-based or perception-driven processing at some point in the ingestion-related loop, we categorize their feedback and intervention strategies in terms of modality, focusing on how information is conveyed and perceived. These modalities are classified based on the primary human sense they target, as summarized in Table 2.
Table 2. Descriptions and category of modality terms in feedback/intervention process
Modality Submodality Description Paper count
Visual Textual Delivering feedback through text-based information, such as messages, notifications, or instructions 85
Graphical Providing feedback using graphical elements like charts, progress bars, icons, or visualizations to convey behavioral states or intervention cues 41
Symbolic Giving visual feedback through abstract or simplified representations, such as symbols, flashing lights, or color changes, to convey states or prompts with minimal cognitive load or linguistic interpretation 21
[…]
Deformation Delivering feedback by altering the physical shape, structure, or volume of an interface, allowing users to perceive behavioral cues through tangible disruptions embedded within their natural interaction actuators 7
[…]
5.2.1. Visual
Visual feedback is one of the most commonly used modalities in ingestive behavior interventions, providing intuitive, accessible, and scalable means of conveying prompts, progress, or state awareness. Based on representation format, we categorize visual feedback into three submodalities: textual, graphical, and symbolic. This layered approach to visual design allows systems to balance clarity, immediacy, and ambient subtlety, tailoring feedback complexity to user needs and situational demands. A complete list of visual feedback systems and their classified modalities is available in Table 8.
Textual: Textual submodality presents information explicitly through written content such as notifications, reminders, or dietary instructions. These systems often use text-based prompts to communicate recommended actions, reflect behavioral progress, or reinforce decision-making. Examples include recommender agents (Zenun Franco, 2017; Arefeen et al., 2022), just-in-time eating assistants (Prasetyo et al., 2020; Carroll et al., 2013), and mobile self-monitoring tools (Burke et al., 2011; Krishna et al., 2025), all of which deliver messages through on-screen text. Many additional systems also employ message notifications or app-based dashboards to present dietary data or track goals in textual form, reinforcing the prominence of this submodality in behavior change design (Faltaous et al., 2021; Lee et al., 2019; Haerens et al., 2007).
Graphical: Graphical methods employ structured visual elements—such as charts, progress bars, or data visualizations—to represent behavior patterns or intervention outcomes. For instance, eat2pic translates bite patterns into abstract visual paintings as a playful form of real-time feedback (Nakamura et al., 2023), while many other systems use app interfaces to graph nutritional balance, meal composition, or eating speed, enhancing user self-awareness and reflection (Faltaous et al., 2021; Scisco et al., 2011; Lo et al., 2007).
Symbolic: Symbolic approaches convey prompts using abstract or simplified visual representations, such as color changes (Bobin et al., 2018), flashing lights (Matsui et al., 2018; Lessel et al., 2016), or emoji-style icons (Kim and Bae, 2018a), which communicate behavioral cues with minimal cognitive load. These signals are often embedded in utensils or ambient displays to indicate states like “ready to eat” or “slow down” without requiring linguistic interpretation (Agarwal et al., 2024; Kim et al., 2016b).
5.2.2. Auditory
Auditory modalities primarily rely on simple sound output devices, most commonly speakers embedded in utensils, tableware, or ambient systems. These systems use auditory feedback to deliver symbolic or affective cues that influence perception, pacing, or user awareness during meals. For instance, enhanced chewing sounds or filtered auditory effects have been applied to modulate food perception and reduce eating speed through cross-modal interaction (Kleinberger et al., 2023). Interactive forks or tableware generate rhythmic or instructive sounds, such as audio riddles or feedback tones, to encourage mindful chewing or promote healthy eating habits in children (Kadomura et al., 2013b; Cai et al., 2024). Auditory reminders are also embedded in game-based or robot-assisted scenarios to regulate intake timing (Xie et al., 2023; Kadomura et al., 2014). A complete list of studies employing auditory approaches can be found in Table 8. Compared to the visual channel, auditory cues provide a lightweight, ambient feedback pathway that can be unobtrusively integrated into multisensory interventions.
5.2.3. Haptic
Haptic modality involves bodily interventions using physical signals such as vibration or shape deformation. Instead of relying solely on visual or auditory channels, this approach engages the user’s sensorimotor experience by delivering tactile cues, modifying utensil mechanics, or altering physical resistance to shape real-time food-related actions.
Vibration: Most systems employ vibration motors or piezoelectric actuators. For instance, smart bottles vibrate when hydration targets are unmet (Faltaous et al., 2021; Varun et al., 2023), chopsticks trigger vibrations upon improper placement (Wang and Zhang, 2021), and wristbands or smartwatches vibrate in response to fast eating or insufficient chewing (Kim et al., 2016b, a; Kim and Bae, 2018b; Kamachi et al., 2023). Some systems provide vibration cues for journaling reminders (Ye et al., 2016), while others use mobile phone vibrations as hydration prompts (Chiu et al., 2009). In utensil-based designs, vibrotactile feedback is provided when eating speed exceeds a threshold (Hermans et al., 2017). Overall, vibration offers an embodied, non-disruptive alternative to visual or auditory alerts.
Deformation: Deformation-based feedback or intervention modifies the physical properties of utensils or wearable devices to guide ingestive behavior. These systems typically implement adjustable deformation through pneumatic or motorized actuation to enable conscious or subconscious intervention during eating. For example, inflatable wearables like the FatBelt simulate abdominal fullness after caloric thresholds to promote self-awareness (Pels et al., 2014), while utensil designs such as the SSpoon (Chen et al., 2022), SWAN (Khot et al., 2020), a pneumatically bending spoon (Zhang et al., 2019), or a motor-stabilized spoon for Parkinson’s patients (Singhal et al., 2024) deform or adjust posture to hinder rapid intake or support stable consumption. In contrast, other approaches rely on materials with inherent deformability, such as VR rubber donuts (Li and Bailenson, 2018) or rubbing scent-releasing tactile surfaces (Cai et al., 2024).
5.2.4. Olfactory
Olfactory modality utilizes scent stimuli to influence perception, modulate appetite, or reinforce associations with food-related behaviors. All studies related to this modality adopt the strategy of intentionally delivering odor cues to induce behavioral or perceptual changes. Methods range from digitally coordinated aroma delivery to embedded scent-triggering materials. Li and Bailenson (2018) used aromatic oil diffusers synchronized with VR scenes to emit food-related smells, enhancing satiation and reducing intake. Narumi et al. (2011) developed an olfactory display that filtered or overlaid real food odors to shift perception and promote healthier choices. Cai et al. (2024) embedded scent-infused ink in puzzle pieces that released odors through friction to support children’s food recognition. Collectively, these interventions illustrate the potential of olfactory feedback to modulate user experience on both affective and cognitive levels within ingestion-related contexts. Although relatively underexplored, scent-based strategies offer a promising channel for multisensory engagement, particularly when integrated with complementary modalities such as visual or tactile input.
5.2.5. Physiological
Among physiological modalities, electrodermal stimulation is the only technique applied in ingestion-related interventions. These systems use intraoral electrical pulses to modulate muscle activity in a closed-loop manner, responding to physiological events. Zheng et al. (2025) delivered synchronized stimulation to the mandibular vestibule based on mastication force, enhancing tactile feedback and chewing strength. Biyani et al. (2019) used tongue pressure to trigger soft palate stimulation during swallowing, aiming to reduce pharyngeal transit time and improve coordination. These sensorimotor interventions directly target neuromuscular regulation rather than perceptual awareness. Although rarely explored, such systems offer a promising path for low-latency modulation of ingestion behavior.
5.3. Design Space Mapping through Modality to Behavior
To further characterize the distribution and coverage of closed-loop approaches for sensing and intervention in ingestion health, we conducted a morphological analysis combining the dimensions of sensing and feedback/intervention modalities with their corresponding target behaviors, inspired by Zwicky (1967) and Jansen et al. (2022).
This results in two-dimensional design matrices, shown in Table 3 and Table 4, which map the occurrence of each modality (x-axis) against specific behavioral targets (y-axis). To reduce redundancy and enhance interpretability, we adopt a mixed-level classification along the modality axis: for modalities with numerous fine-grained subcategories and overlapping device implementations, we collapse them into higher-level groupings to better reflect functional similarity and design relevance. The matrices enable us to identify well-covered modality–behavior pairs, underexplored combinations, and potential design opportunities. We hope this visualization can serve as a generative framework to inspire new behavior-aware intervention designs in ingestion health.
Analysis of Table 3 reveals several key trends in the current design landscape in sensing process. Human-based motion sensing, particularly wrist and object trajectory tracking, demonstrates broad behavioral coverage in oral and finger motor activities and some other context factors. Visual and acoustic modalities are also well represented, especially for food recognition, intake timing, and ingestion pace, often leveraging image, sound, or AR-based inputs.
By contrast, environment-based modalities like electrical or magnetic sensing remain underutilized, appearing in only a few cases. Physiological sensing (e.g., EMG, EDA, CGM) offers rich internal state signals but is mostly limited to specific domains such as emotional state or metabolic monitoring. This suggests potential in integrating these modalities for broader behavioral targets.
The matrix also highlights underexplored areas, including food hygiene, emotional triggers, and environmental dynamics. These gaps point to opportunities for multimodal fusion and novel device placements. Overall, the mapping reveals both saturated modality–behavior pairs and areas where new sensing strategies can enhance behavior-aware intervention design.
Table 3. Mapping of Sensing Modalities to Target Factors and Behaviors in Ingestion Health
Sensing Modality
Human-based Environment-based
Physiological Electrical
Visual Auditory Haptic Motion Cardiac Metabolic Autonomic Visual Acoustic Force Spectral Conductivity Magnetic
Food Choices facial expression (Robinson et al., 2020; Triantafyllidis et al., 2019), gaze (VR) (Karkar et al., 2018) speech (Prasetyo et al., 2020; Arefeen et al., 2022; Luo et al., 2021; Robinson et al., 2020), intake sound (Bi et al., 2016; Maramis et al., 2014; Kalantarian et al., 2014) touch (Cai et al., 2024; Robinson et al., 2020), oral behavior (Abisha and Rajalakshmy, 2017), body weight (Harous et al., 2017; Ludwig et al., 2011; Bastida et al., 2023; Nedungadi et al., 2018) wrist trajectory (Faltaous et al., 2021; Dragoni et al., 2017; Maramis et al., 2014), Object trajectory (Kadomura et al., 2013b; Nakamura et al., 2023; Joi et al., 2016; Kadomura et al., 2014, 2013a; Bedri et al., 2022), exercise (Kumbhare et al., 2023; Mishra et al., 2024; Harous et al., 2017; Dragoni et al., 2017; Maramis et al., 2014; Bastida et al., 2023; Chukwu, 2011; Badawi et al., 2012a; Fuchs et al., 2019; Tirasirichai et al., 2018; Jung et al., 2007; Vazquez et al., 2015) blood pressure (Huang et al., 2018; Jung et al., 2007; Ludwig et al., 2011; Park and Min, 2007; Nedungadi et al., 2018), heart rate (ECG) (Jung et al., 2007; Park and Min, 2007; Badawi et al., 2012b; Nedungadi et al., 2018), blood oxygen (Park and Min, 2007; Nedungadi et al., 2018) CGM (Liang, 2024; Arefeen et al., 2024; Nedungadi et al., 2018), BIA (Balbin et al., 2020; Chukwu, 2011), cholesterol (Harous et al., 2017) temperature (Park and Min, 2007; Nedungadi et al., 2018; Jung et al., 2007) food picture (Faltaous et al., 2021; Nakamura et al., 2023; Cai et al., 2024; Kadomura et al., 2014, 2013a; Ganesh et al., 2014; Arefeen et al., 2022; YR et al., 2024; Mishra et al., 2024; Alloghani et al., 2016; Hsiao et al., 2011, 2011; Achilleos et al., 2017; Kanjalkar et al., 2024; Kwon et al., 2016; Eigen et al., 2018; Cordeiro et al., 2015; Waltner et al., 2015; Epstein et al., 2016; Akter et al., 2021), Nutrition Label (Hsiao et al., 2011; Fernandez et al., 2013; Gutiérrez et al., 2018; Fuchs et al., 2019; Volkova et al., 2016; Eyles et al., 2017; Mankoff et al., 2002; Vazquez et al., 2015) food removal (Liu et al., 2024) Meals and plate weight (de Vries et al., 2020; Joi et al., 2016; Jung et al., 2017; Zhao et al., 2021; Ritschel et al., 2018; Krishna et al., 2025; Maramis et al., 2014; Kim et al., 2016c) alcohol (Matsui et al., 2018), temperature (Bastida et al., 2023; Kwon et al., 2016; Badawi et al., 2012b), object state (Ritschel et al., 2018), food distance (Bedri et al., 2022) humidity (Bastida et al., 2023; Kwon et al., 2016; Badawi et al., 2012b)
Intake Timing intake sound (Bi et al., 2016) button click (Fortmann et al., 2014) wrist gesture (Ye et al., 2016) CGM (Liang, 2024) heart rate (Varun et al., 2023) liquid level (Varun et al., 2023) food weight (Lessel et al., 2016) temperature (Varun et al., 2023) humidity (Varun et al., 2023)
Environment indoor position (Faltaous et al., 2021)
Context Factors Activity Around Ingestion postprandial exercise (Alexander et al., 2017)
Intake Quantity body gesture (Xie et al., 2023) oral behavior (Farooq et al., 2017) bite count (Scisco et al., 2011), object trajectory (Zhou et al., 2021; Li and Bailenson, 2018) food picture (Chiu et al., 2009) leftovers (Faltaous et al., 2021), liquid level (Varun et al., 2023; Tommy et al., 2017; Ravindran et al., 2022) liquid level (Lessel et al., 2016) liquid level (Poddar et al., 2024) liquid level (Kaner et al., 2018; Kreutzer et al., 2015; Ko et al., 2007; Pankajavalli et al., 2017; Wijanarko et al., 2019), alcohol (Bobin et al., 2018) liquid level (Yıldız and Coşkun, 2019)
Ingestion Pace and Oral Motor Behavior facial expression (Kim and Bae, 2018a) intake sound (Kleinberger et al., 2023; Koizumi et al., 2011; Bi et al., 2016; Maramis et al., 2014; Kamachi et al., 2023; Chen and Yen, 2022) oral behavior (Kim et al., 2016b; Biyani et al., 2019; Zheng et al., 2025) wrist trajectory (Mendi et al., 2013; Lee et al., 2019; Scisco et al., 2011; Kim et al., 2016a; Kim and Bae, 2018b; Maramis et al., 2014; Turner-McGrievy et al., 2017), object trajectory (Kadomura et al., 2013b; Nakamura et al., 2023; Biyani et al., 2019; Hermans et al., 2017; Zhang et al., 2019), skin distance (Hori et al., 2023) wrist action (EMG) (Lee et al., 2019; Kim et al., 2016b; Sugita et al., 2018) food weight (de Vries et al., 2020; Krishna et al., 2025; Lo et al., 2007; Kim et al., 2016c; Ioakimidis et al., 2009; Mitchell et al., 2015; Hermans et al., 2017)
Food Awareness or Emotional State intake sound (Kleinberger et al., 2023; Koizumi et al., 2011; Zhang and Parker, 2020), speech (Mayumi et al., 2023; Parra et al., 2023; Zhang and Parker, 2020) wrist trajectory (Wang et al., 2025), object trajectory (Khot et al., 2020) emotion state (EDA) (Carroll et al., 2013) food picture (AR) (Narumi et al., 2011) food weight (Mitchell et al., 2015) physical contact (Han and Kang, 2017)
Target Factors and Behaviors Process Behaviors Safe Eating Practices hygiene (Wang and Zhang, 2021) temperature (Singhal et al., 2024)
Analysis of Table 4 reveals key trends in feedback modality distribution. Textual and graphical visual feedback dominate, especially for food choices, intake quantity, and ingestion pace. These channels are favored for their accessibility and clarity.
Symbolic feedback appears less frequently but with growing diversity. LED indicators are used across hydration, eating pace, and safety contexts. Symbolic auditory feedback—such as chewing sounds or metaphorical cues—emerges as a lightweight and engaging method for promoting awareness. Tactile symbolic forms, like thermochromic or pneumatic responses, remain niche but offer ambient and intuitive signaling.
Auditory feedback via speakers and alarms spans nearly all behavior types, while haptic feedback, mostly motor-based, is concentrated in emotional regulation and pace control. More advanced haptics (e.g., piezoelectric, pneumatic) remain underutilized. Experimental modalities like olfactory and physiological feedback appear in a few behaviorally rich cases, offering unique multisensory pathways.
Overall, the table highlights the dominance of visual modalities but also points to symbolic and sensory alternatives as promising design directions. Gaps in environmental, emotional, and hygiene-related feedback suggest potential for broader and more creative feedback strategies in closed-loop ingestion health systems.
Notably, despite the variety of sensing pathways, few systems reuse the same sensors for both perception and intervention. Most existing feedback channels rely heavily on smartphones or static screens, rather than leveraging interactive or perceptual sensors for embodied or context-aware interventions. This reflects a substantial unexploited design space, which we further discuss and explore in the following Chapter 7.
Table 4. Mapping of feedback/intervention Modalities to Target Factors and Behaviors in Ingestion Health
Feedback/Intervention Modality
Visual
Textual Graphical Symbolic Auditory Haptic Olfactory Physiological
Food Choices (total 62 pubs) SMS, app/screen message, AR/MR message (Gutiérrez et al., 2018; Fuchs et al., 2019) (total 26 pubs) app/screen graph, AR/MR graph (Ganesh et al., 2014; Gutiérrez et al., 2018; Fuchs et al., 2019) LED (de Vries et al., 2020; Matsui et al., 2018; Ganesh et al., 2014; Kim et al., 2016c; Ritschel et al., 2018; Triantafyllidis et al., 2019) microphone speaker (Kadomura et al., 2013b; Cai et al., 2024; Kadomura et al., 2014; Robinson et al., 2021; Kadomura et al., 2013a; Zhao et al., 2021; De Carolis et al., 2024; Liu et al., 2024; Ritschel et al., 2018; Chukwu, 2011; Terzioğlu et al., 2023; Triantafyllidis et al., 2019; Agarwal et al., 2024; Luo et al., 2021; Robinson et al., 2020) piezoelectric vibration (Faltaous et al., 2021), Aerodynamic deformation (Pels et al., 2014), friction (Cai et al., 2024) Scented liquid (Cai et al., 2024)
Intake Timing SMS (Bi et al., 2016; Varun et al., 2023), app/screen message (Liang, 2024; Faltaous et al., 2021; Ye et al., 2016) app/screen graph (Liang, 2024; Faltaous et al., 2021) pneumatic membrane (Xie et al., 2023), LED (Fortmann et al., 2014) microphone speaker (Xie et al., 2023), alarm (Varun et al., 2023) piezoelectric vibration (Faltaous et al., 2021), motor vibration (Varun et al., 2023; Ye et al., 2016) LED (Lessel et al., 2016; Ravindran et al., 2022)
Environment microphone speaker (Faltaous et al., 2021)
Context Factors Activity Around Ingestion SMS (Alexander et al., 2017)
Intake Quantity SMS (Poddar et al., 2024; Kreutzer et al., 2015; Moses et al., 2023; Pankajavalli et al., 2017; Wijanarko et al., 2019), app/screen message (Faltaous et al., 2021; Chaudhry et al., 2016; Chiu et al., 2009; Alshurafa et al., 2015; Maramis et al., 2014; Varun et al., 2023; Tommy et al., 2017), LED message (Ravindran et al., 2022) app/screen graph (Faltaous et al., 2021; Chaudhry et al., 2016; Scisco et al., 2011; Maramis et al., 2014; Varun et al., 2023; Chiu et al., 2009; Ko et al., 2007; Wijanarko et al., 2019; Tommy et al., 2017), projector (Yıldız and Coşkun, 2019) pneumatic membrane (Xie et al., 2023), thermo -chromic (Kaner et al., 2018), LED (Bobin et al., 2018; Lessel et al., 2016; Ravindran et al., 2022), living-media (Zhou et al., 2021) microphone speaker (Xie et al., 2023; Farooq et al., 2017), alarm (Varun et al., 2023; Tommy et al., 2017), symbolic auditory (Zhou et al., 2021) piezoelectric vibration (Faltaous et al., 2021), motor vibration (Li and Bailenson, 2018; Chiu et al., 2009; Varun et al., 2023), pneumatic Deformation (Chen et al., 2022) essential oil (Li and Bailenson, 2018)
Ingestion Pace and Oral Motor Behavior SMS (Lee et al., 2019; Mendi et al., 2013; Krishna et al., 2025; Moses et al., 2023), app/screen message (Bi et al., 2016; Kim and Bae, 2018b; Maramis et al., 2014; Kamachi et al., 2023; Turner-McGrievy et al., 2017; Hori et al., 2023; Lo et al., 2007; Ioakimidis et al., 2009) app/screen graph (Nakamura et al., 2023; Scisco et al., 2011; Kim and Bae, 2018b; Maramis et al., 2014; Turner-McGrievy et al., 2017; Hori et al., 2023; Lo et al., 2007), AR/MR graph (Kim and Bae, 2018a; Sugita et al., 2018) LED (de Vries et al., 2020; Kim et al., 2016b, a; Hermans et al., 2017; Kim et al., 2016c) chewing sound feedback (Kleinberger et al., 2023; Chen and Yen, 2022), microphone speaker (Ioakimidis et al., 2009; Sugita et al., 2018; Mendi et al., 2013; Koizumi et al., 2011; Kadomura et al., 2013b), physical symbolic feedback (Mitchell et al., 2015) piezoelectric vibration (Faltaous et al., 2021; Zhang et al., 2019), motor vibration (Kim et al., 2016b, a; Kim and Bae, 2018b; Kamachi et al., 2023; Hermans et al., 2017) Intraoral electrical stimulation (Biyani et al., 2019; Zheng et al., 2025)
Food Awareness or Emotional State SMS(Carroll et al., 2013), app/screen message (Mayumi et al., 2023; Sun et al., 2020; Parra et al., 2023; Wang et al., 2025) app/screen graph(Sun et al., 2020; Parra et al., 2023; Zhang and Parker, 2020), AR/MR graph (Narumi et al., 2011) LED (Han and Kang, 2017; Khot et al., 2020) microphone speaker (Kadomura et al., 2013b; Koizumi et al., 2011; Han and Kang, 2017; Mayumi et al., 2023; Parra et al., 2023), chewing sound feedback (Kleinberger et al., 2023; Zhang and Parker, 2020), physical symbolic feedback (Mitchell et al., 2015) motor vibration (Khot et al., 2020) odor filter (Narumi et al., 2011)
Target Factors and Behaviors Process Behaviors Safe Eating Practices motor vibration (Wang and Zhang, 2021; Singhal et al., 2024)
6. Effectiveness and User Acceptance of Current Ingestion Health Intervention
In this section, we examine papers that include user studies to assess the effectiveness and user acceptance of their work. To systematically understand their results, we have categorized these papers into four types: behavioral feedback/intervention, educational feedback/intervention, social feedback/intervention, and other specific feedback/intervention. Furthermore, we have divided each type into three paradigms: those that include feedback only, those that include intervention only, and those that include both feedback and intervention. In total, 106 papers are included.
6.1. Behavioral Feedback/Intervention
Behavioral feedback/intervention refers to approaches that primarily focus on altering specific ingestive behaviors of users, such as eating speed, chewing counts, and drinking frequency, ultimately assisting them in achieving healthier eating habits. The technical approaches in this category primarily rely on sensor-based technologies (e.g., IMU, piezoelectric sensors, RFID) and real-time feedback mechanisms (e.g., visual, haptic, and auditory cues).
6.1.1. Behavioral Feedback Only
Behavioral feedback refers to the presentation of real-time or retrospective data that directly reflects a user’s actions, enabling self-awareness and self-regulation without explicit instructions or forced behavioral changes. It objectively mirrors user behavior, allowing individuals to draw their own conclusions and adjust their habits autonomously. Systems that provide behavioral feedback only — such as continuous glucose monitoring (Liang, 2024) and smart hydration bracelets (Fortmann et al., 2014) — offer unobtrusive data visualization without explicit instructions, relying on users’ self-reflection to drive behavioral changes. Similarly, ambient logging devices (Tang et al., 2015) and smart nutrition trackers (Alshurafa et al., 2015) collect and display consumption patterns through minimal interfaces, facilitating autonomous behavior adjustment. A total of 20 papers are included in this category.
Feedback effectiveness is evidenced by measurable behavioral changes. For instance, real-time glucose feedback (Liang, 2024) reduced average glucose levels in 50% of participants and prompted adjustments in meal timing, aligning with the ”action” stage of the transtheoretical model. Bite-rate feedback (Kim et al., 2016b; Scisco et al., 2011) led to a 57.7% reduction in energy intake among fast eaters, although the effects varied based on baseline consumption habits. Isomorphic feedback, such as the waist inflation from FatBelt (Pels et al., 2014) and ambient displays (Kaner et al., 2018), utilized physical or visual cues to reduce calorie intake and increase water consumption by 20–30%, respectively. Notably, the modality of feedback matters: younger users (ages 20s) preferred vibration feedback, while older adults (ages 50–70) favored light-based feedback (Kim et al., 2016b). Additionally, manual input requirements, such as the time-interval logging in the SAL logger (Tang et al., 2015), can reduce adherence, particularly among older adults.
The success of behavioral feedback presumably hinges on balancing autonomy (minimal manual input) and personalization (age-appropriate feedback and aesthetic adaptability). Devices that integrate seamlessly into daily routines typically outperform those that require active engagement. Behavioral feedback demonstrates consistent patterns in user acceptance and effectiveness across multiple studies. For instance, wrist-worn devices, such as the WaterJewel bracelet (Fortmann et al., 2014) and the IDEA wristband (Lee et al., 2019), achieve a long-term usage willingness of 70–83% due to their non-intrusive design and plug-and-play functionality. In contrast, neck-worn (Fortmann et al., 2014) or head-mounted sensors face lower acceptance rates due to discomfort or social stigma (Lee et al., 2019). Additionally, aesthetic customization, such as WaterJewel’s design catering to both male and female users (Fortmann et al., 2014) and the GROW bottle’s thermo-chromic display (Kaner et al., 2018), further enhances acceptance by aligning with personal preferences.
6.1.2. Behavioral Intervention Only
Behavioral interventions actively modify ingestive behaviors through direct mechanisms, such as physical constraints or real-time prompts. Physical constraint systems, like the pneumatic shape-changing spoon (Chen et al., 2022), mechanically regulate portion sizes and chewing rates by modifying the tangible interface. Digital guidance platforms, such as portion estimation interfaces (Chaudhry et al., 2016), provide prescriptive dietary recommendations via mobile apps and interactive displays. More immersive interventions, including the sensory interactive table (de Vries et al., 2020) and the smart tray system (Jung et al., 2017), combine environmental sensing with real-time behavioral prompts to actively reshape dining interactions. Unlike passive feedback that merely presents data, these interventions enforce immediate behavioral changes by altering how users interact with food. A total of 15 papers are included in this category.
Behavioral intervention systems actively reshape user actions through physical, computational, or gamified means. The mandatory approach proves particularly effective for users needing structured support, as it bypasses reliance on self-motivation to drive habitual change. Interventions that physically alter user behavior demonstrate particularly strong efficacy. For instance, the Spoon’s shape-changing mechanism reduced bite size by 13.7–16.1% and total intake by 4.4–4.6% without affecting appetite (Chen et al., 2022). Additionally, intraoral closed-loop systems increased mastication force by 26% through real-time electrical stimulation (Biyani et al., 2019) and improved swallowing duration via neuromuscular triggers (Zheng et al., 2025). Structured guidance approaches have also proven effective; for example, the DIMA app enhanced portion estimation accuracy across various literacy levels (Chaudhry et al., 2016), and computer-tailored interventions achieved a twofold greater reduction in fat intake compared to improvements in physical activity through enforced rules (Vandelanotte et al., 2005). Furthermore, WaterCoaster increased drinking frequency from 2–3 to 4–6 times daily by combining personalized hydration goals with virtual rewards (Lessel et al., 2016).
Regarding user acceptance of behavioral interventions, physical manipulation systems like the SSpoon (Chen et al., 2022) achieved high acceptance by subtly altering ingestive behavior without disrupting appetite. Structured guidance approaches, such as computer-tailored interventions, revealed user segmentation; simultaneous interventions were more effective for non-compliant users, while sequential approaches suited compliant individuals (Vandelanotte et al., 2005). Contextually embedded systems exhibited the widest variation in acceptance. For instance, WaterCoaster’s gamified approach engaged 70.6% users (Lessel et al., 2016), whereas the complex gamification of the Sensory Interactive Table likely led to user fatigue (de Vries et al., 2020).
6.1.3. Behavioral Feedback with Intervention
Behavioral feedback with intervention combines the features of real-time data representation and mandatory guidance. It typically takes the form of comprehensive health system applications, such as the smart home-based IoT nutritional intake monitoring system (Faltaous et al., 2021), which provides unobtrusive health intervention advice through a smartphone application along with symbolic feedback mechanisms. This dual approach simultaneously offers users visualized insights while enforcing behavioral modifications through structured guidance. A total of 7 papers are included in this category.
Integrated systems that combine behavioral feedback with active interventions demonstrate enhanced effectiveness and user acceptance when they balance automation with personalization. For instance, a smart home-based nutritional monitoring system achieved over 50% habit adjustment through goal-aligned mobile reminders, although privacy concerns were noted by 66.6% users (Faltaous et al., 2021). Similarly, wristband-based eating speed systems reduced bite size with a detection accuracy of 96.7%, while maintaining high acceptance through haptic or visual feedback (Kim et al., 2016a). In clinical populations, the Parkinson’s smart spoon improved stability (X/Y-axis deviation ¡15% under simulated tremors) while integrating diagnostic alerts; however, its niche design limits broader adoption (Singhal et al., 2024). Low-cost solutions, such as receipt-based nutrition tracking, proved feasible for promoting dietary awareness, particularly for monitoring calcium intake, but relied heavily on user compliance (Mankoff et al., 2002).
6.2. Educational Feedback/Intervention
Educational feedback/intervention refers to feedback or interventions that have educational significance or involve educational content. This includes food and recipe recommendations, nutrition education, educational gaming, and guidance for users to adopt healthy eating habits. Ultimately, these approaches aim to enhance food literacy among users through educational approach.
6.2.1. Educational Feedback Only
Educational feedback focuses exclusively on providing knowledge-based guidance to enhance learning outcomes, without direct instructional intervention prompts. It prioritizes pedagogical value over real-time behavioral modification and delivers structured information (e.g., nutritional facts, eating patterns) to cultivate cognitive awareness and self-reflection, empowering users to make informed dietary decisions. For example, a puzzle game can be utilized to enhance children’s food literacy, encouraging self-driven exploration of healthy eating patterns without explicit behavioral directives (Cai et al., 2024). Additionally, an acoustic sensor can detect ingestion and deliver real-time nutritional insights via a mobile application, thereby raising awareness of everyday dietary habits through unobtrusive monitoring (Bi et al., 2016). This type of feedback emphasizes knowledge dissemination over actual modification, leveraging playful interaction (Cai et al., 2024) or ambient data capture (Bi et al., 2016) to implicitly influence ingestive behaviors. As such, it positions itself as an educational tool for long-term nutrition self-regulation. A total of 11 papers are included in this category.
Educational feedback demonstrates significant potential in improving food literacy and dietary behaviors, particularly when leveraging engaging, multi-sensory interfaces. For instance, multi-sensory puzzle games increased children’s food literacy by 59.16% through tangible interactions that fostered cognitive connections with food (Cai et al., 2024). Similarly, mixed reality (MR) food labels reduced beverage energy content by 34% and sugar content by 28% by making nutritional information more accessible and interactive (Fuchs et al., 2019). Speech-based journaling applications like FoodScrap (Luo et al., 2021) encouraged mindful eating by reducing data entry burdens, with users reporting increased self-reflection on their ingestive behaviors. In clinical populations, the SaltSwitch application (Eyles et al., 2017) successfully reduced salt purchases among patients with cardiovascular disease by providing real-time, barcode-based traffic light nutrition labels.
Educational feedback excels when it reduces cognitive load (e.g., speech input in FoodScrap (Luo et al., 2021) compared to manual logging) and leverages intuitive interfaces (e.g., mixed reality labels versus traditional charts (Fuchs et al., 2019)). It achieves higher user acceptance by prioritizing intuitive interactions and minimizing user burden.
6.2.2. Educational Intervention Only
Educational interventions emphasize the delivery of structured knowledge and instructional guidance through direct conversational interactions, without incorporating data-driven feedback mechanisms. Their design aligns with didactic principles, leveraging conversational interfaces to scaffold learning. This type of intervention is typically exemplified by dialogue-based chatbots, such as Foodbot (Prasetyo et al., 2020), and humanoid social robots (Robinson et al., 2021), which provide goal-oriented educational content (e.g., nutrition advice and dietary management) through scripted or AI-generated conversations. A total of 19 papers are included in this category.
Social robots and embodied agents have proven particularly effective in delivering interventions, increasing intervention effectiveness by 23% compared to digital interfaces. Clinic-based humanoid robots, for instance, have shown moderate effectiveness in reducing alcohol intake while maintaining high usability (Robinson et al., 2021; Mayumi et al., 2023). Similarly, social robots outperformed chatbots in promoting healthier eating intentions among children, leveraging their physical presence to enhance persuasiveness (De Carolis et al., 2024). Interactive tableware, such as FunEat’s gamified plate (Zhao et al., 2021), improved children’s food choices through projected animations, although parents noted potential distractions during meals. Other ubiquitous computing approaches, like sensor-embedded forks, achieved 89% child engagement by replacing abstract lessons with playful interactions (Kadomura et al., 2014).
Personalized educational dietary interventions demonstrate strong effectiveness when incorporating adaptive algorithms. The SmartDiet system (Hsiao and Chang, 2010) generates nutritionally balanced meal plans within seconds while maintaining high user acceptance through interactive optimization. Linear optimization frameworks have achieved less than 1.2% calorie recommendation errors by dynamically adjusting to user deviations (Arefeen et al., 2022). Intelligent nutrition services further enhance adoption by combining image recognition with flexible service selection, although ecosystem factors did not significantly impact perceived educational-targeted behavioral control (Hsiao et al., 2011). Therefore, future systems should balance guidance intensity (Liu et al., 2024) with user autonomy (Ritschel et al., 2018), while leveraging AI optimization for sustainable personalization (Arefeen et al., 2022; Hsiao and Chang, 2010).
6.2.3. Educational Feedback with Intervention
Educational feedback and intervention combine knowledge-based data insights with structured guidance to create a comprehensive learning support system. This approach delivers real-time educational feedback (e.g., nutrition analytics) while providing direct instructional interventions (e.g., personalized study plans and goal-setting prompts) to reinforce desired behaviors through a standardized framework (Kadomura et al., 2013b). Gamification is also widely applied in such scenarios (Nakamura et al., 2023). This dual mechanism ensures that learners not only understand their performance metrics but also receive actionable, curriculum-aligned strategies to bridge knowledge-practice gaps, distinguishing it from standalone feedback or intervention approaches. Related studies demonstrate a convergence of real-time monitoring and pedagogical guidance in dietary feedback and interventions. These implementations collectively operate within the ”observe-educate-act” framework through embedded sensors and algorithmic personalization, although they vary in modality (tangible vs. conversational) and intervention immediacy (real-time vs. retrospective). A total of 13 papers are included in this category.
Integrated systems that combine real-time feedback with structured educational interventions achieve behavioral change through multimodal engagement, expert-guided personalization, and interactive reinforcement. Multimodal engagement proves particularly effective in pediatric contexts. For instance, systems like EducaTableware’s sound-emitting tableware (Kadomura et al., 2013b) and FoodWorks’ AR meal augmentations (Ganesh et al., 2014) increased children’s willingness to try new foods by 45-60% while reducing mealtime distractions. Meanwhile, eat2pic’s color-based food visualization significantly slowed eating speed (meal time/intervals) (Nakamura et al., 2023), although effectiveness varied with users’ artistic interests. Nutrifit’s sensor-tracked meals improved the consumption of nutrient-dense foods by 3-8% through real-time weight monitoring and app feedback (Krishna et al., 2025). Expert-guided systems also demonstrate strong clinical validity. Therefore, if such educational systems combine both embedded sensing and adaptive guidance, they will maximize user acceptance and effective outcomes.
Regarding user acceptance of this combined educational approach, the OCARIoT platform reduced childhood obesity prevalence by 75.5% through IoT-collected nutrition and activity data, along with coordination from healthcare professionals, achieving over 80% user acceptance (Bastida et al., 2023). Web-based health monitoring platforms maintained 95% daily user engagement through real-time expert feedback, although 85% of nutritionists preferred weekly reviews over daily ones (Achilleos et al., 2017). Additionally, non-invasive bioelectric devices like ”Keep and Eat” demonstrated 98.45% accuracy (±1.55% error) in body composition measurements, enabling precise dietary recommendations (Balbin et al., 2020).
6.3. Social Feedback/Intervention
Social feedback/intervention refers to feedback or interventions that involve more than one person, incorporating social support or the eating environment. This type of feedback/intervention leverages group dynamics and shared experiences to reinforce closed-loop paradigms through social norms. They often incorporate interactive elements, such as family meal tracking or community network challenges, to foster collective accountability. By embedding feedback and interventions in natural social contexts, they tap into intrinsic motivations for conformity and belonging that individual ingestion-related approaches cannot address.
6.3.1. Social Feedback Only
Social feedback creates shared accountability through real-time group comparisons (e.g., displaying community members’ vegetable intake rates) or peer endorsements. It leverages social norms and observational learning—showing how similar individuals achieve balanced diets to motivate self-adjustment without prescriptive rules. This type of feedback makes healthy choices feel collectively validated rather than individually enforced. A total of 2 papers are included in this category.
HealthyStadium (Eigen et al., 2018) and Crumbs (Epstein et al., 2016) utilize social feedback—through meal photo rankings and shared challenges—to encourage healthier eating habits more effectively than solo tracking. Both systems demonstrate how passive peer interaction fosters dietary mindfulness and self-adjustment. Their designs intentionally avoid direct intervention, instead relying on social dynamics such as normative comparison and gentle competition to organically shift ingestive behaviors. Both systems demonstrate the effectiveness of social feedback in promoting healthier ingestive habits through lightweight, non-intrusive engagement rather than active intervention. Critically, this lightweight, feedback-driven approach has proven to be more acceptable than traditional interventions, as users tend to resist rigid rules but embrace socially reinforced habits.
Users in HealthyStadium exhibited higher engagement by striving for better health ratings, with participants consciously opting for healthier meals due to peer evaluations. One user noted, “When I was wondering whether to have a delicious meal or a healthy meal, I chose a healthy one” (Eigen et al., 2018). Meanwhile, Crumbs leveraged daily challenges and social interaction, with 72% of users enjoying the challenges and 46% expressing intent to continue using the platform. Additionally, the social group reported motivational effects, such as ”not wanting to fall behind others” (Epstein et al., 2016).
6.3.2. Social Intervention Only
Social intervention for healthy ingestion actively leverages interpersonal dynamics to reshape dietary behaviors through actual engagement. This type of intervention employs real-time social triggers—such as parent-child tactile feedback—that directly mediate ingestive actions within relational contexts. By embedding behavioral guidance within social interactions (e.g., family meals and shared environments), these interventions transform abstract nutritional awareness into contextually anchored behavior change. A total of 4 papers are included in this category.
Child-focused systems like CHIBO (Xie et al., 2023) and interactive tableware (Joi et al., 2016) utilize real-time sensory cues to improve vegetable consumption during parent-child feeding. School-based programs combine environmental changes with personalized feedback, and parental involvement has been shown to boost effectiveness (Haerens et al., 2007). For older groups, auto-generated nutritional postcards link eating patterns with emotional associations (Sun et al., 2020), offering a subtle behavioral nudge.
The interactive tableware system (Joi et al., 2016) successfully encouraged vegetable consumption in children aged 3–4, with observed behavioral shifts, such as a 4-year-old transitioning from being spoon-fed to self-feeding. Additionally, it improved parent-child interaction through gamified learning. Similarly, a middle-school healthy eating intervention demonstrated gender-dependent effects, notably reducing fat intake in girls—particularly when parents were involved—though it had a limited impact on fruit and water intake as well as soft drink consumption (Haerens et al., 2007). Meanwhile, the auto-generated postcard system facilitated self-reflection, leading to expressed intent to modify behaviors (Sun et al., 2020).
A key strength of these interventions is their embedded social context: family dynamics (Joi et al., 2016), peer environments (Haerens et al., 2007), and personal nostalgia (Sun et al., 2020) all serve as external motivators that extend beyond individual willpower, highlighting the importance of the ”social” element. However, effectiveness varied by target group and intervention design—young children responded well to playful tableware (Joi et al., 2016), while adolescents exhibited gender-specific dietary changes (Haerens et al., 2007). In contrast, adults engaged more with emotionally resonant self-reflection tools (Sun et al., 2020).
6.3.3. Social Feedback with Intervention
Social feedback with intervention uniquely hybridizes data-driven insights with social dynamics to synergistically promote healthy ingestive behaviors. These approaches integrate real-time nutritional tracking and visualization with socially embedded prompts—such as dietary monitoring applications for parents (Alloghani et al., 2016) or office hydration competitions (Chiu et al., 2009)—creating a dual mechanism for behavior change. This ensures that targeted nudges are both empirically grounded and socially reinforced. A total of 2 papers are included in this category.
Regarding the outcomes of this approach, the mobile health monitoring application successfully enhanced user trust and acceptance by combining QR-code food tracking, geo-location, and wearable sensors. Users prioritized ease of use and security as key determinants of perceived usefulness (Alloghani et al., 2016). Similarly, Playful Bottle leveraged social reminders alongside system-based hydration games, demonstrating that social prompts were more effective than automated reminders. This approach not only increased water intake volume but also improved regularity and response speed among office workers (Chiu et al., 2009). Notably, social elements consistently outperformed system-only interventions. Users of Playful Bottle responded faster and more reliably to peer reminders than to impersonal alerts (Chiu et al., 2009), echoing findings from the obesity app study that indicated trust and usability are heightened by socially embedded design choices (Alloghani et al., 2016).
6.4. Other Specific Feedback/Intervention
Other specific approaches not covered above are categorized into three key themes: Perceptual Augmentation (i.e., enhancing sensory engagement with food), Mindful and Psychological (i.e., fostering awareness and emotional connections to ingestion), and Smart Monitoring (i.e., automated tracking systems). These frameworks illustrate a range of technological strategies for promoting changes in ingestive behavior.
6.4.1. Perceptual Augmentation
Perceptual augmentation actively modifies ingestion experiences by technologically enhancing sensory feedback during consumption. These approaches employ real-time cross-modal stimulation—such as altering chewing sounds to influence flavor perception (Kleinberger et al., 2023) or using augmented reality (AR) to visually enhance food texture and taste (Koizumi et al., 2011; Narumi et al., 2011)—to directly reshape how food is sensorially experienced. By dynamically manipulating auditory, visual, or haptic cues during the act of ingestion, such interventions create immersive sensory modifications rather than relying on traditional feedback or interventions. A total of 3 papers are included in this category.
Emerging perceptual augmentation technologies demonstrate how non-invasive sensory enhancement can effectively alter ingestive behaviors and perceptions. Auditory feedback systems leverage cross-modal effects to enhance flavor perception—significantly improving texture ratings such as crispiness and reducing hunger (Kleinberger et al., 2023; Koizumi et al., 2011). Meanwhile, Metacookie+ utilizes AR-driven visual and olfactory cues to manipulate taste perception, achieving a 79.3% success rate in altering familiar flavors, although unfamiliar tastes, such as mushroom, saw lower recognition at 42.9% (Narumi et al., 2011).
6.4.2. Mindful & Psychological
Mindful and psychological approaches reshape ingestive behaviors by targeting cognitive-emotional processes and fostering real-time awareness during consumption. These interventions employ evidence-based techniques, such as future-self visualization to motivate long-term behavior change (Rho et al., 2017), emotion-aware wearables to detect and interrupt emotional eating (Carroll et al., 2013), and conversational agents to guide mindful food preparation (Parra et al., 2023). These strategies directly address the psychological roots of maladaptive eating patterns. Ultimately, these interventions cultivate mindful eating by transforming abstract nutritional knowledge into embodied psychological awareness—where every bite becomes an opportunity to nourish both body and mind, and where present-moment consciousness bridges our current selves with the healthier future selves we aspire to become. A total of 4 papers are included in this category.
Mindful and psychological approaches demonstrate strong potential in promoting healthier ingestive behaviors. FutureSelf effectively motivated users by forecasting weight outcomes, with participants reporting increased goal awareness and behavioral conscientiousness (Rho et al., 2017). EmoTree’s Just-In-Time interventions improved emotional eating awareness by 37.5% and received an averaged rating of 3 out of 5 for effectiveness in stress reduction, although its impact on preventing eating was limited, with an averaged rating of 2.2 out of 5 (Carroll et al., 2013). These interventions succeed by leveraging introspection and emotional engagement but require refinement in adaptive personalization and long-term habit formation (Carroll et al., 2013; Rho et al., 2017). Future work should aim to balance immediate support with sustained behavioral reinforcement.
6.4.3. Smart Monitoring
Smart monitoring leverages wearable technologies and ambient sensors—from alcohol-detecting ice cubes (Matsui et al., 2018) to eyeglass-mounted detectors (Bedri et al., 2022) and wrist-worn motion trackers (Ye et al., 2016)—to autonomously capture ingestive behaviors (e.g., drinking and chewing) and environmental interactions (e.g., food selection (Volkova et al., 2016)). This approach transforms passive observation into actionable behavioral datasets. A total of 4 papers are included in this category.
Key strengths in this category include reduced cognitive load (via automated detection) and enhanced self-awareness, particularly regarding snack consumption (Bedri et al., 2022; Ye et al., 2016). However, challenges remain, such as false positives—evidenced by 50% of users receiving 6 or more unnecessary alerts in FitNibble (Bedri et al., 2022)—and issues related to device wearability. Smart monitoring systems have demonstrated effectiveness in improving dietary tracking compliance. FitNibble reduced missed logs by 19.6% and lowered journaling difficulty (Bedri et al., 2022). Similarly, the wrist-worn Pebble system received positive usability scores (2.7 on a scale from -3 to 3), particularly benefiting irregular eaters through visual logging prompts (Ye et al., 2016).
7. Discussion
We synthesize insights from the reviewed systems to reflect on broader limitations and opportunities in the design of behavioral sensing and feedback/intervention in ingestion health. By revisiting the conceptual and technical landscape through the lens of system completeness, behavioral scope, and interdisciplinary integration, we offer a summary analysis that responds to our fourth research question. The discussion is organized around three key challenges: (1) the difficulty of accurately measuring and intervening in real-world ingestive behavior, (2) the disconnect between sensing-focused and intervention-driven research efforts, and (3) the methodological and epistemological gap between HCI and clinical health domains. For each, we outline design directions aimed at informing future development of more effective and ecologically embedded closed-loop systems.
7.1. Challenges in Measuring and Intervening in Real-World Ingestive Behavior
Despite recent advances in sensing technologies and intervention strategies, significant challenges remain in accurately measuring and applying ingestion behavior interventions in real-world settings. These reflections are informed by our previous analysis of intervention effectiveness and user acceptance (Chapter  6), where we observed overall evaluation methods and user feedback in ecological validity, evaluation consistency, and behavioral coverage. In particular, the absence of robust and context-sensitive evaluation frameworks continues to hinder the translation of promising closed-loop sensing–intervention systems into everyday practice. These challenges can be broadly categorized into three areas:
7.1.1. Validity and Consistency of Ingestive Behavior Measurement Indicators
As observed in our analysis of feedback and intervention effectiveness (Chapter 6), the lack of consistency in how ingestive behaviors are defined and evaluated poses a major barrier to assessing and comparing intervention outcomes. Although many studies focus on similar behavioral constructs, the measurement indicators they employ vary widely, resulting in inconsistencies across interpretation, validation, and generalization. Traditional instruments including questionnaires and scales are prevalent in the field of ingestion health intervention, such as the Three-Factor Eating Questionnaire (TFEQ) (Stunkard and Messick, 1985), which assesses cognitive restraint, disinhibition and hunger, and the Eating Behavior Inventory (EBI) (O’Neil et al., 1979), which targets weight-control behaviors. Despite their widespread use, their applicability across diverse populations and contexts has still been questioned, remaining non-universal; for emerging eating behavior interventions that rely on sensing and computational methods, the challenge is further compounded by diverse evaluation metrics and the lack of standardized assessment frameworks. Even when interventions address the same behavioral target—such as eating speed or intake pace—the metrics used to quantify outcomes are highly heterogeneous. To date, medical literature offers no standardized clinical definition for many of these behavioral constructs. For example, although rapid eating is commonly linked to adverse health outcomes, there is no standardized definition or measurement approach; current systems rely on varied metrics such as chewing rate, swallowing frequency, bite intervals, or total meal duration to approximate eating speed. This diversity highlights not only methodological flexibility but also the absence of a shared ground truth in computationally mediated behavior evaluation. Developing valid, standardized, and context-aware behavioral metrics is therefore a critical direction for ensuring scientific rigor and comparability in ingestion health research.
7.1.2. Laboratory-Based Research vs. Real-World Ingestive Behaviors
A substantial portion of ingestive behavior research has been conducted in controlled laboratory environments, which, though offering experimental rigor, often fail to reflect the complexity of real-world contexts. Lab protocols typically exclude environmental, social, and emotional factors that influence natural eating, leading to potential over- or underestimation of intervention effectiveness when deployed outside the lab (De Castro, 2000; Bell et al., 2020; Hiraguchi et al., 2023).
For instance, bite and chewing detection systems are often validated in constrained conditions—such as fixed posture, single food types, and minimized distraction—within laboratory settings [e.g., (Scisco et al., 2011; Kim et al., 2016b; Hori et al., 2023; Zhang et al., 2019)]. However, in real-world environments, behaviors are shaped by multitasking, social interaction, and contextual variability, which can significantly impact system performance and user adherence (Bell et al., 2020; Hiraguchi et al., 2023; Jansen et al., 2022). Moreover, some systems using head-mounted cameras, chest-worn sensors, or bulky wearables often perform well in short-term lab trials, but can feel intrusive or uncomfortable in daily life, potentially disrupting natural behavior and lowering compliance (Bedri et al., 2022; Bell et al., 2020). These real-world deployment issues, such as behavioral reactivity and long-term disengagement, are frequently underestimated in lab-based evaluations (Jansen et al., 2022). Bridging this gap calls for validation in situation that accounts for behavioral variability, contextual complexity, and sustained user acceptance (Hiraguchi et al., 2023).
7.1.3. Potential Adverse Effects of Intervention Deployments
While designed to promote healthier behaviors, eating interventions may produce unintended negative effects. A key concern is ”food noise”, which means persistent and intrusive thoughts about eating triggered by constant monitoring. It can elevate anxiety and contribute to disordered eating patterns (Hayashi et al., 2023). Meanwhile, interventions requiring user interaction during meals, such as smart utensils or app prompts, may also disrupt social interactions or conflict with cultural norms, especially in shared settings (Robinson et al., 2020; Kadomura et al., 2014). Moreover, long-term use of intrusive devices like head-mounted sensors or chewing monitors has been linked to discomfort and user dropout (Bedri et al., 2022). These risks underscore the need for balancing behavioral influence with psychological comfort and contextual fit.
7.2. Gaps between Sensing-Based and Intervention-Based Studies on Ingestive Behavior
Recent reviews on sensor-based research in ingestive behavior [e.g., (Hiraguchi et al., 2023; Bell et al., 2020)] have highlighted that while the target behaviors addressed in sensing studies overlap with those identified in intervention-focused research, significant gaps remain between the two approaches. These gaps primarily stem from differences in modality alignment and the application of behavior change strategies.
7.2.1. Modality Misalignment between Sensing and Intervention Phases
Sensing studies have made significant strides in utilizing a wide array of modalities, including wearable devices, ambient sensors, and smartphone-integrated tools, to detect ingestive behaviors with increasing granularity and accuracy [e.g., (Bedri et al., 2022; Zhang et al., 2020)]. These tools offer real-time monitoring of behaviors like food intake, bite frequency, and meal duration. In contrast, interventions rely on a narrower set of modalities, mostly using visual cues like text prompts, graphical feedback, or auditory signals delivered via mobile devices. Although some studies [e.g., (Kim et al., 2016b)] have used wearable devices for both sensing and delivering interventions, such integrated applications remain relatively scarce, as shown in Chapter 5.3. This disconnect between detection and intervention modalities creates a lack of continuity in the intervention process, limiting the potential for closed-loop systems where detection and feedback are fully integrated. For instance, while detection technologies capture behavioral data, the feedback often remains confined to basic prompts, which may not effectively engage users or encourage sustained behavior change.
7.2.2. Disconnect Between Detection-Centric Studies and Actual Intervention Implementation
A second significant gap exists in the large body of research only focus on improving detection accuracy ,but without progressing to actionable interventions. Many sensor-based studies mention behavior change as a long-term goal but fall short of translating their findings into real-world applications. This disconnection is largely due to the differing complexities of sensing and intervention research. Detection-centric studies typically focus on technical performance metrics, such as the accuracy of behavior detection and system usability, which are easier to evaluate. On contrary, intervention studies must deal with a broader range of challenges, including incorporating behavior change theories, designing meaningful user interactions, and establishing long-term engagement. The higher threshold for evaluating interventions has contributed to a bottleneck, where many studies stop after validating detection methods, while fewer studies advance to actual intervention implementation. As a result, there is an imbalance in the literature, with a higher volume of studies dedicated to detection accuracy, and fewer addressing the complexity of behavior modification and sustained interventions.
7.3. Human-Computer Interaction vs. Medical Perspectives on Ingestive Behavior
Although both Human-Computer Interaction (HCI) and medical research aim to improve ingestive behaviors, such as dietary patterns and meal composition, they differ in their assessment frameworks and metrics. HCI typically focuses on real-time, sensor-based interventions and user experience, whereas medical research emphasizes clinical assessments and physiological measures. These differences pose challenges in aligning the methodologies and measuring the effectiveness of intervention. Additionally, while both fields acknowledge the importance of subjective experiences and physiological states, their integration into behavior metrics differs. HCI research tends to incorporate subjective feedback via interactive devices, while medical research relies more heavily on objective physiological data. This gap highlights the limitations of current HCI technologies in capturing the complex physiological aspects of ingestive behavior.
7.3.1. Divergent Assessment Frameworks
From a medical perspective, healthy ingestive behaviors are often linked to specific health conditions, such as cardiovascular disease, type 2 diabetes, obesity, and cancer. Medical practitioners use comprehensive nutritional assessments that combine food history, biochemical data, medical tests, and clinical procedures (Kesari and Noel, 2022). Tools like the Food Frequency Questionnaire (FFQ) (Cade et al., 2004), Three-Factor Eating Questionnaire (TFEQ) (Karmali et al., 2013), and Adult Eating Behavior Questionnaire (AEBQ) (Hunot et al., 2016) offer a validated, multi-domain approach for assessing ingestive behaviors and their relationship to health outcomes.
In contrast, HCI research often relies on isolated validation methods, lacking the cross-validation mechanisms inherent in medical assessments. While HCI interventions focus on sensor-based technologies and user behavior tracking, they do not integrate the multi-domain assessments seen in medical protocols. This difference limits HCI-based interventions in capturing the complex interactions between physiological states and behavioral patterns. HCI tools primarily focus on real-time data acquisition (e.g., food intake, eating speed), but they often overlook broader health contexts, such as digestive symptoms or stress levels. This results in an incomplete understanding of the long-term impact of HCI interventions on health outcomes.
7.3.2. Limited Perception of Subjective Experience and Physiological States in Target Behavior Metrics
Medical assessments integrate subjective experiences (e.g., emotional eating, satiety cues) with physiological states, offering a comprehensive evaluation of ingestive behaviors essential for disease diagnosis (Kesari and Noel, 2022). These tools benefit from (1) long-term validation, (2) structured clinical environments that enhance patient compliance, and (3) professional interpretation by healthcare specialists. These assessments are particularly effective in documenting longitudinal changes in variables such as pain, function, and psychosocial distress.
In contrast, early HCI research relied on questionnaires that (1) had limited long-term validation, (2) faced low accuracy and compliance due to experimental constraints, and (3) lacked professional interpretation. While HCI interventions track behaviors like chewing or swallowing, they fail to capture subjective experiences (e.g., satiety or emotional states), essential for understanding ingestive behaviors comprehensively. This gap highlights a key limitation: sensors can tell us how we eat, but they do not address why we eat or how we feel during eating, which are crucial aspects for understanding ingestive behaviors in the context of health outcomes.
Thus, while HCI tools focus on quantifiable metrics (e.g., food intake or eating speed), they overlook the subjective aspects emphasized by medical assessments. To address this gap, HCI interventions should integrate both subjective and physiological data to offer a more holistic understanding of ingestive behaviors, particularly in real-world contexts.
7.4. Design Recommendations
7.4.1. Optimizing Experimental Design and Intervention Implementation
To bridge the gap between laboratory-based research and real-world interventions, clear and standardized measurement metrics should be established, drawing from validated constructs in the existing literature. Combining qualitative and quantitative methods provides a comprehensive view of intervention outcomes, helping reduce bias and ensuring a more accurate evaluation. Additionally, integrating self-reported data with objective physiological measures offers a balanced approach to long-term tracking, avoiding participant fatigue. For example, targeted micro-assessments, triggered by unusual behavior patterns, can provide timely and context-specific insights without overwhelming participants.
When designing experimental protocols, it is essential to consider real-world conditions, user experience, and long-term feasibility. Instead of excluding common unhealthy behaviors, such as continuous eating without proper pauses, interventions should directly address them as part of the behavior change process. This ensures that interventions are more relevant to natural eating contexts, promoting sustainable behavior change.
For long-term engagement, the technical design should prioritize usability and integration with daily life. Embedding sensing capabilities into familiar, unobtrusive devices—such as smartwatches with inertial sensors or bone conduction earphones—can enhance natural interaction and reduce disruption. This approach will ensure continuous monitoring while preserving comfort and user adherence.
Furthermore, ethical safeguards must be incorporated into the intervention design. Participants should be clearly informed that they can withdraw at any time if they feel discomfort or emotional distress, ensuring their psychological well-being is not compromised by the intervention.
Finally, given the inconsistency in measurement tools across studies, future interventions should consider context-sensitive evaluation methods. These should account for cultural differences and ensure interventions align with both natural ingestive behaviors and social norms, improving applicability across diverse populations.
7.4.2. Enhancing Integration and Adaptability in Sensing-Driven Interventions for Ingestion Health
To effectively bridge the gap between sensing and intervention in ingestion health, future interventions should focus on the seamless integration of real-time data from sensing technologies into personalized, actionable, and adaptive interventions. To enhance the continuity of the closed-loop system, interventions should increasingly leverage the same or compatible devices for both sensing and feedback delivery. Such consistency in device use not only improves the intuitiveness and coherence of the intervention process but also fosters user familiarity and acceptance over time. More personalized and actionable intervention requires moving beyond basic detection capabilities, not only by gathering more granular physiological and behavioral data but also by ensuring that feedback mechanisms are tailored to the specific needs, preferences, and circumstances of each user. For instance, interventions must consider factors such as individual eating patterns, meal timing, emotional states, and physiological markers like glucose levels, enabling a level of personalization that ensures more effective engagement and results (Liang, 2024; Reynolds et al., 2019).
Moreover, interventions should be designed with adaptability in mind. Success in behavior change depends not only on tracking short-term ingestive behaviors but also on continually adapting the intervention based on long-term data. This means creating systems that can evolve their feedback strategies in response to behavioral trends and progress over time. Such adaptive interventions maintain their relevance and motivational power, preventing user fatigue and promoting long-term engagement. A crucial aspect of this adaptability involves understanding the dynamic nature of ingestive behaviors. For example, certain behaviors may be triggered by emotional states or specific environmental factors, which require intervention systems that can quickly detect and respond to these variables. This requires the continuous collection and analysis of both physiological and behavioral data, ensuring that interventions are timely and appropriate.
Finally, the long-term success of such systems hinges on the ability to sustain behavior change. The intervention should encourage not just short-term compliance but foster lasting habits by gradually increasing the complexity or frequency of the feedback, thereby encouraging the user to internalize healthier behaviors and develop autonomous control over their eating habits. This continuous, evolving support through adaptive interventions will ensure that users are not only supported in their current behavior change but are also equipped with the tools to sustain these changes in the long run.
7.4.3. Aligning with Medical Goals and Practices
To bridge the gap between HCI-based interventions and medical objectives, personalized interventions should account for the diversity in ingestive behaviors, physiological baselines, and cultural backgrounds among participants (Mirmiran et al., 2021). Incorporating both subjective and physiological data, especially baseline parameters like blood glucose and heart rate variability, will improve the precision and effectiveness of interventions (Liang, 2024; Harous et al., 2017).
Future studies should adopt dynamic intervention models and extend trial durations to better capture long-term effects. Exploring dose-response relationships, including varying intervention frequency, duration, and intensity, will help identify the optimal parameters for sustained health benefits (Pankajavalli et al., 2017; Kim et al., 2016b).
Emerging behavioral and physiological targets such as postprandial heart rate variability, blood glucose response curves, and digestive symptoms (e.g., bloating, nausea) offer valuable insights into metabolic health. Continuous glucose monitoring (CGM) has shown promise in real-time glucose tracking, and integrating it with wearable technologies, including smartwatches and fitness trackers, can offer multi-dimensional feedback on ingestive behavior and metabolic response (Liang, 2024; Bedri et al., 2022; Zhao et al., 2021).
Multimodal sensing approaches, integrating wearable and minimally invasive technologies, could enable more comprehensive tracking of physiological and behavioral signals, enhancing intervention outcomes (Zhang and Parker, 2020). Future sensor fusion techniques, combining CGM with emotion detection (EDA) and activity tracking, will offer a complete view of ingestive behavior and its impact on health (Carroll et al., 2013).
As sensor technology progresses, real-time systems could detect subtle digestive processes, like esophageal peristalsis during swallowing, to assess intervention effects on digestive function (Tirasirichai et al., 2018; Biyani et al., 2019). These advancements will improve our understanding of nutrition-physiology relationships and provide actionable insights for chronic disease prevention (Reynolds et al., 2019; Bi et al., 2016).
8. Conclusion and Limitations
In this review, we systematically analyzed 136 studies focused on behavioral sensing and intervention paradigms in ingestion health, with an emphasis on sensor-based systems and interactive interventions. These studies were categorized into various types based on the modality of feedback and the behavioral targets of the interventions. A key contribution of this review is the development of a behavioral closed-loop framework, which highlights the importance of seamlessly integrating sensing, reasoning, and feedback mechanisms for effective behavior change. This framework offers a detailed structure for the design of adaptive, context-aware interventions, targeting both physiological and behavioral aspects of ingestive health.
However, despite the comprehensive nature of this review, there are several limitations that should be acknowledged. Our keyword search strategy, centered around the term “ingestion health,” may have excluded relevant studies that do not explicitly focus on ingestion but contribute to broader research in behavior modification or health interventions. Moreover, the majority of the studies in HCI domain reviewed were short-term, laboratory-based experiments, which limits the generalizability of our findings to real-world settings. The dynamic nature of real-life eating behaviors requires more longitudinal studies to better understand the long-term effectiveness and challenges of these interventions.
Acknowledgements.
Appendix A RESULT COUNTS OF PAPER VENUES
Table 5. Venues and number of publications found after applying the SLR search queries.
Conference / Journal Venue Number
ACM Conference on Human Factors in Computing Systems (CHI) 16
ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT/UbiComp) 16
ACM Conference on Recommender Systems (RecSys) 1
ACM Interaction Design and Children Conference (IDC) 2
ACM Transactions on Interactive Intelligent Systems (TiiS) 1
ACM Symposium on User Interface Software and Technology (UIST) 1
ACM Designing Interactive Systems Conference (DIS) 4
ACM International Conference on Bioinformatics, Computational Biology, and Health Informatics 1
ACM Southeast Conference 1
ACM/IEEE International Conference on Human-Robot Interaction (HRI) 2
IEEE Sensors 2
IEEE International Conference on Robot and Human Interactive Communication (RO-MAN) 2
IEEE International Workshop on Advances in Sensors and Interfaces (IWASI) 1
IEEE International Conference on Innovations in Electrical, Electronics, Instrumentation and Media Technology (ICEEIMT) 1
IEEE/ACM Conference on Connected Health: Applications, Systems and Engineering Technologies (CHASE) 1
IEEE North Karnataka Subsection Flagship International Conference (NKCon) 1
IEEE International Conference on Intelligent Computing and Control Systems (ICICCS) 1
IEEE International Conference on Computer Science and Engineering (UBMK) 1
IEEE International Conference on Digital Information Processing and Communications (ICDIPC) 1
IEEE International Wireless Communications and Mobile Computing Conference (IWCMC) 1
IEEE International Symposium on Computer-Based Medical Systems (CBMS) 1
IEEE International Conference on Contemporary Computing and Communications (InC4) 1
IEEE International Conference on Robotics and Biomimetics (ROBIO) 1
IEEE International Conference on Industrial Engineering and Engineering Management (IEEM) 1
IEEE World Forum on Internet of Things (WF-IoT) 1
IEEE/EMBS International Conference on Neural Engineering (NER) 1
IEEE International Conference on Information, Intelligence, Systems and Applications (IISA) 1
IEEE Healthcare Innovation Conference (HIC) 1
IEEE Transactions on Neural Systems and Rehabilitation Engineering 1
IEEE International Conference on Convergence Information Technology (ICCIT) 1
IEEE Journal of Translational Engineering in Health and Medicine 1
IEEE International Conference on Integrated Intelligence and Communication Systems (ICIICS) 1
IEEE Consumer Communications and Networking Conference (CCNC) 1
IEEE International Conference on Virtual Rehabilitation (ICVR) 1
IEEE International Conference on Bioinformatics and Bioengineering (BIBE) 1
IEEE Conference on e-Learning, e-Management and e-Services (IC3e) 1
IEEE International Symposium on Medical Measurements and Applications 1
IEEE EMBS International Conference on Biomedical and Health Informatics (BHI) 1
IEEE International Conference on Disruptive Technologies (ICDT) 1
IEEE International Conference on Multimedia and Expo Workshops 1
IEEE Latin America Transactions 1
International Conference on Multimodal Interaction (ICMI) 1
International Conference on Data Science and Information Technology (DSIT) 1
International Conference on Advances in Computer Entertainment Technology (ACE) 1
International Conference on Mobile and Ubiquitous Multimedia (MUM) 1
International Conference on the Internet of Things (IoT) 1
International Conference on Advanced Visual Interfaces (AVI) 1
International Workshop on Multisensory Approaches to Human-Food Interaction (MHFI) 1
International Conference on Trends in Quantum Computing and Emerging Business Technologies 1
International Conference on Intelligent Data Communication Technologies and Internet of Things (IDCIoT) 1
International Conference on Design Innovation for 3 Cs Compute Communicate Control (ICDI3C) 1
International Conference on Wireless Mobile Communication and Healthcare (MOBIHEALTH) 1
International Conference on Computer Science and Education 1
International Conference on Human-Computer Interaction with Mobile Devices and Services Adjunct 1
International Conference on Computing Engineering and Design (ICCED) 1
International Conference on Computer and Information Technology (CIT) 1
International Conference on Intelligent User Interfaces 1
International Conference on Computing Communication and Networking Technologies (ICCCNT) 2
International Conference on Tangible, Embedded, and Embodied Interaction (TEI) 2
International Joint Conference on Biomedical Engineering Systems and Technologies (BIOSTEC) 1
International Joint Conference on Service Sciences 1
International Workshop on Multimedia Assisted Dietary Management (MADiMa) 1
International Workshops: BioFor, CTMR, RHEUMA, ISCA, MADiMa, SBMI, and QoEM (ICIAP) 1
International Joint Conference on Computer Science and Software Engineering (JCSSE) 1
Annals of Behavioral Medicine 1
Augmented Human International Conference 1
Appetite 1
Behavior Research Methods 1
European Journal of Preventive Cardiology 1
Humaine Association Conference on Affective Computing and Intelligent Interaction 1
Internet Interventions 1
Innovations in Power and Advanced Computing Technologies (i-PACT) 2
Journal of the American Dietetic Association 1
Journal of Dental Research 1
Journal of Medical Systems 1
JMIR mHealth and uHealth 1
Mensch und Computer (MuC) 1
Nordic Conference on Human-Computer Interaction (NordiCHI) 1
Obesity 1
Proceedings of the Behavior Transformation by IoT International Workshop (BTIW) 1
Public Health Nutrition 1
Presence 1
PervasiveHealth 2
SIGGRAPH Asia 2
Smart Health 1
Symposium on Applied Computing 1
Systems and Information Engineering Design Symposium (SIEDS) 1
Studies in Health Technology and Informatics 1
Valuable Visualization of Healthcare Information (VVH) 1
World Congress on Medical Physics and Biomedical Engineering 2
Total 136
Appendix B DISTRIBUTION OF PUBLICATIONS ACROSS TARGET FACTORS AND BEHAVIORS
Table 6. Taxonomy of Target Factors and Behaviors with Involved Publications
Target Factor or Behavior References
Contextual Factors
    Food Choices
       Nutrition Structure Control (Kadomura et al., 2013b; Nakamura et al., 2023; de Vries et al., 2020; Zenun Franco, 2017; Cai et al., 2024; Joi et al., 2016; Kadomura et al., 2014; Haerens et al., 2007; Jung et al., 2017; Kadomura et al., 2013a; Zhao et al., 2021; De Carolis et al., 2024; Ganesh et al., 2014; Butscher et al., 2016; Arefeen et al., 2022; YR et al., 2024; Krishna et al., 2025; Alloghani et al., 2016; Hsiao and Chang, 2010; Gaikwad et al., 2024; Hsiao et al., 2011; Dragoni et al., 2017; Ludwig et al., 2011; Achilleos et al., 2017; Kalantarian et al., 2014; Park and Min, 2007; Bastida et al., 2023; Triantafyllidis et al., 2019; Qu and Yu, 2014; Karkar et al., 2018; Badawi et al., 2012a; Agarwal et al., 2024; Sun et al., 2020; Casas et al., 2018; Gutiérrez et al., 2018; Kwon et al., 2016; Bedri et al., 2022; Badawi et al., 2012b; Volkova et al., 2016; Waltner et al., 2015; Epstein et al., 2016; Mankoff et al., 2002; Huang et al., 2010; Vazquez et al., 2015; Akter et al., 2021; Jung et al., 2007; Lee et al., 2007)
       Calorie Target Control (Zenun Franco, 2017; Balbin et al., 2020; Burke et al., 2011; Pels et al., 2014; Rho et al., 2017; Kumbhare et al., 2023; YR et al., 2024; Mishra et al., 2024; Harous et al., 2017; Ludwig et al., 2011; Kanjalkar et al., 2024; Cordeiro et al., 2015; Kim et al., 2016c; Tirasirichai et al., 2018; Jung et al., 2007; Lee et al., 2007)
       Macronutrient and Micronutrient Control
                Carbohydrate Intake (Arefeen et al., 2024)
                Sugar Intake (Liang, 2024)
                Fat Intake (Vandelanotte et al., 2005; Haerens et al., 2007; Burke et al., 2011; Chukwu, 2011)
                Salt Intake (Eyles et al., 2017; Kim et al., 2016c)
       Limiting Specific Food Categories
                Snacks (Bi et al., 2016; Liu et al., 2024; Maramis et al., 2014; Bedri et al., 2022; Robinson et al., 2020)
                Unhealthy Drinks (Matsui et al., 2018; Robinson et al., 2021; Faltaous et al., 2021; Haerens et al., 2007; Ritschel et al., 2018; Fuchs et al., 2019)
       Unspecified Food Choices (Prasetyo et al., 2020; Abisha and Rajalakshmy, 2017; Huang et al., 2018; Terzioğlu et al., 2023; Fernandez et al., 2013; Eigen et al., 2018; Luo et al., 2021; Nedungadi et al., 2018)
    Intake Timing (Xie et al., 2023; Bi et al., 2016; Fortmann et al., 2014; Varun et al., 2023; Lessel et al., 2016; Ravindran et al., 2022; Liang, 2024; Faltaous et al., 2021; Ye et al., 2016)
    Environment (Faltaous et al., 2021)
    Activity Around Ingestion (Alexander et al., 2017)
Process Behaviors
    Intake Quantity
       Solid Food (Xie et al., 2023; Chaudhry et al., 2016; Chen et al., 2022; Alshurafa et al., 2015; Scisco et al., 2011; Maramis et al., 2014; Farooq et al., 2017; Moses et al., 2023; Li and Bailenson, 2018)
       Liquid (Faltaous et al., 2021; Alshurafa et al., 2015; Kaner et al., 2018; Poddar et al., 2024; Kreutzer et al., 2015; Varun et al., 2023; Bobin et al., 2018; Lessel et al., 2016; Yıldız and Coşkun, 2019; Chiu et al., 2009; Ko et al., 2007; Pankajavalli et al., 2017; Wijanarko et al., 2019; Tommy et al., 2017; Ravindran et al., 2022; Zhou et al., 2021)
    Ingestion Pace and Oral Motor Behavior
       Chewing and Swallowing Behaviors (Kadomura et al., 2013b; Kleinberger et al., 2023; Koizumi et al., 2011; Bi et al., 2016; Scisco et al., 2011; Maramis et al., 2014; Chen and Yen, 2022; Sugita et al., 2018; Turner-McGrievy et al., 2017; Hori et al., 2023; Kim et al., 2016b; Kamachi et al., 2023; Biyani et al., 2019; Zheng et al., 2025)
       Bite Intervals and Size (Nakamura et al., 2023; de Vries et al., 2020; Lee et al., 2019; Kim et al., 2016a; Kim and Bae, 2018b, a; Mendi et al., 2013; Hermans et al., 2017; Chen et al., 2022)
       Total Meal Duration (Moses et al., 2023; Zhang et al., 2019; Kim et al., 2016c; Ioakimidis et al., 2009; Krishna et al., 2025; Lo et al., 2007; Mitchell et al., 2015)
    Food Awareness or Emotional State (Kleinberger et al., 2023; Koizumi et al., 2011; Parra et al., 2023; Wang et al., 2025; Khot et al., 2020; Kadomura et al., 2013b; Han and Kang, 2017; Sun et al., 2020; Zhang and Parker, 2020; Tang et al., 2015; Mayumi et al., 2023; Carroll et al., 2013; Mitchell et al., 2015; Narumi et al., 2011)
    Safe Eating Practices (Wang and Zhang, 2021; Singhal et al., 2024)
Appendix C DISTRIBUTION OF PUBLICATIONS IN SENSING MODALITY
Table 7. Taxonomy of Sensing Modality with Involved Publications
Sensing Modality References
Human-based
    Visual
       Gaze (Karkar et al., 2018; Khot et al., 2020)
       Facial expression (Mayumi et al., 2023; Kim and Bae, 2018a; Triantafyllidis et al., 2019; Robinson et al., 2020)
       Body (Xie et al., 2023; Mayumi et al., 2023; Zhang and Parker, 2020)
    Auditory
       Speech (Prasetyo et al., 2020; Mayumi et al., 2023; Robinson et al., 2020; Arefeen et al., 2022; Luo et al., 2021; Parra et al., 2023; Kamachi et al., 2023)
       Sound (Zhang and Parker, 2020; Kleinberger et al., 2023; Koizumi et al., 2011; Bi et al., 2016; Chen and Yen, 2022; Kalantarian et al., 2014; Kamachi et al., 2023; Maramis et al., 2014)
    Haptic (Harous et al., 2017; Robinson et al., 2020; Cai et al., 2024; Fortmann et al., 2014; Alshurafa et al., 2015; Kim et al., 2016b; Zheng et al., 2025; Biyani et al., 2019; Ludwig et al., 2011; Tang et al., 2015; Abisha and Rajalakshmy, 2017; Farooq et al., 2017; Bastida et al., 2023; Nedungadi et al., 2018)
    Motion
       Gesture (Kadomura et al., 2013a; Mishra et al., 2024; Harous et al., 2017; Fuchs et al., 2019; Singhal et al., 2024; Li and Bailenson, 2018; Faltaous et al., 2021; Nakamura et al., 2023; Lee et al., 2019; Kadomura et al., 2014; Joi et al., 2016; Scisco et al., 2011; Kim et al., 2016a; Kim and Bae, 2018b; Mendi et al., 2013; Kumbhare et al., 2023; Gaikwad et al., 2024; Dragoni et al., 2017; Bobin et al., 2018; Alexander et al., 2017; Chukwu, 2011; Badawi et al., 2012a; Bedri et al., 2022; Wang et al., 2025; Hermans et al., 2017; Ye et al., 2016; Turner-McGrievy et al., 2017; Zhang et al., 2019; Khot et al., 2020; Chiu et al., 2009; Tirasirichai et al., 2018; Tommy et al., 2017; Zhou et al., 2021; Vazquez et al., 2015; Jung et al., 2007; Maramis et al., 2014; Bastida et al., 2023)
       Chewing and Micro-head Motion (Biyani et al., 2019; Kadomura et al., 2013b, a; Hori et al., 2023; Han and Kang, 2017)
    Physiological
       Autonomic (Lee et al., 2019; Sugita et al., 2018; Kim et al., 2016b; Park and Min, 2007; Carroll et al., 2013; Jung et al., 2007; Nedungadi et al., 2018)
       Metabolic (Chukwu, 2011; Harous et al., 2017; Liang, 2024; Balbin et al., 2020; Arefeen et al., 2024; Nedungadi et al., 2018)
       Cardiac (Huang et al., 2018; Park and Min, 2007; Ludwig et al., 2011; Jung et al., 2007; Varun et al., 2023; Badawi et al., 2012b; Nedungadi et al., 2018)
    Others (Zenun Franco, 2017; Chaudhry et al., 2016; Prasetyo et al., 2020; Burke et al., 2011; Vandelanotte et al., 2005; Haerens et al., 2007; Robinson et al., 2021; Pels et al., 2014; Rho et al., 2017; Butscher et al., 2016; Mishra et al., 2024; Harous et al., 2017; Huang et al., 2018; Kanjalkar et al., 2024; Sun et al., 2020; Moses et al., 2023; Arefeen et al., 2022; Ludwig et al., 2011; Jung et al., 2007; Bastida et al., 2023; Nedungadi et al., 2018)
Environment-based
    Visual (Kadomura et al., 2013a; Ganesh et al., 2014; Mishra et al., 2024; Kanjalkar et al., 2024; Fuchs et al., 2019; Cai et al., 2024; Arefeen et al., 2022; Faltaous et al., 2021; Nakamura et al., 2023; Kadomura et al., 2014; Achilleos et al., 2017; Chiu et al., 2009; Vazquez et al., 2015; Wang and Zhang, 2021; YR et al., 2024; Alloghani et al., 2016; Hsiao et al., 2011; Fernandez et al., 2013; Gutiérrez et al., 2018; Eigen et al., 2018; Narumi et al., 2011; Cordeiro et al., 2015; Volkova et al., 2016; Waltner et al., 2015; Epstein et al., 2016; Eyles et al., 2017; Mankoff et al., 2002; Akter et al., 2021; Kwon et al., 2016)
    Acoustic (Faltaous et al., 2021; Tommy et al., 2017; Liu et al., 2024; Poddar et al., 2024; Wijanarko et al., 2019; Ravindran et al., 2022; Varun et al., 2023)
    Force (Ritschel et al., 2018; Mitchell et al., 2015; Joi et al., 2016; Hermans et al., 2017; Jung et al., 2017; Zhao et al., 2021; Krishna et al., 2025; Lessel et al., 2016; Lo et al., 2007; Kim et al., 2016c; Ioakimidis et al., 2009; de Vries et al., 2020; Maramis et al., 2014)
    Spectral (Ritschel et al., 2018; Singhal et al., 2024; Bedri et al., 2022; Tommy et al., 2017; Wijanarko et al., 2019; Matsui et al., 2018; Bastida et al., 2023; Varun et al., 2023; Badawi et al., 2012b; Kwon et al., 2016)
    Electrical
       Conductivity (Bobin et al., 2018; Tommy et al., 2017; Kaner et al., 2018; Kreutzer et al., 2015; Ko et al., 2007; Pankajavalli et al., 2017; Bastida et al., 2023; Varun et al., 2023; Badawi et al., 2012b; Kwon et al., 2016)
       Magnetic (Vazquez et al., 2015; Han and Kang, 2017; Huang et al., 2010; Jung et al., 2017; Yıldız and Coşkun, 2019)
Appendix D DISTRIBUTION OF PUBLICATIONS IN FEEDBACK/INTERVENTION MODALITY
Table 8. Taxonomy of Feedback/Intervention Modality with Involved Publications
Feedback/Intervention Modality References
Visual
    Textual (Liang, 2024; Faltaous et al., 2021; Zenun Franco, 2017; Chaudhry et al., 2016; Balbin et al., 2020; Lee et al., 2019; Prasetyo et al., 2020; Joi et al., 2016; Burke et al., 2011; Vandelanotte et al., 2005; Alshurafa et al., 2015; Bi et al., 2016; Haerens et al., 2007; Jung et al., 2017; De Carolis et al., 2024; Kim and Bae, 2018b; Rho et al., 2017; Butscher et al., 2016; Mendi et al., 2013; Abisha and Rajalakshmy, 2017; Carroll et al., 2013; Kumbhare et al., 2023; Arefeen et al., 2022; YR et al., 2024; Mishra et al., 2024; Poddar et al., 2024; Krishna et al., 2025; Singhal et al., 2024; Alloghani et al., 2016; Harous et al., 2017; Hsiao and Chang, 2010; Gaikwad et al., 2024; Kreutzer et al., 2015; Hsiao et al., 2011; Huang et al., 2018; Maramis et al., 2014; Dragoni et al., 2017; Varun et al., 2023; Ludwig et al., 2011; Achilleos et al., 2017; Alexander et al., 2017; Kalantarian et al., 2014; Park and Min, 2007; Bastida et al., 2023; Kanjalkar et al., 2024; Chukwu, 2011; Terzioğlu et al., 2023; Fernandez et al., 2013; Qu and Yu, 2014; Badawi et al., 2012a; Arefeen et al., 2024; Agarwal et al., 2024; Sun et al., 2020; Parra et al., 2023; Casas et al., 2018; Gutiérrez et al., 2018; Kwon et al., 2016; Bedri et al., 2022; Fuchs et al., 2019; Eigen et al., 2018; Badawi et al., 2012b; Kamachi et al., 2023; Moses et al., 2023; Ye et al., 2016; Turner-McGrievy et al., 2017; Hori et al., 2023; Lo et al., 2007; Volkova et al., 2016; Waltner et al., 2015; Epstein et al., 2016; Eyles et al., 2017; Robinson et al., 2020; Kim et al., 2016c; Mankoff et al., 2002; Nedungadi et al., 2018; Tirasirichai et al., 2018; Pankajavalli et al., 2017; Wijanarko et al., 2019; Tommy et al., 2017; Ravindran et al., 2022; Ioakimidis et al., 2009; Vazquez et al., 2015; Akter et al., 2021; Jung et al., 2007; Lee et al., 2007)
    Graphical (Cai et al., 2024; Liang, 2024; Faltaous et al., 2021; Nakamura et al., 2023; Chaudhry et al., 2016; Balbin et al., 2020; Joi et al., 2016; Kadomura et al., 2014; Tang et al., 2015; Scisco et al., 2011; Kadomura et al., 2013a; Zhao et al., 2021; Kim and Bae, 2018b; Ganesh et al., 2014; Kumbhare et al., 2023; YR et al., 2024; Mishra et al., 2024; Alloghani et al., 2016; Hsiao et al., 2011; Maramis et al., 2014; Kalantarian et al., 2014; Bastida et al., 2023; Kanjalkar et al., 2024; Agarwal et al., 2024; Sun et al., 2020; Parra et al., 2023; Gutiérrez et al., 2018; Kwon et al., 2016; Yıldız and Coşkun, 2019; Fuchs et al., 2019; Sugita et al., 2018; Turner-McGrievy et al., 2017; Hori et al., 2023; Lo et al., 2007; Cordeiro et al., 2015; Chiu et al., 2009; Ko et al., 2007; Tirasirichai et al., 2018; Wijanarko et al., 2019; Zhang and Parker, 2020; Lee et al., 2007)
    Symbolic (de Vries et al., 2020; Matsui et al., 2018; Fortmann et al., 2014; Tang et al., 2015; Kaner et al., 2018; Kim et al., 2016b; Han and Kang, 2017; Kim et al., 2016a; Kim and Bae, 2018a; Ganesh et al., 2014; Ritschel et al., 2018; Bobin et al., 2018; Lessel et al., 2016; Triantafyllidis et al., 2019; Mitchell et al., 2015; Hermans et al., 2017; Khot et al., 2020; Chiu et al., 2009; Kim et al., 2016c; Ravindran et al., 2022; Zhou et al., 2021)
Auditory (Xie et al., 2023; Kadomura et al., 2013b; Kleinberger et al., 2023; Cai et al., 2024; Kadomura et al., 2014; Koizumi et al., 2011; Robinson et al., 2021; Han and Kang, 2017; Kadomura et al., 2013a; Mayumi et al., 2023; Zhao et al., 2021; De Carolis et al., 2024; Liu et al., 2024; Ritschel et al., 2018; Mendi et al., 2013; Singhal et al., 2024; Varun et al., 2023; Chukwu, 2011; Terzioğlu et al., 2023; Triantafyllidis et al., 2019; Agarwal et al., 2024; Parra et al., 2023; Luo et al., 2021; Chen and Yen, 2022; Farooq et al., 2017; Sugita et al., 2018; Zhang and Parker, 2020; Robinson et al., 2020; Huang et al., 2010; Tommy et al., 2017; Zhou et al., 2021; Ioakimidis et al., 2009)
Haptic
    Vibration (Faltaous et al., 2021; Wang and Zhang, 2021; Kim et al., 2016b, a; Chiu et al., 2009; Kim and Bae, 2018b; Varun et al., 2023; Kamachi et al., 2023; Ye et al., 2016; Hermans et al., 2017)
    Deformation (Chen et al., 2022; Pels et al., 2014; Cai et al., 2024; Li and Bailenson, 2018; Zhang et al., 2019; Khot et al., 2020; Singhal et al., 2024)
Olfactory (Cai et al., 2024; Narumi et al., 2011; Li and Bailenson, 2018)
Physiological
    Electrodermal (Zheng et al., 2025; Biyani et al., 2019)

### v4

@@
-skipabove=1em,skipbelow=1em
+A Review of Behavioral Closed-Loop Paradigm from Sensing to Intervention for Ingestion Health
 Abstract.
-Figure 1. Overall framework of ingestion health sensing and intervention
+Figure 1. Paradigm of ingestion health from sensing to intervention
 Ingestive behavior plays a critical role in health, yet many existing interventions remain limited to static guidance or manual self-tracking.
-With the increasing integration of sensors and perceptual computing, recent systems have begun to support closed-loop interventions that dynamically sense user behavior and provide feedback during or around ingestion episodes.
-In this survey, we review 136 studies that leverage sensor-enabled or interaction-mediated approaches to influence eating behavior.
-We propose a behavioral closed-loop paradigm comprising three core components: target behaviors, sensing modalities, and feedback strategies.
+With the increasing integration of sensors, context-aware computing, and perceptual computing, recent systems have begun to support closed-loop interventions that dynamically sense user behavior and provide feedback during or around ingestion episodes.
+In this survey, we review 136 studies that leverage sensor-enabled or interaction-mediated approaches to influence ingestive behavior.
+We propose a behavioral closed-loop paradigm rooted in context-aware computing and inspired by HCI behavior change frameworks, comprising four components: target behaviors, sensing modalities, reasoning and intervention strategies.
 A taxonomy of sensing and intervention modalities is presented, organized along human- and environment-based dimensions.
@@
 1. Introduction
-Ingestion health has been shown to significantly influence quality of life, longevity, and overall well-being (Jovanov et al., 2014; Mesas et al., 2012).
-Given its critical importance, it is essential to develop effective strategies that integrate both sensing and intervention for ingestive behaviors.
-Traditional approaches to ingestion monitoring and intervention, such as food diaries (Thompson and Subar, 2017), self-report questionnaires (Shim et al., 2014), and periodic counseling sessions (Greaves et al., 2011), are heavily dependent on user self-discipline and engagement.
-Moreover, these methods are typically static, offering one-time or episodic interventions without continuous feedback mechanisms (Raber et al., 2021; Takemoto et al., 2018).
+Ingestion health plays a critical role in shaping health outcomes, influencing quality of life, longevity, and overall well-being (Jovanov et al., 2014; Mesas et al., 2012).
+Despite its importance, current ingestion health interventions often rely on static tools such as food diaries (Thompson and Subar, 2017), self-report questionnaires (Shim et al., 2014), and periodic counseling sessions (Greaves et al., 2011), which place a heavy burden on users and provide limited behavioral feedback (Raber et al., 2021; Takemoto et al., 2018).
 As a result, they often fail to dynamically support individuals in developing sustainable and healthy ingestive behaviors over time.
-With the advancement of sensing technologies, new opportunities have emerged for low-burden and continuous sensing of ingestive behaviors.
-These technologies not only support the development of systems capable of capturing rich behavioral data with minimal user effort, but also enable more personalized and adaptive feedback and intervention strategies (Bell et al., 2020).
-Consequently, there is a growing body of research that leverages sensing and interactive technologies to bridge the gap between data collection and real-time intervention in ingestion health.
-This shift represents a critical evolution from static, user-driven methods toward context-aware, behavior-informed support systems.
-Building on these advances, research in behavioral sensing and intervention has expanded considerably in recent years, alongside growing interest in ingestion health.
-Several recent reviews have addressed related domains, which can be broadly categorized into three research streams: (1) dietary intake and nutrition assessment, (2) wearable ingestion detection technologies, and (3) behavior change interventions.
-Dietary Intake Nutrition Assessment
-Traditional dietary assessment methods face challenges in both accuracy and user burden, prompting researchers to explore sensor-based technologies as a more objective alternative.
-Skinner et al.
-(2020) proposed a framework that integrates physical, chemical, and biosensors to objectively capture what, how, and when individuals eat.
-Wang et al.
-(2022) reviewed real-time sensor systems for nutritional support, while emphasizing deployment challenges.
-While these advances improve the reliability of intake measurement objectivity, they remain largely confined to assessment purposes and lack extensions toward adaptive behavioral interventions in ingestion health.
-Wearable Ingestion Detection Technologies
-Wearable-based ingestion detection has been extensively reviewed from different perspectives.
-Bell et al.
-(2020) surveyed automatic intake detection studies in free-living settings, highlighting sensor modalities and validation metrics.
-He et al.
-(2020) categorized detection systems by sensor type, while Hiraguchi et al.
-(2023) provided a comprehensive survey covering both eating and drinking behaviors.
-Beyond general ingestion detection, specialized methods for chewing detection have been explored.
-Wei et al.
-(2022) and Mortazavi and Gutierrez-Osuna (2023) reviewed wearable-based chewing sensing techniques and signal processing approaches.
-Despite these efforts, existing studies and reviews consistently emphasize detection performance, without advancing toward behavioral feedback or adaptive intervention frameworks.
-Behavior Change Intervention
-Previous studies have reviewed a variety of digital intervention strategies aimed at promoting healthy eating behaviors.
-Zhang et al.
-(2020) proposed a generative design framework for technology-mediated eating interventions, while Pan et al.
-(2021) surveyed digital tools, including mobile apps and wearable devices, designed to support healthy eating in occupational settings.
-However, these reviews primarily focus on organizing design considerations and intervention strategies, often relying on traditional education or standalone gamification without incorporating sensing technologies or perceptual computing.
-As a result, they typically lack integration with behavioral sensing or interactive feedback mechanisms.
-In summary, while many existing reviews emphasize isolated stages such as intake detection, dietary assessment, or single-instance feedback, few focus on the complete behavioral closed-loop process from sensing to intervention.
-This fragmentation highlights the urgent need for a systematic review that frames ingestion health interventions within a behavioral closed-loop paradigm, grounded in perceptual computing and spanning the full pipeline from sensing to feedback and intervention.
-To address these gaps, we present a structured review guided by the following research questions (RQs), aimed at evaluating how ingestion health interventions are conceptualized and implemented within behavioral closed-loop systems:
+The emergence of ubiquitous computing has long envisioned systems that seamlessly integrate into everyday life, sensing and responding to context to enhance human activity (Weiser, 1991).
+Advances in sensing technologies have further enabled low-burden, continuous monitoring of ingestive behaviors.
+In particular, multimodal sensor fusion has made unobtrusive ingestion monitoring feasible (Bell et al., 2020), allowing for the collection of rich behavioral data with minimal user effort.
+Context-aware computing is a core capability of ubiquitous computing (Abowd and Mynatt, 2000), and it advocates modular architectures in which continuous sensing informs higher-level reasoning to enable adaptive interventions.
+Dey (2001) defined context-aware systems as those that ”use context to provide relevant services to the user, where relevancy depends on the user’s task”.
+Accordingly, sensing is not merely about acquiring raw data, but about capturing context to enable situation inference.
+Intervention, then, corresponds to the application component that, as Abowd and Mynatt (2000) emphasized, makes decisions and takes actions based on inferred context.
+These principles support personalized, adaptive interventions grounded in real-time behavioral context (Bell et al., 2020), marking a shift from static, user-driven methods to context-aware, behavior-informed systems.
+Studies of ingestion health related to the aforementioned context have been systematically reviewed, with prior reviews primarily focusing on three key aspects: dietary assessment, wearable ingestion detection, and behavior change intervention.
+i.
+Dietary Assessment: Skinner et al.
+(2020) and Wang et al.
+(2022) reviewed sensor-based frameworks for improving dietary assessment accuracy.
+However, their scope remains limited to intake monitoring, without addressing how such data can inform adaptive interventions.
+ii.
+Wearable Ingestion Detection: Reviews by Bell et al.
+(2020) and He et al.
+(2020) systematically analyzed sensor types and detection pipelines, while others focused on chewing-specific sensing (Wei et al., 2022; Mortazavi and Gutierrez-Osuna, 2023).
+These works emphasize detection performance but rarely consider downstream behavior change frameworks.
+iii.
+Behavior Change Intervention: Zhang et al.
+(2020) and Pan et al.
+(2021) reviewed digital tools for eating-related interventions but primarily centered on static strategies such as education or gamification, with limited integration of sensing or context-aware intervention.
+In summary, prior reviews tend to isolate sensing, reasoning, or intervention, without linking them into an integrated behavior change pipeline.
+This highlights the need for a systematic review that explicitly integrates ingestion health interventions and perceptual computing into a cohesive system.
+To address this, we propose a behavioral closed-loop paradigm for ingestion health that encompasses sensing, reasoning, and intervention.
+This paradigm builds on Salber’s Context Toolkit (Salber et al., 1999), inspired by Personal Informatics Framework (Li et al., 2010) and Digital Behavior Change Intervention (Hekler et al., 2016), aiming to bridge the scattered components identified in previous reviews.
+Moreover, it advances from passive context display toward more active and adaptive forms of behavioral scaffolding.
+Building on this conceptual foundation, we conduct a structured review guided by the following research questions (RQs), aimed at evaluating how ingestion health interventions are conceptualized and implemented within behavioral closed-loop systems:
   • RQ1: What behaviors are targeted in the existing literature on ingestion health intervention?
@@
   • We conduct a systematic review of ingestion health interventions that employ behavioral sensing and interactive systems, and develop a taxonomy of key components across different stages of the behavioral closed-loop paradigm.
-  • We propose a behavioral closed-loop framework tailored to ingestive behavior interventions, and review the evaluation of effectiveness and user acceptance in current literature.
+  • We propose a behavioral closed-loop paradigm, grounded in context-aware computing and informed by HCI behavior change frameworks, to guide ingestive behavior interventions, and review the evaluation of effectiveness and user acceptance in current literature.
   • We identify critical gaps between current ingestion health interventions and real-world behavioral complexities, and provide design insights for better bridging behavioral targets, sensing, reasoning, and intervention stages toward effective closed-loop management.
@@
 2.1. Paper Retrieval
-Given that most relevant publications place greater emphasis on the feedback or intervention aspect, we also include studies that do not cover the full pipeline from sensing to intervention.
-After careful discussion, we decided to include all publications in which sensor-based interactive devices or perceptual computing methods are involved at any stage within the behavioral closed-loop process, ranging from sensing to feedback or intervention.
+Given that most relevant publications place greater emphasis on the intervention aspect, we also include studies that do not cover the full pipeline from sensing to intervention.
+After careful discussion, we decided to include all publications in which sensor-based interactive devices or perceptual computing methods are involved at any stage within the behavioral closed-loop process, ranging from sensing to intervention.
 To ensure consistency in inclusion and classification throughout in this paper, we first examined three core terms and finalized their definitions as follows:
   • Ingestive Behavior: The process by which an individual ingests nutrition from food or liquid intake, excluding the intake of medications or medical nutrition supplements, and comprising appetitive behaviors, consummatory behaviors, and post-ingestive physiological responses;
-  • Health Intervention: A system-generated feedback mechanism, whether active or passive, that aims to alter the user’s behavior with the goal of improving health from a medical perspective;
+  • Health Intervention: A system-generated mechanism aimed at altering user behavior for health improvement, categorized as passive interventions — providing feedback to foster self-awareness, and active interventions — delivering structured guidance or direct mechanisms to proactively guide behavior.
   • Perceptual Computing: Technologies that enable machines to perceive, compute, and reason with data from real-world sensory inputs.
-Specifically, Benoit and Tracy (2008) defined ingestive behavior as ”a complex process controlled by multiple physiological and psychological mechanisms”, and divided it to two general categories, including ”behaviors involved in seeking, locating, and acquiring foods” and ”behaviors associated with the act of ingestion itself, including chewing and swallowing”.
+While the paradigm is conceptually grounded in context-aware computing, we adopt ”perceptual computing” as a core operational term.
+This term provides a more implementation-oriented perspective on behavioral data processing in closed-loop systems building on context-aware principles.
+It focuses on gaps in prior reviews where perception and intervention were often treated as disconnected components.
+Specifically, Benoit and Tracy (2008) defined ingestive behavior as ”a complex process controlled by multiple physiological and psychological mechanisms”, and divided it into two general categories, including ”behaviors involved in seeking, locating, and acquiring foods” and ”behaviors associated with the act of ingestion itself, including chewing and swallowing”.
 We refer to these two categories as ”appetitive behaviors” and ”consummatory behaviors”, and additionally added ”post-ingestive behaviors” to cover physiological feedback mechanisms and related behaviors after ingestive behavior.
-Moreover, Mendel and Wu (2010) described perceptual computing as a sensor-based method “associated with machinery for computing and reasoning with perceptions and data,” while Poole (2013) noted that “health interventions leverage technology to deliver timely support, monitoring, and guidance to influence health behaviors and outcomes.” These definitions align closely with the conceptual scope of our study.
+Moreover, Mendel and Wu (2010) described perceptual computing as a sensor-based method ”associated with machinery for computing and reasoning with perceptions and data”, while Poole (2013) noted that ”health interventions leverage technology to deliver timely support, monitoring, and guidance to influence health behaviors and outcomes”.
+These definitions align closely with the conceptual scope of our study.
 Accordingly, we adopted and refined the above terms to form the foundation of our inclusion criteria and analytical framework.
@@
 To illustrate the outcome of this refinement process, we present the top five high-frequency terms extracted via TF-IDF and Bi-gram analysis (excluding semantically uninformative combinations such as function words or prepositional phrases) as follows:
-{mdframed}
-[backgroundcolor=gray!15] TF-IDF: eating, behavior, feedback, food, intake
-Bi-gram: behavior(al) change, eating rate, eating habits, eating behavior, to improve
+TF-IDF: eating, behavior, feedback, food, intake Bi-gram: behavior(al) change, eating rate, eating habits, eating behavior, to improve
 We selected three primary databases for our literature retrieval: the ACM Digital Library, IEEE Xplore, and Google Scholar.
@@
 Given the variation in keyword syntax and thesaurus constraints across platforms, we tailored the search strings and scoping targets for each database accordingly.
-The final keyword formulations and corresponding search results for each source are summarized below.
+The final keyword formulations are summarized below, with corresponding search details for each source shown in Appendix A.
 All searches were conducted on April 4, 2025, without restricting the publication time interval.
-{mdframed}
-[backgroundcolor=gray!15] Keywords (combined with AND):
-  • eating-related: eat*, ”food”, diet*, ”intake”
-  • intervention-related: interven*, ”feedback”, ”behavior(al) change”, promot*, adjust*, design*, improv*, alter*
-  • technology-related: ”wearable”, ”sensor(s)”, ”device(s)”, ”interactive”
-ACM Digital Library
-  • query target: Title, Keywords, Abstract
-  • filters: No keynote, tutorial, invited talk, panel
-  • inquiry result: 679
-IEEE Explore
-  • query target: Document Title (eating), Abstract (eating, intervention), All Metadata (technology)
-  • filters: Conferences, Journals
-  • inquiry result: 911
-Google Scholar
-  • filters: No books
-  • inquiry result: 100
+Keywords (combined with AND): eating-related: eat*, ”food”, diet*, ”intake” intervention-related: interven*, ”feedback”, ”behavior(al) change”, promot*, adjust*, design*, improv*, alter* technology-related: ”wearable”, ”sensor(s)”, ”device(s)”, ”interactive”
 2.2. Screening
-To further narrow down the set of retrieved articles to those aligned with the defined scope of our review, two authors conducted a structured two-stage screening process in accordance with PRISMA guidelines.
+To further narrow down the set of retrieved articles to those aligned with the defined scope of our review, two authors conducted a structured two-stage screening process with PRISMA guidelines.
 During the initial screening, which involved title and abstract review followed by full-text assessment, each article was evaluated against a predefined set of inclusion and exclusion criteria.
-Articles were labeled as “include”, “exclude”, or “uncertain” and any disagreements were resolved through discussion until consensus was reached.
+Articles were labeled as ”include”, ”exclude”, or ”uncertain” and any disagreements were resolved through discussion until consensus was reached.
 The selection criteria used for this process are outlined below and have been adapted to follow PRISMA recommendations.
@@
 Annual publication count is shown in Figure 3, illustrating a growing research interest in ingestion health in recent years.
-We identified 93 distinct publication venues across all included studies, with full details provided in LABEL:tab:venues_list.
+We identified 93 distinct publication venues across all included studies, with full details provided in Table LABEL:tab:venues_list.
 Among them, the following venues featured two or more included papers: ACM Conference on Human Factors in Computing Systems (CHI, n=16), ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT)/UbiComp (N=16), ACM Interaction Design and Children Conference (IDC, n=2), ACM Designing Interactive Systems Conference (DIS, n=4), IEEE Sensors (n=2), IEEE International Conference on Robot and Human Interactive Communication (RO-MAN, n=2), ACM/IEEE International Conference on Human-Robot Interaction (HRI, n=2), PervasiveHealth (n=2), SIGGRAPH Asia (n=2), International Conference on Tangible, Embedded, and Embodied Interaction (TEI, n=2), Innovations in Power and Advanced Computing Technologies (i-PACT, n=2), International Conference on Computing Communication and Networking Technologies (ICCCNT, n=2), World Congress on Medical Physics and Biomedical Engineering (n=2).
@@
 To facilitate systematic analysis, we created a Google Sheets table with 57 columns.
-The column design and the taxonomy of core features (corresponding to Chapter 4 and Chapter 5) were iteratively developed.
+The column structure was iteratively developed to support the discovery and refinement of the taxonomy of core features (corresponding to Chapter 4 and Chapter 5).
 Two authors first defined preliminary categories during the filtering stage, followed by collaborative refinement of the taxonomy through a full review of the papers.
@@
 3. Result Overview
-Following the review methodology and the research questions outlined above, this section presents an overview derived from the selected studies, including a classification based on Just-in-Time (JIT), and the primary behavioral closed-loop framework that synthesizes the key stages from sensing to intervention in ingestion health.
+Following the review methodology and research questions outlined above, this section presents a two-part overview of papers selected in Section 2.2.
+We first classify the publications based on Just-in-Time (JIT) characteristics, highlighting their temporal and contextual responsiveness.
+Then, we introduce our behavioral closed-loop paradigm grounded in context-aware computing and inspired by HCI behavior change frameworks.
+This paradigm outlines key stages from sensing to intervention and serves as the analytical foundation for the subsequent taxonomy and discussion.
 3.1. Primary Taxonomy based on JIT
-Given the critical role of timing and context-awareness in optimizing the effectiveness of behavioral interventions, as highlighted in previous reviews (Hsu et al., 2025), we classified the included studies based on whether their intervention strategies exhibit JIT characteristics.
-With a primary overview based on JIT, this classification offers a high-level understanding of the integration mechanisms across existing ingestive behavior interventions, encompassing both sensing and feedback/intervention components.
-JIT interventions are defined as approaches that deliver support dynamically at behaviorally or contextually opportune moments, relying on real-time assessments of users’ behavioral, physiological, or environmental states (Nahum-Shani et al., 2018; Sarker et al., 2014).
-Determining the appropriate timing for intervention typically involves identifying periods when users are vulnerable to adverse health outcomes and are physically, cognitively, and socially available to receive support (Choi et al., 2019; Sarker et al., 2014).
-As shown in Figure 4(a), early researches in this domain primarily focused on non-JIT interventions, exhibiting a steady but slow growth over the years.
-Beginning from 2007, studies incorporating JIT intervention strategies began to emerge, and by 2011, the cumulative number of JIT publications surpassed that of non-JIT publications.
-A marked acceleration in the growth of JIT-related work can be observed around 2015–2016, which can be attributed to the rapid advancement of wearable sensing technologies and the increasing availability of real-time behavioral and contextual data.
-The distribution of JIT versus non-JIT studies is further illustrated in Figure 4(b), where JIT interventions account for a substantial majority (73.7%) of the included literature.
-Together, these findings provide a high-level understanding of the temporal trends and overall characteristics of research in the field of ingestion health interventions.
+Building on prior work emphasizing the importance of timing and context-awareness in behavioral interventions (Hsu et al., 2025), we classified the included studies based on whether their intervention strategies exhibit JIT characteristics.
+This classification offers a high-level understanding of the integration mechanisms across existing ingestion health systems, encompassing both sensing and intervention components.
+JIT interventions are defined as approaches that deliver support dynamically at behaviorally or contextually opportune moments, determined by real-time assessments of users’ behavioral, physiological, or environmental states—particularly when individuals are vulnerable and receptive to intervention (Nahum-Shani et al., 2018; Sarker et al., 2014; Choi et al., 2019).
+As illustrated in Figure 4(a), JIT interventions have gained increasing attention since their emergence around 2007, surpassing non-JIT studies by 2011.
+A notable surge occurred after 2015, driven by advances in wearable sensing and the availability of real-time contextual data.
+Overall, JIT interventions account for 73.7% of the reviewed studies, as shown in Figure 4(b), highlighting their central role in contemporary ingestion health intervention design.
 (a) Cumulative publication count
@@
 Figure 4. Cumulative publication count of ingestion health research (a) and publications pie (b).
-3.2. Behavioral Closed-loop Framework
-After providing an overview of the presence of JIT characteristics in the reviewed literature, we shift our focus to systematically examining behavior sensing and intervention paradigms, structured through a behavioral closed-loop framework.
-This framework serves as the analytical foundation for addressing the review questions outlined earlier, providing a coherent structure for understanding the current landscape of ingestive behavior interventions throughout the remainder of the paper.
-In the broader field of human-computer interaction (HCI), ”closed-loop” architectures are defined as systems in which signals from the user’s brain and body are continuously collected, analyzed, and categorized as inputs to an adaptive controller, which then triggers appropriate responses at the interface (Fairclough, 2015).
-Such systems facilitate a continuous, bidirectional dialogue between the user and the technology.
-Furthermore, interaction loops in HCI can be conceptualized as exchanges between mutually interacting agents, where both the user and the system perceive and act upon each other’s states over time (Murray-Smith et al., 2024).
-In the context of ingestive behavior interventions discussed in this paper, we define ”behavioral closed-loop” as dynamic systems or approaches modeled as exchanges between interacting agents.
-In these systems, user ingestive behavior is monitored, interpreted, and utilized to deliver adaptive feedback or interventions, with subsequent behavioral changes tracked to iteratively refine future measures.
-This framework emphasizes the dynamic, responsive, and iterative nature of behavioral closed-loop interventions, capturing the essential stages involved in intervening from perception to ingestive behavior.
-Specifically, the stages in behavioral closed-loop system include identifying target behaviors, sensing behaviorally relevant data, interpreting the observed phenomena through reasoning processes, developing feedback or intervention strategies, and monitoring subsequent behavioral adjustments to close the loop.
-This structure guides the subsequent analysis and provides a detailed understanding of how perceptual computing techniques are applied across different stages of intervention design.
-An overview of the primary behavioral closed-loop paradigm and its staged evolution is illustrated in Figure 5.
-While existing studies often focus on specific stages rather than achieving a fully closed-loop design, we structure the subsequent analysis and synthesis based on the ideal of a complete behavioral closed-loop framework.
+3.2. Behavioral Closed-loop Paradigm
+We propose a behavioral closed-loop paradigm that integrates various components — sensing, reasoning, and intervention — into a cohesive system.
+This paradigm is conceptually grounded in Human-Computer Interaction (HCI), drawing inspiration from context-aware computing and behavior change fields.
+Specifically, context-aware theories provide the theoretical foundation for system components, while HCI behavior change frameworks inform the behavioral logic of the closed-loop paradigm, resulting in an integrated closed-loop system.
+3.2.1. Grounded in Context-Aware Computing
+Context-aware computing emphasizes how interactive systems continuously sense, interpret, and respond to contextual information to enable adaptive, task-relevant interventions.
+Building on Dey’s foundational definition, context is defined as ”any information that can be used to characterize the situation of an entity”, and a context-aware system ”uses context to provide relevant information and/or services to the user, where relevancy depends on the user’s task(Dey, 2001).
+These two definitions align with our paradigm’s components in ingestion health systems: Sensing and Reasoning.
+Additionally, the Context Toolkit proposed by Salber et al.
+(1999) highlights a key principle: the separation of sensing, context interpretation, and application logic into modular components.
+Consequently, we have derived three main components in our paradigm:
+  • Sensing: Raw sensor inputs are first collected as behavioral signals.
+  • Reasoning: Contextual signals are combined and interpreted to infer the user’s situation.
+  • Intervention: Application components make task-relevant decisions and deliver adaptive actions based on interpreted context.
+3.2.2. Inspired from Behavior Change Frameworks
+We then turn to behavioral change frameworks for a process-oriented perspective that clarifies how components can be organized temporally and functionally.
+The Personal Informatics Framework proposed by Li et al.
+(2010) describes a five-stage cyclical process: preparation, collection, integration, reflection, and action.
+This framework inspired us to conceptualize our paradigm as a behavioral loop, mapping collection to Sensing, integration to Reasoning, and the reflection–action sequence to Intervention, framing the paradigm as a loop that enables continuous adjustment.
+Building on this, the Digital Behavior Change Interventions (DBCI) framework (Hekler et al., 2016) provides system-level insights by integrating sensing, tailoring, action, and interaction into a unified architecture for adaptive interventions (Yardley et al., 2016).
+This approach led us to design the components to function not as isolated modules but as an integrated system.
+While our paradigm focuses on three system-level components, it draws inspiration from the five-stage PIF process.
+Some stages, such as preparation, are embedded implicitly within broader sensing or reasoning phases.
+3.2.3. Strengthened with Closed-Loop Concept
+To reinforce the ”loop” structure, we draw from broader ”closed-loop” architectures in HCI.
+According to Fairclough (2015), closed-loop systems continuously monitor user states and feed those signals into adaptive controllers to trigger context-appropriate responses.
+Murray-Smith et al.
+(2024) further conceptualize such systems as mutually interactive agents, where both the user and the system perceive and act on each other’s changing states over time.
+Drawing from these perspectives, we develop our behavioral closed-loop paradigm as a dynamic, sensing-driven system that continuously interprets user behavior and delivers adaptive interventions, enabling iterative feedback over time and context.
+This feedback can arise from real-time behavioral data, user-state inference, or performance evaluation, enabling adaptive recalibration across the loop.
+Instead of a monolithic process, this paradigm comprises recurring stages: identifying target behaviors, sensing relevant data, delivering interventions, and evaluating outcomes to guide future actions.
+When applied across everyday environments, these loops form dense, time- and context-sensitive patterns, realizing the vision of ubiquitous computing through distributed, context-aware interventions.
+Figure 5 illustrates the core structure and evolution of the behavioral closed-loop paradigm, informing our subsequent taxonomy and system-level analysis.
 Figure 5. Behavior closed-loop overview
@@
 Daily ingestive behavior interventions can play an important role in promoting healthier eating and drinking habits.
-The targets of these interventions reflect the key behaviors believed to be critical for promoting healthy eating and improving overall health outcomes.
-Understanding the specific factors and behaviors targeted in the existing intervention literature is crucial, as they not only highlight important aspects of ingestive behaviors relevant to health, but also provide insight into the intervention strategies designed to address them.
-By categorizing and analyzing these targeted behaviors, we can better understand how current interventions aim to modify eating habits, assess the effectiveness of different approaches, and identify gaps in existing methods.
-This chapter aims to explore the targeted behaviors in current ingestive behavior intervention studies, categorized into contextual factors (external to the human body) and process behaviors (individual actions during the ingestion process).
+Identifying the specific factors and behaviors targeted in prior research helps clarify health-relevant aspects of ingestion and associated intervention strategies.
+By analyzing these behaviors, we can evaluate how current interventions modify eating patterns, compare approach effectiveness, and uncover gaps in existing methods.
+This chapter aims to explore the targeted behaviors in current ingestive behavior intervention studies, and classify them into contextual factors (external to the human body) and process behaviors (individual actions during the ingestion process).
 These two broad categories encompass a wide range of interventions designed to optimize various target factors or behaviors for ingestion and promote healthier ingestion habits.
-Figure 3 presents trends in the number of intervention studies over the years which can be categorized into three groups: contextual factors only, process behaviors only, and studies targeting both.
+Figure 3 presents the yearly distribution of intervention studies which can be categorized into three groups: contextual factors only, process behaviors only, and studies targeting both.
 Both categories have garnered significant attention from researchers, with a greater number of interventions focusing on contextual factors.
-Figure 6. Taxonomy of target factors and behaviors with the number of studies involving each target.
-Building on this, Figure 6 displays the taxonomy of these target factors and behaviors, providing a structured overview of the specific intervention targets within each category and the number of studies associated with each.
+Building on this, Figure 6 displays the taxonomy of these target factors and behaviors, providing a structured overview of the specific intervention targets within each category and the number of associated studies.
 This taxonomy organizes the diverse intervention targets into a hierarchical structure, clarifying the relationship between broader categories and more specific target behaviors.
 A summary of these categories and the corresponding publications included in this review is shown in Table 2.
+Figure 6. Taxonomy of target factors and behaviors with the number of studies involving each target.
 4.1. Contextual Factors
 Contextual factors refer to external influences that affect ingestive behaviors, which are usually not limited to the duration of the ingestion process itself.
-These factors, which include food choices, intake timing, the environment in which eating takes place, and activity around ingestion process, can shape an individual’s dietary patterns and habits.
-Since these influences often operate outside the immediate context of the meal, they play a crucial role in determining overall ingestive behaviors and are frequently targeted in interventions aimed at promoting healthier eating habits.
+These include food choices, intake timing, the eating environment, and activities surrounding the ingestion process, all of which can shape an individual’s dietary patterns and habits.
 4.1.1. Food Choices
@@
 Poor food choices are strongly associated with the development of chronic diseases, such as obesity, diabetes, cardiovascular disease, and certain cancers, while healthier selections promote better metabolic health and disease prevention.
-Given the central role of food choices in overall dietary health, many current ingestive behavior interventions prioritize food choices as a primary target.
-In fact, among the reviewed literature, interventions targeting food choices constitute the largest proportion.
+In fact, interventions focusing on food choices represent the largest portion of the reviewed literature.
 These interventions aim to improve dietary quality by modifying nutrient intake, promoting healthier eating habits, and restricting the consumption of unhealthy foods.
-Within these interventions, food choice-related targets are commonly categorized into several specific areas: Nutrition Structure Control (aiming for balanced nutrition and dietary variety), Calorie Control (regulating total energy intake), Macronutrient and Micronutrient Control (focusing on specific nutrients such as carbohydrates, fats, sugars, and salt), Limiting Specific Food Categories (e.g., reducing intake of snacks and sugary beverages), and Unspecified Food Choices (general recommendations for healthier eating without specific nutrient guidelines).
-Among these categories, interventions focusing on nutrition balance and calorie control are particularly prevalent, reflecting their perceived critical role in promoting healthy eating and preventing diet-related chronic diseases.
-Nutrition Structure Control: This target behavior focuses on ensuring balanced nutrient intake and promoting dietary variety.
-Nutrition structure control interventions focus on fostering a balanced intake of various nutrients by encouraging dietary diversity.
-Rather than emphasizing individual food items or nutrients, these interventions encourage individuals to consume a wide range of food groups, including fruits, vegetables, whole grains, proteins, and dairy.
-A significant portion of these interventions specifically target children, with the goal of instilling balanced eating habits early in life.
-Many studies aim to prevent picky ingestive behaviors and encourage the development of healthy, varied eating patterns from a young age.
-By promoting a broader range of food intake, these interventions help address nutrient deficiencies and support long-term health maintenance.
-Calorie Target Control: This target behavior aims to regulate energy intake through portion size estimation or meal planning.
-Calorie control interventions primarily aim to manage overall energy intake to prevent overweight, obesity, and related chronic diseases such as diabetes and cardiovascular conditions.
-These interventions guide individuals in selecting lower-calorie, nutrient-dense foods, such as fruits, vegetables, and lean proteins.
-Strategies often include reducing portion sizes and replacing high-calorie foods with healthier alternatives.
-In the reviewed literature, calorie control is frequently employed as a strategy for weight management and reducing obesity risk, underscoring its central role in ingestion health interventions.
+Within these interventions, food choice-related targets are commonly categorized into several specific areas: Nutrition Structure Control, Calorie Control, Macronutrient and Micronutrient Control, Limiting Specific Food Categories, and Unspecified Food Choices.
+Among these categories, interventions focusing on nutrition balance and calorie control are particularly prevalent, reflecting their perceived critical role in promoting healthy ingestion and preventing diet-related chronic diseases.
+i.
+Nutrition Structure Control: This target behavior aims to promote balanced nutrient intake by encouraging dietary diversity across major food groups, such as fruits, vegetables, whole grains, proteins, and dairy.
+A significant portion of interventions targets children, seeking to cultivate healthy eating habits early in life.
+Additionally, some studies address picky eating behaviors that limit dietary variety.
+By broadening food intake, these interventions help prevent nutrient deficiencies and support long-term health maintenance.
+ii.
+Calorie Target Control: This target behavior focuses on regulating energy intake through portion control, meal planning, and calorie management.
+Interventions guide individuals toward lower-calorie, nutrient-dense food, such as fruits, vegetables, and lean proteins, while encouraging the replacement of high-calorie items with healthier alternatives.
+In the reviewed literature, calorie control is frequently employed as a strategy for weight management and obesity prevention, underscoring its central role in ingestion health interventions.
+iii.
 Macronutrient and Micronutrient Control: This category includes interventions targeting the regulation of specific nutrient intakes, such as carbohydrates, fats, sugars, and salt.
 Many interventions focus on adjusting these components to address health issues like diabetes, hypertension, and cardiovascular diseases.
-For example, regulating carbohydrate intake is often emphasized for individuals at risk of or diagnosed with diabetes, while reducing fat (particularly saturated fat) and sodium intake is associated with cardiovascular risk reduction.
 Current intervention targets include control over specific types of nutrients, such as:
@@
   • Salt Intake — Limiting sodium intake helps prevent hypertension and cardiovascular risks (Jaques et al., 2021).
-Current interventions focus on controlling salt intake to promote healthier eating habits in patients with metabolic syndrome, such as those with cardiovascular disease (Eyles et al., 2017; Kim et al., 2016c).
+Current interventions focus on controlling salt intake in patients with metabolic syndrome, such as those with cardiovascular disease (Eyles et al., 2017; Kim et al., 2016c).
+iv.
 Limiting Specific Food Categories: Some interventions specifically target the reduction or elimination of certain food categories considered unhealthy, such as snacks high in sugar or fat, and sugary beverages like soft drinks and alcohol.
-By limiting the intake of these foods, the aim is to reduce empty calorie consumption, lower the risk of obesity, and improve overall diet quality.
-This approach often complements broader nutrition and calorie-focused strategies by removing key sources of dietary excess.
 Current research focuses on reducing intake of:
@@
   • Unhealthy Drinks — Interventions mainly focus on alcohol intake (Matsui et al., 2018; Robinson et al., 2021) or sugar-sweetened soft drink intake control (Faltaous et al., 2021; Haerens et al., 2007; Ritschel et al., 2018; Fuchs et al., 2019).
+v.
 Unspecific Food Choices: This refers to general recommendations that do not focus on specific nutrients or food categories.
-A smaller subset of interventions addresses food choices in a more general way, providing broad guidance to ”eat healthier” without specifying particular nutrients or food categories.
+A smaller subset of interventions addresses food choices in a more general way, providing guidance to ”eat healthier” without specifying particular nutrients or food categories.
 These interventions typically focus on encouraging awareness about healthy eating principles and promoting better food decision-making habits, sometimes through educational tools (Prasetyo et al., 2020; Fernandez et al., 2013), reminders (Abisha and Rajalakshmy, 2017; Luo et al., 2021), or behavioral nudges (Huang et al., 2018; Terzioğlu et al., 2023; Eigen et al., 2018).
@@
 Appropriate meal frequency may promote weight management and improve insulin sensitivity and blood glucose levels (Holmbäck et al., 2010; Carlson et al., 2007).
-Along with maintaining regular meal times, consistent fluid intake is also essential for supporting bodily functions, such as temperature regulation, nutrient transportation, and waste removal.
 In current approaches, target intake timing is typically defined by the regular intervals between meals or water intake for appropriate eating frequency or to prevent dehydration (Xie et al., 2023; Bi et al., 2016; Fortmann et al., 2014; Varun et al., 2023; Lessel et al., 2016; Ravindran et al., 2022), as well as the optimal timing for meals consumption throughout the day to promote regular and healthy eating habits (Liang, 2024; Faltaous et al., 2021; Ye et al., 2016).
 4.1.3. Environment
-The environment in which an individual eats can significantly impact ingestive behaviors.
-Both social, cultural, and physical environments contribute to shaping food choices, eating patterns, and the overall eating experience.
-For example, the availability of healthy food options, social influences, and food cues in the surroundings can either promote or hinder healthy ingestive behaviors.
+Ingestive behaviors are shaped by a combination of social, cultural, and physical environments, which influence food choices, eating patterns, and the broader eating experience.
 Research has shown that eating location and the surrounding environment can notably affect the intake of unhealthy foods.
@@
 Interventions targeting process behaviors typically occur within the duration of an ingestion episode, although they may also span a longer ingestion period.
-These behaviors, including intake quantity, ingestion pace and oral motor activities, food awareness or emotional state, and safe eating practices, directly impact food intake and digestion.
-argeting these aspects aims to improve eating habits, prevent overeating, promote mindful consumption, and reduce eating-related safety risks.
+These behaviors-including intake quantity, ingestion pace, oral motor patterns, food awareness, emotional state, and safe eating practices—directly influence food intake and digestion.
 4.2.1. Intake Quantity
@@
 Interventions targeting intake quantity typically focus on two key aspects: controlling solid food portion sizes and managing liquid intake to promote healthier eating and drinking behaviors.
-Solid Food: In this context, solid food refers to the ingestion of meals or snacks, excluding water and other beverages.
-By promoting better awareness and control of food quantities, these interventions aim to prevent overeating, reduce the risk of obesity, and support healthy ingestive behaviors, including the prevention of eating disorders (ED) (Rolls et al., 2002; Livingstone and Pourshahidi, 2014).
+i.
+Solid Food: Solid food refers to the ingestion of meals or snacks, excluding water and other beverages.
+These interventions aim to enhance awareness and control of food quantities to prevent overeating and reduce obesity risk.
+They also support healthy ingestive behaviors, including the prevention of eating disorders (Rolls et al., 2002; Livingstone and Pourshahidi, 2014).
 Interventions targeting solid food portion control often measure intake quantity by assessing changes in food mass before and after a meal (Xie et al., 2023; Chaudhry et al., 2016; Chen et al., 2022; Alshurafa et al., 2015; Scisco et al., 2011; Maramis et al., 2014; Farooq et al., 2017; Moses et al., 2023; Li and Bailenson, 2018).
-Liquid: In this category, liquid intake specifically refers to the consumption of water and beverages.
-Most interventions targeting liquid intake emphasize the importance of maintaining adequate hydration throughout the day.
-Sufficient fluid intake is crucial for supporting various physiological functions, including temperature regulation, cardiovascular and cognitive health, and preventing conditions such as kidney stones and urinary tract infections.
-It may also aid in managing energy intake and weight by promoting healthier beverage choices (Popkin et al., 2010; Daniels and Popkin, 2010).
+ii.
+Liquid: Liquid intake refers to the consumption of water and beverages.
+Most interventions in this category emphasize maintaining adequate hydration to support physiological functions and promote healthier beverage choices.
+These strategies may also aid in energy balance and weight management (Popkin et al., 2010; Daniels and Popkin, 2010).
 Interventions focusing on fluid intake typically measure water consumption by tracking changes in volume, such as the amount consumed between two time points or the total daily intake (Faltaous et al., 2021; Alshurafa et al., 2015; Kaner et al., 2018; Poddar et al., 2024; Kreutzer et al., 2015; Varun et al., 2023; Bobin et al., 2018; Lessel et al., 2016; Yıldız and Coşkun, 2019; Chiu et al., 2009; Ko et al., 2007; Pankajavalli et al., 2017; Wijanarko et al., 2019; Tommy et al., 2017; Ravindran et al., 2022; Zhou et al., 2021).
 4.2.2. Ingestion Pace and Oral Motor Activities
-Ingestion pace and oral motor behaviors play a crucial role in regulating energy intake and promoting healthy eating patterns.
-Interventions targeting these behaviors focus on modifying the speed and manner in which food is consumed to enhance satiety, prevent overeating, and support weight management.
+Interventions targeting ingestion pace and oral motor behaviors focus on modifying the speed and manner in which food is consumed to enhance satiety, prevent overeating, and support weight management.
 Faster eating rates have been consistently associated with higher energy intake, increased risk of overweight and obesity, and poorer appetite regulation.
@@
 Current approaches typically target three main aspects: chewing and swallowing behaviors, bite intervals and size, and total meal duration.
+i.
 Chewing and swallowing behaviors: Chewing and swallowing behaviors are key components influencing the pace of eating and the overall ingestion process.
-Many interventions target chewing behavior specifically by encouraging individuals to chew more times per bite or to slow down the rate of chewing.
-By increasing the number of chews per mouthful, individuals can naturally slow their eating pace, enhance satiety signaling, and reduce overall food intake, which are important strategies for managing weight and preventing overeating (Kadomura et al., 2013b; Kleinberger et al., 2023; Koizumi et al., 2011; Bi et al., 2016; Scisco et al., 2011; Maramis et al., 2014; Chen and Yen, 2022; Sugita et al., 2018; Turner-McGrievy et al., 2017; Hori et al., 2023).
-Some interventions combine the regulation of chewing and swallowing, aiming to control the full oral processing cycle rather than focusing solely on chewing.
+Many interventions encourage individuals to chew more times per bite or slow down the rate of chewing.
+These practices can naturally slow their eating pace, enhance satiety signaling, and reduce overall food intake, which supports weight management and prevention of overeating (Kadomura et al., 2013b; Kleinberger et al., 2023; Koizumi et al., 2011; Bi et al., 2016; Scisco et al., 2011; Maramis et al., 2014; Chen and Yen, 2022; Sugita et al., 2018; Turner-McGrievy et al., 2017; Hori et al., 2023).
+Others combine the regulation of chewing and swallowing, aiming to control the full oral processing cycle rather than focusing solely on chewing.
 These approaches aim to slow ingestion by coordinating the timing between chewing and swallowing, ensuring thorough oral processing before taking the next bite (Kim et al., 2016b; Kamachi et al., 2023).
@@
 These interventions promote safer, more mindful eating practices, improving digestion and reducing the risk of eating-related complications (Biyani et al., 2019; Zheng et al., 2025).
-Bite Intervals and Size: Bite size and the interval between bites are critical aspects of eating pace regulation and have been targeted in current interventions.
-This strategy allows more time for satiety signals to develop, helping individuals recognize fullness earlier and thus reducing overall food intake.
+ii.
+Bite Intervals and Size: As critical components of eating pace regulation, bite size and inter-bite intervals have been intervened to allow more time for satiety signals to develop, helping individuals recognize fullness earlier and thus reducing overall food intake.
 Most interventions focus on extending the time interval between successive bites, encouraging individuals to pause longer before taking the next bite (Nakamura et al., 2023; de Vries et al., 2020; Lee et al., 2019; Kim et al., 2016a; Kim and Bae, 2018b, a; Mendi et al., 2013; Hermans et al., 2017).
-In addition to manipulating bite intervals, the amount of food consumed in a single bite is also a key target behavior.
-By reducing bite size, individuals naturally slow down their eating pace and consume fewer calories during a meal (Chen et al., 2022).
-Total Meal Duration: Total meal duration — the overall time taken to complete a meal — is another important target for interventions aimed at regulating eating pace.
-Most studies monitor and estimate total meal duration by tracking changes in food weight over time, providing an objective measure of ingestion pace.
-Many interventions focus on prolonging meal duration to slow down eating speed, helping individuals better recognize satiety cues and prevent overeating, which supports healthier dietary behaviors (Moses et al., 2023; Zhang et al., 2019; Kim et al., 2016c; Ioakimidis et al., 2009).
-Some interventions, particularly those targeting young children, aim to prevent meals from extending excessively long.
-These interventions help children establish structured, healthy eating routines early in life (Krishna et al., 2025; Lo et al., 2007).
+Bite size is another key behavioral target.
+Reducing the amount of food per bite naturally slows eating pace and lowers overall caloric intake during a meal (Chen et al., 2022).
+iii.
+Total Meal Duration: The overall time taken to complete a meal is typically estimated by tracking changes in food weight over time to provide an objective measure of ingestion pace.
+Many interventions focus on prolonging meal duration to slow down eating speed, helping individuals better recognize satiety cues and prevent overeating (Moses et al., 2023; Zhang et al., 2019; Kim et al., 2016c; Ioakimidis et al., 2009).
+Others, particularly those involving young children, seek to avoid excessively long meals to help establish structured eating routines early in life (Krishna et al., 2025; Lo et al., 2007).
 Additionally, some interventions are designed to synchronize eating paces in social settings, where regulating total meal duration across a group helps reduce discomfort caused by uneven eating speeds (Mitchell et al., 2015).
@@
 Interventions targeting food awareness or emotional state focus on enhancing individuals’ cognitive and emotional engagement with their eating experiences to promote healthier behaviors.
-A key approach in this area is encouraging greater attention to food during meals, either through practices like mindful eating, which emphasizes being fully present during eating (Kleinberger et al., 2023; Koizumi et al., 2011; Parra et al., 2023; Wang et al., 2025; Khot et al., 2020), or through interventions designed to help children develop better focus on their food (Kadomura et al., 2013b; Han and Kang, 2017).
-In addition to promoting awareness during ingestion, some interventions focus on improving individuals’ self-perception and cognitive evaluation of their ingestive behaviors after meals.
+A key approach is encouraging greater attention to food during meals, either through practices like mindful eating, which emphasizes being fully present during eating (Kleinberger et al., 2023; Koizumi et al., 2011; Parra et al., 2023; Wang et al., 2025; Khot et al., 2020), or through interventions designed to help children develop better focus on their food (Kadomura et al., 2013b; Han and Kang, 2017).
+Other interventions focus on improving individuals’ self-perception and cognitive evaluation of their ingestive behaviors after meals.
 By increasing reflective thinking about eating habits, these interventions help individuals build healthier long-term dietary patterns and strengthen their self-regulation capacity (Sun et al., 2020; Zhang and Parker, 2020; Tang et al., 2015).
-Moreover, another line of interventions aims to enhance the emotional quality of the eating experience itself, such as promoting positive emotions through novel experiences and reducing negative emotions like anxiety during eating.
+Moreover, a separate line of interventions aims to enhance the emotional quality of the eating experience itself, such as promoting positive emotions through novel experiences and reducing negative emotions like anxiety during eating.
 A healthy emotional state during ingestion has been linked to better digestion, improved dietary choices, and a lower risk of disordered ingestive behaviors (Mayumi et al., 2023; Carroll et al., 2013; Mitchell et al., 2015; Narumi et al., 2011).
 4.2.4. Safe Eating Practices
-Safe eating practices aim to ensure that the ingestion process does not pose additional health risks.
-Current interventions in this area primarily target two aspects: hygiene behaviors and food temperature control.
+Safe eating practices are intended to minimize health risks during the ingestion process.
+Current interventions primarily target two aspects: hygiene behaviors and food temperature control.
 Hygiene-related interventions focus on promoting safe practices during eating to prevent foodborne infections (Wang and Zhang, 2021).
 Other interventions address food temperature control to prevent oral injuries, such as burns (Singhal et al., 2024).
-5. Structures and Design Space for Behavioral Sensing and Intervention Paradigms
-This chapter analyzes the structural and modality-level characteristics of ingestion health paradigms, as outlined in Chapter 3.2, with a focus on two essential stages of the behavioral loop: sensing and feedback/intervention.
-While reasoning processes typically mediate between perception and action, they are often embedded within system pipelines, reflected through how sensing data is interpreted and applied.
-Therefore, our analysis centers on the sensing and feedback/intervention components, where modality design plays a direct and observable role.
-Inspired by Jaimes and Sebe (2007) and Jansen et al.
-(2022), we adopt modality as the classification standard for both the sensing and intervention components, encompassing interactive devices and perceptual computing methods.
-Specifically, Jaimes and Sebe (2007) define modality as “mode of communication according to human senses and computer input devices activated by humans or measuring human qualities,” emphasizing that modality serves as a communication channel designed around human perception and computational input-output mechanisms.
-Furthermore, they introduce a human-based perspective on modality, identifying core modalities corresponding to human senses, including vision, touch, hearing, smell, and taste, and categorizing more complex examples within these sensory categories.
-Building on this foundation, Jansen et al.
-(2022) expand the concept of human-based modality by introducing two additional categories—cerebral and cardiac modalities—to accommodate advances in interactive device technologies.
-Following insights from Jaimes and Sebe (2007), we recognize that, in the sensing process, certain modalities do not directly correspond to individual human sensory channels but involve interactions mediated through external devices or environmental signals.
-Thus, we introduce the category of environment-based modalities, capturing input and feedback mechanisms targeting environmental objects independently of human senses.
-This extension provides a more comprehensive classification for sensing and intervention designs in ingestion health systems, covering modalities arising from device manipulation, environmental interactions, or indirect physiological monitoring.
-We adapt and refine this modality classification to suit ingestion health interventions, distinctly addressing both sensing and feedback/intervention stages.
-The primary classification human-based modality contains visual, auditory, haptic, gustatory, olfactory, motion, and physiological, and environment-based modality contains Visual, Acoustic, Force, Spectral, and Electrical.
-Almost each modality further includes submodalities considering focus and implications different from sensing and feedback/intervention.
-Accordingly, slight structural and descriptive variations in modality classifications are introduced to align precisely with each phase of the behavioral closed-loop, as shown in following subsections.
+5. Structures and Design Space from Behavioral Sensing to Intervention Paradigms
+This section analyzes the modality-level characteristics of ingestion health systems, focusing on two essential stages of the behavioral loop: sensing and intervention.
+While reasoning processes mediate between perception and action, they are typically embedded within pipelines, reflected in how sensing data is interpreted and applied.
+Therefore, our analysis centers on the sensing and intervention components, where the design of modality plays a direct and observable role in shaping user experience and effectiveness of the system.
+Prior research has established several perspectives on modality.
+Jaimes and Sebe (2007) define modality as ”modes of communication according to human senses and computer input devices activated by humans or measuring human qualities”, emphasizing the centrality of perception and input-output channels.
+Building on this view, Jansen et al.
+(2022) extend the concept by introducing cerebral and cardiac modalities to accommodate emerging interactive technologies.
+However, in ingestion-related systems, certain modalities do not correspond directly to individual human sensory channels but instead involve environmental interactions or indirect physiological monitoring.
+To address this, we introduce the category of environment-based modalities, which encompass input and feedback mechanisms targeting environmental objects independently of human senses.
+This extension enables a more comprehensive classification of sensing and intervention designs in ingestion health interventions.
+Building on these foundations, we classify human-based modalities into visual, auditory, haptic, gustatory, olfactory, motion, and physiological, and environment-based modalities into Visual, Acoustic, Force, Spectral, and Electrical.
+Slight structural and descriptive variations in modality classifications are introduced below to align precisely with each phase of the behavioral closed-loop.
 5.1. Sensing
-This subsection focuses on the sensing modalities within overall ingestion health intervention paradigms, where behavioral data are primarily acquired through various sensor-based or perceptual computing-driven modalities.
+This subsection focuses on sensing modalities within overall ingestion health intervention paradigms, where behavioral data are primarily acquired through various sensor-based or perceptual computing-driven modalities.
 The detailed structure and descriptions of each modality, along with its associated submodalities in the sensing process, are summarized in Table 5.
 Notably, gustatory and olfactory modalities are absent in the sensing stage across existing systems.
-We believe that sensing gustatory or olfactory information through perceptual or sensor-based means remains technically challenging, and these modalities are mainly employed in the feedback/intervention phase rather than for direct behavioral sensing.
-The following sections present the identified sensing modalities, categorized according to their classifications, with representative approaches within ingestion health intervention paradigms.
-Table 1. Descriptions and category of modality terms in sensing process
-Modality Submodality Description Paper count
-Visual Human-based visual captures human-derived visual cues to infer the user’s attention, emotional state, or physical actions during ingestion-related behaviors 8
-[…]
-Facial expression Facial expression analyzes dynamic changes in facial muscle movements to detect emotional responses, physiological states, or behavioral reactions associated with eating activities 4
-Body posture Body posture captures the orientation, alignment, and movement of the user’s body or head to interpret ingestion-related behaviors via camera, such as eating gestures, drinking actions, or body positioning during meals 3
-[…]
-[…]
-Sound Sound captures non-verbal acoustic signals generated during ingestion activities, such as chewing sounds, swallowing noises, or utensil interactions, to infer ingestive behaviors or consumption events 8
-[…]
-[…]
-[…]
-Chewing and Micro-Head Motion Chewing and micro-head motion monitors rhythmic jaw movements and subtle head or neck dynamics associated with eating, drinking, or swallowing 5
-[…]
-[…]
-Metabolic Metabolic captures nutritional and biochemical markers such as blood glucose, cholesterol, and body fat rate, typically using biosensors to assess metabolic responses and long-term ingestion-related health status 6
-Human-based Autonomic Autonomic captures involuntary signals, including electrodermal activity (EDA) and skin temperature, using biosensors to monitor autonomic activity during ingestion 7
-Visual Environment-based visual captures information about food items, environments, or objects through image-based techniques such as classification, segmentation, or object detection 29
-[…]
-[…]
-[…]
-[…]
-[…]
-Environment-based Magnetic Magnetic sensing detects the presence, identity, or proximity of ingestion-related objects using techniques such as Radio-Frequency Identification (RFID) or Near-Field Communication (NFC) technologies 5
+We believe that sensing gustatory or olfactory information through perceptual or sensor-based means remains technically challenging, and these modalities are mainly employed in the intervention phase rather than for direct behavioral sensing.
+The identified sensing modalities is categorized according to their classifications below.
+A summary of sensing modalities and corresponding publications included in this review is shown in Table 3.
 5.1.1. Human-based
-Visual: Visual sensing has been widely adopted in ingestion-related systems to extract behavioral and attentional cues through camera-based observation.
-We identified three common submodalities within visual sensing: gaze, facial features, and body movement.
-Gaze-based sensing tracks eye movements and fixation points to assess users’ attention during food-related interactions.
-For example, Karkar et al.
-(2018) integrated an eye tracker into a VR system to analyze children’s visual attention toward nutrition-themed stimuli.
-Webcam-based gaze estimation has also been used to infer attentional disengagement during meals, which can inform timely interventions (Khot et al., 2020).
-Facial-expression-based sensing is primarily used to detect emotional or behavioral responses and confirm user presence.
-A webcam was employed to extract facial expressions and assess affective states during meals (Mayumi et al., 2023).
-Moreover, Kim and Bae (2018a) used lip and jaw motion analysis to model bite rhythm.
-Face detection has also supported social presence recognition in multi-user contexts, helping to determine whether a child is present (Triantafyllidis et al., 2019) or enabling socially responsive camera-based interaction (Robinson et al., 2020).
-Body-based visual sensing targets ingestion-related gestures and postures, such as hand movements and upper-body orientation.
-Xie et al.
-(2023) tracked caregivers’ feeding gestures to estimate feeding frequency and potential satiety cues in children.
-Skeleton tracking with depth cameras has been used to monitor upper-body posture during meals (Mayumi et al., 2023).
-Additionally, egocentric video data were collected for post-hoc analysis of mealtime behaviors and social dynamics (Zhang and Parker, 2020).
-Auditory: Auditory sensing in ingestion health systems typically involves capturing either verbal input or eating-related sounds, which can be classified into two main submodalities: speech and non-verbal sound, covering existing approaches.
-Speech-based auditory sensing is commonly used for semantic data acquisition or enabling conversational interfaces.
-Spoken input has been leveraged to extract structured dietary information using linguistic pattern recognition (Prasetyo et al., 2020).
-Systems such as those proposed by Mayumi et al.
-(2023) and Robinson et al.
-(2020) receive users’ queries for food-related facts or recommendations and extract key information from speech signals.
-Speech has also been employed as an input modality for logging intake events or interacting with nutrition tracking systems (Luo et al., 2021; Arefeen et al., 2022).
-Parra et al.
-(2023) used Wizard-of-Oz setups and processed users’ responses for further system-guided prompting.
-Sound-based sensing focuses on detecting physiological cues, such as chewing or swallowing, using microphones.
-Some researchers have placed microphones near the throat (Kalantarian et al., 2014) or collar (Chen and Yen, 2022) to capture oral sound signals during ingestion, which can be analyzed for bite detection (Bi et al., 2016), food classification (Kleinberger et al., 2023), or intake rhythm estimation.
-Microphones have also been implemented in earpiece form factors to detect rhythmic mastication events, such as chewing sensors (Maramis et al., 2014) and bone conduction microphones (Kamachi et al., 2023), which capture sound with reduced ambient interference.
-Additionally, although categorized separately above, some microphones are designed for specific contexts, such as capturing chewing in denture wearers (Koizumi et al., 2011) or jointly recording speech and chewing sounds as contextual data (Zhang and Parker, 2020; Kamachi et al., 2023).
+i. Visual: Visual sensing has been widely adopted in ingestion-related systems to extract behavioral and attentional cues through camera-based observation. We identified three common submodalities:
+Gaze sensing tracks eye movements and fixation points to assess users’ attention during food-related interactions.
+It has been used to measure attentional engagement with food stimuli across settings such as VR eye tracking (Karkar et al., 2018) or webcam-based monitoring (Khot et al., 2020).
+Facial-expression sensing primarily support both emotional and behavioral inference in ingestion-related systems.
+It has been used to detect affective states (Mayumi et al., 2023), model chewing through lip and jaw motion (Kim and Bae, 2018a), and recognize user presence in multi-user contexts (Triantafyllidis et al., 2019; Robinson et al., 2020).
+Body posture sensing interprets behaviors by capturing body and head orientation, alignment, and movement during ingestion.
+Posture-related signals were captured by tracking the feeding gestures of children for satiety signals (Xie et al., 2023), monitoring the upper-body orientation using skeleton models (Mayumi et al., 2023), and analyzing egocentric video for mealtime behavior (Zhang and Parker, 2020).
+ii. Auditory: Auditory sensing in ingestion health systems typically captures either verbal input or eating-related sounds, which can be classified into two main submodalities:
+Speech sensing commonly supports semantic data acquisition and conversational interfaces.
+Spoken input has been used to extract structured dietary data via linguistic pattern recognition (Prasetyo et al., 2020), facilitate food-related recommendations from spoken queries (Mayumi et al., 2023; Robinson et al., 2020), and log intake events with interaction (Luo et al., 2021; Arefeen et al., 2022), as well as to enable system-guided prompting via Wizard-of-Oz protocols (Parra et al., 2023).
+Non-verbal sound sensing focuses on physiological cues like chewing or swallowing.
+Microphones have been placed near the throat (Kalantarian et al., 2014) or collar (Chen and Yen, 2022) to detect bites (Bi et al., 2016), classify food types (Kleinberger et al., 2023), or estimate intake rhythms.
+Earpiece-mounted sensors, including chewing microphones (Maramis et al., 2014) and bone conduction devices (Kamachi et al., 2023), help reduce ambient noise.
+Some systems further adapt to context, such as denture-based chewing detection (Koizumi et al., 2011) or setups capturing both speech and eating sounds (Zhang and Parker, 2020; Kamachi et al., 2023).
+iii.
 Haptic: Haptic sensing aims to detect physical interactions or internal mechanical activity through pressure, force, or contact-based sensors.
-A common form is weight sensing, which tracks changes in body weight as a long-term reflection of dietary intake (Ludwig et al., 2011; Bastida et al., 2023; Nedungadi et al., 2018).
-A substantial body of work has focused on haptic sensing using piezoelectric or force sensors.
-Neck-mounted piezoelectric sensors have been embedded into wearable necklaces to capture throat vibrations associated with swallowing (Alshurafa et al., 2015; Kim et al., 2016b; Abisha and Rajalakshmy, 2017).
-Similar techniques have been used to detect mechanical displacement in the temporalis muscle to infer chewing activity (Farooq et al., 2017).
-Other studies have explored intraoral pressure detection, such as placing sensors on palatal retainers to measure tongue force (Biyani et al., 2019), or using external force sensors to estimate masticatory force (Zheng et al., 2025).
-Additionally, pressure-sensitive buttons or interfaces have been applied to allow users to log ingestion-related events (Fortmann et al., 2014; Tang et al., 2015; Cai et al., 2024; Robinson et al., 2020).
-Motion: Motion sensing captures gestures and subtle physical activity using inertial or proximity sensors.
-Most gesture-oriented systems employ inertial measurement units (IMUs) mounted on utensils (Wang et al., 2025; Khot et al., 2020; Zhang et al., 2019; Li and Bailenson, 2018; Kadomura et al., 2013a, 2014), wristbands (Kim et al., 2016a; Kim and Bae, 2018b; Ye et al., 2016; Turner-McGrievy et al., 2017; Mendi et al., 2013; Kumbhare et al., 2023; Gaikwad et al., 2024; Dragoni et al., 2017; Scisco et al., 2011), or containers (Tommy et al., 2017; Bobin et al., 2018; Chiu et al., 2009; Zhou et al., 2021) to detect hand-to-mouth movement, bite frequency, or drinking actions.
-IMUs have also been integrated into smart forks (Hermans et al., 2017), smart glasses (Bedri et al., 2022), and smartphones (Tirasirichai et al., 2018).
-Additionally, some platforms or frameworks use pedometers (Harous et al., 2017) and activity trackers (Alexander et al., 2017; Vazquez et al., 2015; Fuchs et al., 2019) to estimate ingestion or physical activity context.
-These approaches have made it easier for broader health-monitoring systems to achieve dietary reasoning support (Bastida et al., 2023; Mishra et al., 2024; Maramis et al., 2014; Chukwu, 2011; Badawi et al., 2012a; Jung et al., 2007).
-Chewing and micro-head motion sensing focus on jaw, throat, or mouth-level interactions.
+A common application is weight sensing, which serves as a long-term indicator of dietary intake (Ludwig et al., 2011; Bastida et al., 2023; Nedungadi et al., 2018).
+Wearable piezoelectric and force sensors have been used to capture throat vibrations during swallowing (Alshurafa et al., 2015; Kim et al., 2016b; Abisha and Rajalakshmy, 2017) and temporalis muscle displacement for chewing estimation (Farooq et al., 2017).
+Other methods include intraoral pressure sensors embedded in palatal retainers to measure tongue force (Biyani et al., 2019), external sensors for estimating masticatory force (Zheng et al., 2025), and pressure-sensitive interfaces for logging ingestion events (Fortmann et al., 2014; Tang et al., 2015; Cai et al., 2024; Robinson et al., 2020).
+iv. Motion: Motion sensing captures gestures and subtle physical activity using inertial or proximity sensors.
+Most gesture-oriented systems employ inertial measurement units (IMUs) mounted on utensils (Wang et al., 2025; Khot et al., 2020; Zhang et al., 2019; Li and Bailenson, 2018; Kadomura et al., 2013a, 2014; Hermans et al., 2017), containers (Tommy et al., 2017; Bobin et al., 2018; Chiu et al., 2009; Zhou et al., 2021), wristbands (Kim et al., 2016a; Kim and Bae, 2018b; Ye et al., 2016; Turner-McGrievy et al., 2017; Mendi et al., 2013; Kumbhare et al., 2023; Gaikwad et al., 2024; Dragoni et al., 2017; Scisco et al., 2011), smart glasses (Bedri et al., 2022) and smartphones (Tirasirichai et al., 2018) to detect hand-to-mouth movement, bite frequency, or drinking actions.
+Some platforms also use pedometers (Harous et al., 2017) and activity trackers (Alexander et al., 2017; Vazquez et al., 2015; Fuchs et al., 2019) to estimate ingestion or physical activity context.
+These approaches have made it easier for broader health-monitoring systems to achieve dietary reasoning support.
+Chewing and micro-head motion sensing target jaw, throat, or mouth-level interactions.
 Biyani et al.
-(2019) used a 3D accelerometer to track throat motion, and Hori et al.
-(2023) applied infrared distance sensing to capture subtle skin vibrations.
-Additionally, electrical conductivity sensors (Kadomura et al., 2013a, b) or static charge sensors (Han and Kang, 2017) were embedded in utensils or cups to infer mouth contact and drinking actions.
+(2019) tracked throat motion using a 3D accelerometer, while Hori et al.
+(2023) used infrared sensors to detect subtle skin vibrations.
+Sensors based on electrical conductivity (Kadomura et al., 2013a, b) or static charge (Han and Kang, 2017) have also been embedded in utensils and cups to infer mouth contact and drinking events.
+v.
 Physiological: Physiological sensing provides insight into users’ internal states that influence or contextualize ingestion behavior, including appetite, metabolic status, stress, and cardiovascular conditions.
-Autonomic sensing typically involves EMG electrodes to detect muscle activity during chewing or wrist motion (Lee et al., 2019; Sugita et al., 2018; Kim et al., 2016b), as well as thermistors and respiratory sensors to track body temperature and breathing patterns (Park and Min, 2007; Jung et al., 2007).
-Emotional arousal or stress is commonly inferred through electrodermal activity (EDA) sensors (Carroll et al., 2013; Nedungadi et al., 2018), which may affect eating pace or impulsive behaviors.
+Autonomic sensing typically involves EMG electrodes to detect muscle activity during chewing or wrist motion (Lee et al., 2019; Sugita et al., 2018; Kim et al., 2016b), along with thermistors and respiratory sensors to track body temperature and breathing patterns (Park and Min, 2007; Jung et al., 2007).
+Emotional arousal or stress is inferred through electrodermal activity (EDA) sensors (Carroll et al., 2013; Nedungadi et al., 2018).
 Metabolic sensing employs BIA sensors to estimate body fat composition (Chukwu, 2011; Balbin et al., 2020), and incorporates CGM and cholesterol sensors to continuously monitor glucose levels and lipid markers (Liang, 2024; Harous et al., 2017; Arefeen et al., 2024; Nedungadi et al., 2018).
-Cardiac sensing captures heart rate, blood pressure, and oxygen saturation using ECG (Badawi et al., 2012b), pulse wave (Varun et al., 2023), or optical PPG sensors (Park and Min, 2007; Ludwig et al., 2011; Jung et al., 2007; Huang et al., 2018; Nedungadi et al., 2018), which are frequently used to contextualize ingestive behavior or guide personalized intervention timing.
-Others: Apart from the sensing modalities detailed above, there are several approaches that do not involve sensing or perceptual computing methods.
-These approaches rely entirely on user-reported data collected through manual input.
-Common input methods include the manual entry of dietary events or physical status, image selection or annotation, and questionnaire-style logging via mobile or chatbot interfaces.
-Although these systems may still provide feedback, they fall outside our sensing modality framework due to the lack of automated perceptual input or computational sensing logic.
-A detailed overview of this manual approach is summarized in Table 7, with approximately 20 publications adopting this method.
+Cardiac sensing captures heart rate, blood pressure, and oxygen saturation using ECG (Badawi et al., 2012b), pulse wave (Varun et al., 2023), or optical PPG sensors (Park and Min, 2007; Ludwig et al., 2011; Jung et al., 2007; Huang et al., 2018; Nedungadi et al., 2018) to support context-aware interpretation and intervention timing.
+vi.
+Others: Without sensing or perceptual computing, other approaches rely entirely on manually collected self-reports.
+Common inputs include manual logging of dietary events or physical states, image selection or annotation, and questionnaire-style via mobile or chatbot interfaces.
+Though such systems can support reasoning and intervention, they fall outside our sensing stages due to the lack of automated perceptual input or computational sensing logic.
 5.1.2. Environment-based
+i.
 Visual: Environment-based visual sensing focuses on object-level perception of food and eating-related artifacts to support dietary monitoring and context acquisition.
-Most systems utilize food image analysis, applying computer vision to classify food types, estimate portion sizes, or infer nutritional content from captured images (Kanjalkar et al., 2024; Eigen et al., 2018; Cordeiro et al., 2015; YR et al., 2024; Epstein et al., 2016).
-These methods are commonly implemented through smartphone photography (Mishra et al., 2024; Waltner et al., 2015; Achilleos et al., 2017), utensil-mounted or refrigerator cameras (Nakamura et al., 2023; Kwon et al., 2016; Faltaous et al., 2021), or wearable vision modules (Fuchs et al., 2019).
-Some systems use food color (Kadomura et al., 2013a, 2014), utensil-based visual cues (Wang and Zhang, 2021), or visual distortion sensing (Akter et al., 2021; Chiu et al., 2009) to infer consumption context or eating correctness.
-Other approaches incorporate plate-level dynamics (Ganesh et al., 2014), receipt parsing (Mankoff et al., 2002), puzzle-based interaction (Cai et al., 2024), or packaging and interface-based food input (Arefeen et al., 2022; Hsiao et al., 2011) for supplemental inference and logging.
-Additionally, visual code recognition is widely used to access structured nutritional information through barcode scanning (Volkova et al., 2016; Eyles et al., 2017), QR code interpretation via mobile interfaces (Fernandez et al., 2013; Vazquez et al., 2015; Alloghani et al., 2016), and edible marker detection in AR-enhanced scenarios (Gutiérrez et al., 2018).
+Most systems apply computer vision to classify food types, estimate portions sizes, or infer nutrition from images (Kanjalkar et al., 2024; Eigen et al., 2018; Cordeiro et al., 2015; YR et al., 2024; Epstein et al., 2016) ,via smartphone photography (Mishra et al., 2024; Waltner et al., 2015; Achilleos et al., 2017), utensil- or refrigerator-mounted cameras (Nakamura et al., 2023; Kwon et al., 2016; Faltaous et al., 2021), or wearable modules (Fuchs et al., 2019).
+Other methods use food color (Kadomura et al., 2013a, 2014), utensil-based cues (Wang and Zhang, 2021), visual distortion (Akter et al., 2021; Chiu et al., 2009), or plate dynamics (Ganesh et al., 2014) to infer consumption.
+Additional approaches include receipt parsing (Mankoff et al., 2002), puzzle-based interaction (Cai et al., 2024), packaging input (Arefeen et al., 2022; Hsiao et al., 2011), and visual code recognition such as barcode scanning (Volkova et al., 2016; Eyles et al., 2017), QR code interpretation (Fernandez et al., 2013; Vazquez et al., 2015; Alloghani et al., 2016), and edible marker detection in AR scenarios (Gutiérrez et al., 2018).
+ii.
 Acoustic: Acoustic sensing often utilizes ultrasonic distance sensors to monitor liquid intake or food-related actions.
-Most related works involve embedding ultrasonic transducers into smart bottles or storage containers to measure fluid levels and infer hydration behavior (Poddar et al., 2024; Wijanarko et al., 2019; Ravindran et al., 2022; Varun et al., 2023; Tommy et al., 2017).
-These systems estimate drinking volume based on real-time fluid height or remaining capacity.
-Other systems apply similar sensing techniques to detect discrete ingestion-related events, such as the removal of a snack from a container using ultrasonic signal disruption (Liu et al., 2024), or monitoring food depletion in a storage cabinet based on internal distance measurements (Faltaous et al., 2021).
-Force: Force sensing typically involves the use of load cells or pressure sensors embedded in plates, coasters, or tables to measure changes in food or liquid weight over time.
-This approach enables passive monitoring of eating dynamics, such as intake amount, bite frequency, or consumption rate.
-Smart plates equipped with load cells (Mitchell et al., 2015; Kim et al., 2016c; Maramis et al., 2014; Ioakimidis et al., 2009; de Vries et al., 2020) or force-sensitive resistors (FSRs) (Jung et al., 2017) have been used to continuously track food depletion and infer eating pace or satiety.
-Beyond intake tracking, similar sensing mechanisms have been extended to estimate nutrient intake across meals (Krishna et al., 2025; Zhao et al., 2021; de Vries et al., 2020; Jung et al., 2017), monitor liquid consumption via drink weight changes (Ritschel et al., 2018; Lessel et al., 2016), or support playful interactions through weight-sensitive input (Lo et al., 2007).
-Additionally, multi-plate setups have been applied to detect food preferences and portion control behaviors (Joi et al., 2016; Hermans et al., 2017).
-Spectral: Spectral sensing leverages temperature, proximity, and light-based sensors to capture physical or chemical properties of food, beverages, and environmental conditions.
-Common sensing components include infrared proximity detectors, temperature sensors, and light-spectrum analyzers.
-Specifically, infrared or optical sensors have been used to detect the presence of containers (Ritschel et al., 2018), estimate proximity to food (Bedri et al., 2022), or assess food temperature (Singhal et al., 2024).
-Temperature sensors embedded in bottles or utensils help monitor liquid warmth (Tommy et al., 2017; Wijanarko et al., 2019; Varun et al., 2023) or environmental conditions affecting food stability (Bastida et al., 2023; Badawi et al., 2012b; Kwon et al., 2016).
-In more advanced cases, Matsui et al.
-(2018) applied spectral absorption analysis to estimate the alcohol concentration in beverages and assess drinking speed.
-Electrical: Electrical sensing supports ingestion-related monitoring by detecting changes in conductivity, capacitance, or magnetic response, allowing for the inference of liquid consumption, food identity, or environmental conditions.
-This method enables passive, contact-based tracking of ingestion without requiring user instrumentation.
-Common sensor types include conductivity probes, humidity sensors, capacitive pads, and RFID/NFC modules.
-Specifically, conductivity and humidity sensors have been embedded into bottles or utensils to monitor water levels or detect drinking events (Bobin et al., 2018; Tommy et al., 2017; Kreutzer et al., 2015; Varun et al., 2023), while capacitive level sensing has been used to estimate water volume based on changes in dielectric properties (Kaner et al., 2018).
-Some systems integrate float switches or conductive tilt detection for recognizing bottle use (Pankajavalli et al., 2017; Ko et al., 2007).
-Ambient humidity is also used to characterize dining environments and inform food safety assessments (Bastida et al., 2023; Badawi et al., 2012b; Kwon et al., 2016).
-On the magnetic side, RFID and NFC technologies enable automatic identification of food items, storage states, or utensil presence (Vazquez et al., 2015; Huang et al., 2010; Jung et al., 2017), and advanced NFC designs even estimate consumed volume through passive milliliter tracking (Yıldız and Coşkun, 2019).
+Most systems embed ultrasonic transducers in bottles or containers to track fluid levels and estimate drinking volume (Poddar et al., 2024; Wijanarko et al., 2019; Ravindran et al., 2022; Varun et al., 2023; Tommy et al., 2017).
+Others apply similar ultrasonic sensors to detect discrete events like snack removal (Liu et al., 2024), or monitoring food depletion in storage cabinets (Faltaous et al., 2021).
+iii.
+Force: Force sensing typically uses load cells or pressure sensors embedded to measure changes in food or liquid weight over time.
+Smart utensils equipped with load cells (Mitchell et al., 2015; Kim et al., 2016c; Maramis et al., 2014; Ioakimidis et al., 2009; de Vries et al., 2020; Lessel et al., 2016; Ritschel et al., 2018) or force-sensitive resistors (FSRs) (Jung et al., 2017) have been used to continuously track food or liquid depletion and infer ingestion pace or satiety.
+Building on ingestion tracking, systems aim to estimate nutrient intake (Krishna et al., 2025; Zhao et al., 2021; de Vries et al., 2020; Jung et al., 2017), detect food preferences and portion control (Joi et al., 2016; Hermans et al., 2017), and even gamify eating through weight-sensitive interactions (Lo et al., 2007).
+iv.
+Spectral: Spectral sensing commonly leverages infrared detectors, temperature sensors, and light-spectrum analyzers, to capture physical or chemical properties of food, beverages, and environmental conditions.
+Infrared or optical sensors have been used to detect the presence of containers (Ritschel et al., 2018), estimate proximity to food (Bedri et al., 2022), or assess food temperature (Singhal et al., 2024).
+Temperature sensors embedded in utensils help monitor liquid warmth (Tommy et al., 2017; Wijanarko et al., 2019; Varun et al., 2023) or environmental conditions affecting food stability (Bastida et al., 2023; Badawi et al., 2012b; Kwon et al., 2016).
+In addition, Matsui et al.
+(2018) used spectral absorption to estimate alcohol content and drinking rate.
+v.
+Electrical: Electrical sensing supports ingestion monitoring by detecting changes in conductivity, capacitance, or magnetic response.
+It enables passive contact-based tracking without user instrumentation.
+Common sensors include conductivity probes, humidity sensors, capacitive pads, and RFID/NFC modules.
+Conductivity and humidity sensors embedded in utensils can monitor water levels and drinking events (Bobin et al., 2018; Tommy et al., 2017; Kreutzer et al., 2015; Varun et al., 2023), while capacitive sensing can estimate volume via dielectric change (Kaner et al., 2018).
+Some systems integrate float switches or conductive tilt detection to identify bottle usage (Pankajavalli et al., 2017; Ko et al., 2007).
+Ambient humidity has also been used to characterize dining environments and inform food safety assessments (Bastida et al., 2023; Badawi et al., 2012b; Kwon et al., 2016).
+Magnetic technologies enable automatic identification of food items, storage states, or utensil presence (Vazquez et al., 2015; Huang et al., 2010; Jung et al., 2017).
+Some advanced RFID or NFC designs estimate consumed volume via passive milliliter tracking (Yıldız and Coşkun, 2019).
 Additionally, Han and Kang (2017) used metal contact detection to support real-time eating activity sensing.
-5.2. Feedback and Intervention
-As the final stage in the ingestion behavior loop, feedback and intervention mechanisms convey information back to users to support behavioral awareness, regulation, or change.
-In contrast to sensing, which focuses on data acquisition through sensors or perceptual computing, intervention emphasizes human interpretation and experience.
-Accordingly, the modalities at this stage are inherently human-based, as they operate by stimulating human sensory channels to deliver prompts, guidance, or other affective cues.
-To systematically present the design of the feedback and intervention process, we do not assess whether these mechanisms rely on dedicated sensors.
-Rather, since all systems included in our review engage in sensor-based or perception-driven processing at some point in the ingestion-related loop, we categorize their feedback and intervention strategies in terms of modality, focusing on how information is conveyed and perceived.
-These modalities are classified based on the primary human sense they target, as summarized in Table 2.
-Table 2. Descriptions and category of modality terms in feedback/intervention process
-Modality Submodality Description Paper count
-Visual Textual Delivering feedback through text-based information, such as messages, notifications, or instructions 85
-Graphical Providing feedback using graphical elements like charts, progress bars, icons, or visualizations to convey behavioral states or intervention cues 41
-Symbolic Giving visual feedback through abstract or simplified representations, such as symbols, flashing lights, or color changes, to convey states or prompts with minimal cognitive load or linguistic interpretation 21
-[…]
-[…]
-Deformation Delivering feedback by altering the physical shape, structure, or volume of an interface, allowing users to perceive behavioral cues through tangible disruptions embedded within their natural interaction actuators 7
-[…]
-[…]
+5.2. Intervention
+Intervention mechanisms, whether active or passive, convey information back to users to support behavioral awareness, regulation, or change.
+To systematically present the design of the intervention process, we do not assess whether these mechanisms rely on dedicated sensors.
+Rather, since all systems included in our review engage in sensor-based or perception-driven processing at some point in the ingestion-related loop, we categorize their intervention strategies in terms of modality, focusing on how information is conveyed and perceived.
+To provide an overview of the types of modalities employed, Table 4 summarizes a taxonomy of intervention modalities and their corresponding publications.
+Accordingly, the modalities at this stage are inherently based on the primary human sense they target to deliver prompts, guidance, or other affective cues, as summarized in Table 6.
 5.2.1. Visual
-Visual feedback is one of the most commonly used modalities in ingestive behavior interventions, providing intuitive, accessible, and scalable means of conveying prompts, progress, or state awareness.
-Based on representation format, we categorize visual feedback into three submodalities: textual, graphical, and symbolic.
-This layered approach to visual design allows systems to balance clarity, immediacy, and ambient subtlety, tailoring feedback complexity to user needs and situational demands.
-A complete list of visual feedback systems and their classified modalities is available in Table 8.
-Textual: Textual submodality presents information explicitly through written content such as notifications, reminders, or dietary instructions.
-These systems often use text-based prompts to communicate recommended actions, reflect behavioral progress, or reinforce decision-making.
-Examples include recommender agents (Zenun Franco, 2017; Arefeen et al., 2022), just-in-time eating assistants (Prasetyo et al., 2020; Carroll et al., 2013), and mobile self-monitoring tools (Burke et al., 2011; Krishna et al., 2025), all of which deliver messages through on-screen text.
-Many additional systems also employ message notifications or app-based dashboards to present dietary data or track goals in textual form, reinforcing the prominence of this submodality in behavior change design (Faltaous et al., 2021; Lee et al., 2019; Haerens et al., 2007).
-Graphical: Graphical methods employ structured visual elements—such as charts, progress bars, or data visualizations—to represent behavior patterns or intervention outcomes.
-For instance, eat2pic translates bite patterns into abstract visual paintings as a playful form of real-time feedback (Nakamura et al., 2023), while many other systems use app interfaces to graph nutritional balance, meal composition, or eating speed, enhancing user self-awareness and reflection (Faltaous et al., 2021; Scisco et al., 2011; Lo et al., 2007).
-Symbolic: Symbolic approaches convey prompts using abstract or simplified visual representations, such as color changes (Bobin et al., 2018), flashing lights (Matsui et al., 2018; Lessel et al., 2016), or emoji-style icons (Kim and Bae, 2018a), which communicate behavioral cues with minimal cognitive load.
-These signals are often embedded in utensils or ambient displays to indicate states like “ready to eat” or “slow down” without requiring linguistic interpretation (Agarwal et al., 2024; Kim et al., 2016b).
+Visual cues is one of the most commonly used modalities in ingestive behavior interventions, providing intuitive, accessible, and scalable means of conveying prompts, progress, or state awareness.
+Based on representation format, we categorize visual cues into three submodalities:
+i.Textual: Text-based systems convey dietary information via notifications, reminders, or instructions.
+They often support user behavior by prompting actions, tracking progress, and aiding decisions.
+Examples include recommender agents (Zenun Franco, 2017; Arefeen et al., 2022), JIT assistants (Prasetyo et al., 2020; Carroll et al., 2013), and self-monitoring apps (Burke et al., 2011; Krishna et al., 2025).
+Many others use dashboards or push notifications to present progress and reinforce goals (Faltaous et al., 2021; Lee et al., 2019; Haerens et al., 2007).
+ii.Graphical: Graphical methods use visual elements, such as charts, bars, or data views, to convey behavioral patterns and intervention outcomes.
+These representations support self-awareness by illustrating nutritional balance, eating pace, or meal composition.
+For example, eat2pic transforms bite patterns into visual art (Nakamura et al., 2023), while others visualize ingestion summaries through app dashboards (Faltaous et al., 2021; Scisco et al., 2011; Lo et al., 2007).
+iii.Symbolic: Symbolic approaches convey prompts with simplified visual representations, such as color changes (Bobin et al., 2018), flashing lights (Matsui et al., 2018; Lessel et al., 2016), or emoji-style icons (Kim and Bae, 2018a), which communicate behavioral cues with minimal cognitive load.
+These signals are often embedded in utensils or ambient displays to indicate states like ”ready to eat” or ”slow down (Agarwal et al., 2024; Kim et al., 2016b).
 5.2.2. Auditory
 Auditory modalities primarily rely on simple sound output devices, most commonly speakers embedded in utensils, tableware, or ambient systems.
-These systems use auditory feedback to deliver symbolic or affective cues that influence perception, pacing, or user awareness during meals.
-For instance, enhanced chewing sounds or filtered auditory effects have been applied to modulate food perception and reduce eating speed through cross-modal interaction (Kleinberger et al., 2023).
-Interactive forks or tableware generate rhythmic or instructive sounds, such as audio riddles or feedback tones, to encourage mindful chewing or promote healthy eating habits in children (Kadomura et al., 2013b; Cai et al., 2024).
-Auditory reminders are also embedded in game-based or robot-assisted scenarios to regulate intake timing (Xie et al., 2023; Kadomura et al., 2014).
-A complete list of studies employing auditory approaches can be found in Table 8.
-Compared to the visual channel, auditory cues provide a lightweight, ambient feedback pathway that can be unobtrusively integrated into multisensory interventions.
+These systems deliver symbolic or affective auditory cues that shape perception, pacing, and awareness during meals.
+Such cues influence eating through cross-modal effects — for example, enhancing chewing sounds can help slow intake (Kleinberger et al., 2023).
+Interactive forks and tableware produce feedback tones or playful audio to encourage mindful eating, particularly in children (Kadomura et al., 2013b; Cai et al., 2024).
+Auditory prompts are also employed in game-based or robot-assisted contexts to guide intake timing (Xie et al., 2023; Kadomura et al., 2014).
+This modality provides a lightweight, ambient pathway that can be unobtrusively integrated into multisensory interventions.
 5.2.3. Haptic
-Haptic modality involves bodily interventions using physical signals such as vibration or shape deformation.
-Instead of relying solely on visual or auditory channels, this approach engages the user’s sensorimotor experience by delivering tactile cues, modifying utensil mechanics, or altering physical resistance to shape real-time food-related actions.
-Vibration: Most systems employ vibration motors or piezoelectric actuators.
-For instance, smart bottles vibrate when hydration targets are unmet (Faltaous et al., 2021; Varun et al., 2023), chopsticks trigger vibrations upon improper placement (Wang and Zhang, 2021), and wristbands or smartwatches vibrate in response to fast eating or insufficient chewing (Kim et al., 2016b, a; Kim and Bae, 2018b; Kamachi et al., 2023).
-Some systems provide vibration cues for journaling reminders (Ye et al., 2016), while others use mobile phone vibrations as hydration prompts (Chiu et al., 2009).
-In utensil-based designs, vibrotactile feedback is provided when eating speed exceeds a threshold (Hermans et al., 2017).
-Overall, vibration offers an embodied, non-disruptive alternative to visual or auditory alerts.
-Deformation: Deformation-based feedback or intervention modifies the physical properties of utensils or wearable devices to guide ingestive behavior.
-These systems typically implement adjustable deformation through pneumatic or motorized actuation to enable conscious or subconscious intervention during eating.
-For example, inflatable wearables like the FatBelt simulate abdominal fullness after caloric thresholds to promote self-awareness (Pels et al., 2014), while utensil designs such as the SSpoon (Chen et al., 2022), SWAN (Khot et al., 2020), a pneumatically bending spoon (Zhang et al., 2019), or a motor-stabilized spoon for Parkinson’s patients (Singhal et al., 2024) deform or adjust posture to hinder rapid intake or support stable consumption.
-In contrast, other approaches rely on materials with inherent deformability, such as VR rubber donuts (Li and Bailenson, 2018) or rubbing scent-releasing tactile surfaces (Cai et al., 2024).
+Haptic modality involves bodily interventions using physical signals such as vibration or deformation.
+Rather than relying solely on visual or auditory channels, it engages sensorimotor experience by delivering tactile cues, modifying utensil mechanics, or altering physical resistance to shape real-time food-related actions.
+i.Vibration: Most systems use vibration motors or piezoelectric actuators embedded in utensils, wearables, or phones, to offer a non-disruptive alternative to visual or auditory cues.
+Utensil-based systems include bottles that vibrate to prompt hydration (Faltaous et al., 2021; Varun et al., 2023), chopsticks for posture correction (Wang and Zhang, 2021), and cutlery for eating speed regulation (Hermans et al., 2017).
+Wearables like wristbands or smartwatches deliver vibrotactile intervention for fast eating or poor chewing (Kim et al., 2016b, a; Kim and Bae, 2018b; Kamachi et al., 2023).
+Some systems also use phone vibrations to prompt hydration (Chiu et al., 2009) or journaling (Ye et al., 2016).
+ii.Deformation: Deformative methods typically modify physical properties of devices via pneumatic or motorized actuation to influence ingestive behavior.
+Utensil designs like SSpoon (Chen et al., 2022), SWAN (Khot et al., 2020), and other deformable spoons (Zhang et al., 2019; Singhal et al., 2024) adjust posture or resistance to slow intake or enhance control.
+Wearables such as FatBelt simulate abdominal fullness through inflation after caloric thresholds (Pels et al., 2014).
+Other systems use deformable materials, such as VR rubber donuts (Li and Bailenson, 2018) or scent-releasing tactile surfaces (Cai et al., 2024), to influence ingestive behavior.
 5.2.4. Olfactory
-Olfactory modality utilizes scent stimuli to influence perception, modulate appetite, or reinforce associations with food-related behaviors.
-All studies related to this modality adopt the strategy of intentionally delivering odor cues to induce behavioral or perceptual changes.
-Methods range from digitally coordinated aroma delivery to embedded scent-triggering materials.
-Li and Bailenson (2018) used aromatic oil diffusers synchronized with VR scenes to emit food-related smells, enhancing satiation and reducing intake.
-Narumi et al.
-(2011) developed an olfactory display that filtered or overlaid real food odors to shift perception and promote healthier choices.
-Cai et al.
-(2024) embedded scent-infused ink in puzzle pieces that released odors through friction to support children’s food recognition.
-Collectively, these interventions illustrate the potential of olfactory feedback to modulate user experience on both affective and cognitive levels within ingestion-related contexts.
-Although relatively underexplored, scent-based strategies offer a promising channel for multisensory engagement, particularly when integrated with complementary modalities such as visual or tactile input.
+Olfactory modality uses scent stimuli to influence perception, appetite, or food-related associations, typically through intentional odor delivery.
+Methods range from digital aroma delivery to embedded scent materials, including VR-synced diffusers to curb intake (Li and Bailenson, 2018), odor filters to shift food perception (Narumi et al., 2011), and scent puzzles to aid food recognition in children (Cai et al., 2024).
+These approaches highlight the affective and cognitive potential of olfactory cues, especially when combined with visual or tactile input.
 5.2.5. Physiological
-Among physiological modalities, electrodermal stimulation is the only technique applied in ingestion-related interventions.
-These systems use intraoral electrical pulses to modulate muscle activity in a closed-loop manner, responding to physiological events.
+Among physiological modalities, electrodermal stimulation is the only technique used in ingestion interventions, delivering intraoral electrical pulses to modulate muscle activity.
 Zheng et al.
-(2025) delivered synchronized stimulation to the mandibular vestibule based on mastication force, enhancing tactile feedback and chewing strength.
-Biyani et al.
-(2019) used tongue pressure to trigger soft palate stimulation during swallowing, aiming to reduce pharyngeal transit time and improve coordination.
-These sensorimotor interventions directly target neuromuscular regulation rather than perceptual awareness.
-Although rarely explored, such systems offer a promising path for low-latency modulation of ingestion behavior.
+(2025) stimulated the mandibular vestibule based on chewing force, while Biyani et al.
+(2019) triggered soft palate stimulation via tongue pressure to enhance swallowing.
+Though rarely explored, such sensorimotor methods target neuromuscular control and show promise for low-latency modulation of ingestion behavior.
 5.3. Design Space Mapping through Modality to Behavior
-To further characterize the distribution and coverage of closed-loop approaches for sensing and intervention in ingestion health, we conducted a morphological analysis combining the dimensions of sensing and feedback/intervention modalities with their corresponding target behaviors, inspired by Zwicky (1967) and Jansen et al.
+To further characterize the distribution and coverage of closed-loop paradigms for sensing and intervention in ingestion health, we conducted a morphological analysis combining the dimensions of sensing and intervention modalities with their corresponding target behaviors, inspired by Zwicky (1967) and Jansen et al.
 (2022).
-This results in two-dimensional design matrices, shown in Table 3 and Table 4, which map the occurrence of each modality (x-axis) against specific behavioral targets (y-axis).
-To reduce redundancy and enhance interpretability, we adopt a mixed-level classification along the modality axis: for modalities with numerous fine-grained subcategories and overlapping device implementations, we collapse them into higher-level groupings to better reflect functional similarity and design relevance.
-The matrices enable us to identify well-covered modality–behavior pairs, underexplored combinations, and potential design opportunities.
-We hope this visualization can serve as a generative framework to inspire new behavior-aware intervention designs in ingestion health.
-Analysis of Table 3 reveals several key trends in the current design landscape in sensing process.
-Human-based motion sensing, particularly wrist and object trajectory tracking, demonstrates broad behavioral coverage in oral and finger motor activities and some other context factors.
-Visual and acoustic modalities are also well represented, especially for food recognition, intake timing, and ingestion pace, often leveraging image, sound, or AR-based inputs.
-By contrast, environment-based modalities like electrical or magnetic sensing remain underutilized, appearing in only a few cases.
-Physiological sensing (e.g., EMG, EDA, CGM) offers rich internal state signals but is mostly limited to specific domains such as emotional state or metabolic monitoring.
-This suggests potential in integrating these modalities for broader behavioral targets.
+This results in two-dimensional design matrices, shown in Table 7 and Table 8, mapping each modality (x-axis) to specific behavioral targets (y-axis).
+To reduce redundancy and improve clarity, we adopt a mixed-level classification along the modality axis, collapsing fine-grained subcategories with overlapping implementations into higher-level groupings reflecting functional similarity.
+The matrices enable us to identify well-covered modality–behavior pairs, underexplored combinations, and design opportunities.
+This visualization can serve as a framework to inspire new behavior-aware interventions in ingestion health.
+Analysis of Table 7 reveals trends in current sensing designs.
+Human-based motion sensing, especially wrist and object trajectory tracking, demonstrates broad behavioral coverage of oral and finger motor activities .
+Visual and acoustic modalities are also common for food recognition, intake timing, and pacing, often using image, sound, or AR inputs.
+In contrast, environment-based modalities like electrical or magnetic sensing remain underutilized.
+Physiological sensing (e.g., EMG, EDA, CGM) provides rich internal state signals but is mainly limited to emotional or metabolic monitoring, suggesting potential for broader integration.
 The matrix also highlights underexplored areas, including food hygiene, emotional triggers, and environmental dynamics.
@@
 Overall, the mapping reveals both saturated modality–behavior pairs and areas where new sensing strategies can enhance behavior-aware intervention design.
-Table 3. Mapping of Sensing Modalities to Target Factors and Behaviors in Ingestion Health
-Sensing Modality
-Human-based Environment-based
-Physiological Electrical
-Visual Auditory Haptic Motion Cardiac Metabolic Autonomic Visual Acoustic Force Spectral Conductivity Magnetic
-Food Choices facial expression (Robinson et al., 2020; Triantafyllidis et al., 2019), gaze (VR) (Karkar et al., 2018) speech (Prasetyo et al., 2020; Arefeen et al., 2022; Luo et al., 2021; Robinson et al., 2020), intake sound (Bi et al., 2016; Maramis et al., 2014; Kalantarian et al., 2014) touch (Cai et al., 2024; Robinson et al., 2020), oral behavior (Abisha and Rajalakshmy, 2017), body weight (Harous et al., 2017; Ludwig et al., 2011; Bastida et al., 2023; Nedungadi et al., 2018) wrist trajectory (Faltaous et al., 2021; Dragoni et al., 2017; Maramis et al., 2014), Object trajectory (Kadomura et al., 2013b; Nakamura et al., 2023; Joi et al., 2016; Kadomura et al., 2014, 2013a; Bedri et al., 2022), exercise (Kumbhare et al., 2023; Mishra et al., 2024; Harous et al., 2017; Dragoni et al., 2017; Maramis et al., 2014; Bastida et al., 2023; Chukwu, 2011; Badawi et al., 2012a; Fuchs et al., 2019; Tirasirichai et al., 2018; Jung et al., 2007; Vazquez et al., 2015) blood pressure (Huang et al., 2018; Jung et al., 2007; Ludwig et al., 2011; Park and Min, 2007; Nedungadi et al., 2018), heart rate (ECG) (Jung et al., 2007; Park and Min, 2007; Badawi et al., 2012b; Nedungadi et al., 2018), blood oxygen (Park and Min, 2007; Nedungadi et al., 2018) CGM (Liang, 2024; Arefeen et al., 2024; Nedungadi et al., 2018), BIA (Balbin et al., 2020; Chukwu, 2011), cholesterol (Harous et al., 2017) temperature (Park and Min, 2007; Nedungadi et al., 2018; Jung et al., 2007) food picture (Faltaous et al., 2021; Nakamura et al., 2023; Cai et al., 2024; Kadomura et al., 2014, 2013a; Ganesh et al., 2014; Arefeen et al., 2022; YR et al., 2024; Mishra et al., 2024; Alloghani et al., 2016; Hsiao et al., 2011, 2011; Achilleos et al., 2017; Kanjalkar et al., 2024; Kwon et al., 2016; Eigen et al., 2018; Cordeiro et al., 2015; Waltner et al., 2015; Epstein et al., 2016; Akter et al., 2021), Nutrition Label (Hsiao et al., 2011; Fernandez et al., 2013; Gutiérrez et al., 2018; Fuchs et al., 2019; Volkova et al., 2016; Eyles et al., 2017; Mankoff et al., 2002; Vazquez et al., 2015) food removal (Liu et al., 2024) Meals and plate weight (de Vries et al., 2020; Joi et al., 2016; Jung et al., 2017; Zhao et al., 2021; Ritschel et al., 2018; Krishna et al., 2025; Maramis et al., 2014; Kim et al., 2016c) alcohol (Matsui et al., 2018), temperature (Bastida et al., 2023; Kwon et al., 2016; Badawi et al., 2012b), object state (Ritschel et al., 2018), food distance (Bedri et al., 2022) humidity (Bastida et al., 2023; Kwon et al., 2016; Badawi et al., 2012b)
-Intake Timing intake sound (Bi et al., 2016) button click (Fortmann et al., 2014) wrist gesture (Ye et al., 2016) CGM (Liang, 2024) heart rate (Varun et al., 2023) liquid level (Varun et al., 2023) food weight (Lessel et al., 2016) temperature (Varun et al., 2023) humidity (Varun et al., 2023)
-Environment indoor position (Faltaous et al., 2021)
-Context Factors Activity Around Ingestion postprandial exercise (Alexander et al., 2017)
-Intake Quantity body gesture (Xie et al., 2023) oral behavior (Farooq et al., 2017) bite count (Scisco et al., 2011), object trajectory (Zhou et al., 2021; Li and Bailenson, 2018) food picture (Chiu et al., 2009) leftovers (Faltaous et al., 2021), liquid level (Varun et al., 2023; Tommy et al., 2017; Ravindran et al., 2022) liquid level (Lessel et al., 2016) liquid level (Poddar et al., 2024) liquid level (Kaner et al., 2018; Kreutzer et al., 2015; Ko et al., 2007; Pankajavalli et al., 2017; Wijanarko et al., 2019), alcohol (Bobin et al., 2018) liquid level (Yıldız and Coşkun, 2019)
-Ingestion Pace and Oral Motor Behavior facial expression (Kim and Bae, 2018a) intake sound (Kleinberger et al., 2023; Koizumi et al., 2011; Bi et al., 2016; Maramis et al., 2014; Kamachi et al., 2023; Chen and Yen, 2022) oral behavior (Kim et al., 2016b; Biyani et al., 2019; Zheng et al., 2025) wrist trajectory (Mendi et al., 2013; Lee et al., 2019; Scisco et al., 2011; Kim et al., 2016a; Kim and Bae, 2018b; Maramis et al., 2014; Turner-McGrievy et al., 2017), object trajectory (Kadomura et al., 2013b; Nakamura et al., 2023; Biyani et al., 2019; Hermans et al., 2017; Zhang et al., 2019), skin distance (Hori et al., 2023) wrist action (EMG) (Lee et al., 2019; Kim et al., 2016b; Sugita et al., 2018) food weight (de Vries et al., 2020; Krishna et al., 2025; Lo et al., 2007; Kim et al., 2016c; Ioakimidis et al., 2009; Mitchell et al., 2015; Hermans et al., 2017)
-Food Awareness or Emotional State intake sound (Kleinberger et al., 2023; Koizumi et al., 2011; Zhang and Parker, 2020), speech (Mayumi et al., 2023; Parra et al., 2023; Zhang and Parker, 2020) wrist trajectory (Wang et al., 2025), object trajectory (Khot et al., 2020) emotion state (EDA) (Carroll et al., 2013) food picture (AR) (Narumi et al., 2011) food weight (Mitchell et al., 2015) physical contact (Han and Kang, 2017)
-Target Factors and Behaviors Process Behaviors Safe Eating Practices hygiene (Wang and Zhang, 2021) temperature (Singhal et al., 2024)
-Analysis of Table 4 reveals key trends in feedback modality distribution.
-Textual and graphical visual feedback dominate, especially for food choices, intake quantity, and ingestion pace.
-These channels are favored for their accessibility and clarity.
-Symbolic feedback appears less frequently but with growing diversity.
+Analysis of Table 8 reveals trends in intervention modality distribution.
+Textual and graphical visual intervention dominate, especially for food choices, intake quantity, and ingestion pace, favored for their accessibility and clarity.
+Symbolic intervention appears less frequently but with increasing diversity.
 LED indicators are used across hydration, eating pace, and safety contexts.
-Symbolic auditory feedback—such as chewing sounds or metaphorical cues—emerges as a lightweight and engaging method for promoting awareness.
+Symbolic auditory cues, such as chewing sounds or metaphorical signals, serve as a lightweight and engaging method to promote awareness.
 Tactile symbolic forms, like thermochromic or pneumatic responses, remain niche but offer ambient and intuitive signaling.
-Auditory feedback via speakers and alarms spans nearly all behavior types, while haptic feedback, mostly motor-based, is concentrated in emotional regulation and pace control.
-More advanced haptics (e.g., piezoelectric, pneumatic) remain underutilized.
-Experimental modalities like olfactory and physiological feedback appear in a few behaviorally rich cases, offering unique multisensory pathways.
-Overall, the table highlights the dominance of visual modalities but also points to symbolic and sensory alternatives as promising design directions.
-Gaps in environmental, emotional, and hygiene-related feedback suggest potential for broader and more creative feedback strategies in closed-loop ingestion health systems.
+Auditory interventions via speakers and alarms spans nearly all behavior types, while haptic interventions, mainly motor-based, focus on emotional regulation and pace control.
+More advanced haptics (e.g., piezoelectric, pneumatic) remain underused.
+Experimental modalities like olfactory and physiological interventions appear in a few behaviorally rich cases, offering unique multisensory pathways.
+Overall, the table highlights the dominance of visual modalities but also suggests symbolic and sensory alternatives as promising design directions.
+Gaps in environmental, emotional, and hygiene-related cues indicate potential for broader and more creative intervention strategies in closed-loop ingestion health systems.
 Notably, despite the variety of sensing pathways, few systems reuse the same sensors for both perception and intervention.
-Most existing feedback channels rely heavily on smartphones or static screens, rather than leveraging interactive or perceptual sensors for embodied or context-aware interventions.
-This reflects a substantial unexploited design space, which we further discuss and explore in the following Chapter 7.
-Table 4. Mapping of feedback/intervention Modalities to Target Factors and Behaviors in Ingestion Health
-Feedback/Intervention Modality
-Visual
-Textual Graphical Symbolic Auditory Haptic Olfactory Physiological
-Food Choices (total 62 pubs) SMS, app/screen message, AR/MR message (Gutiérrez et al., 2018; Fuchs et al., 2019) (total 26 pubs) app/screen graph, AR/MR graph (Ganesh et al., 2014; Gutiérrez et al., 2018; Fuchs et al., 2019) LED (de Vries et al., 2020; Matsui et al., 2018; Ganesh et al., 2014; Kim et al., 2016c; Ritschel et al., 2018; Triantafyllidis et al., 2019) microphone speaker (Kadomura et al., 2013b; Cai et al., 2024; Kadomura et al., 2014; Robinson et al., 2021; Kadomura et al., 2013a; Zhao et al., 2021; De Carolis et al., 2024; Liu et al., 2024; Ritschel et al., 2018; Chukwu, 2011; Terzioğlu et al., 2023; Triantafyllidis et al., 2019; Agarwal et al., 2024; Luo et al., 2021; Robinson et al., 2020) piezoelectric vibration (Faltaous et al., 2021), Aerodynamic deformation (Pels et al., 2014), friction (Cai et al., 2024) Scented liquid (Cai et al., 2024)
-Intake Timing SMS (Bi et al., 2016; Varun et al., 2023), app/screen message (Liang, 2024; Faltaous et al., 2021; Ye et al., 2016) app/screen graph (Liang, 2024; Faltaous et al., 2021) pneumatic membrane (Xie et al., 2023), LED (Fortmann et al., 2014) microphone speaker (Xie et al., 2023), alarm (Varun et al., 2023) piezoelectric vibration (Faltaous et al., 2021), motor vibration (Varun et al., 2023; Ye et al., 2016) LED (Lessel et al., 2016; Ravindran et al., 2022)
-Environment microphone speaker (Faltaous et al., 2021)
-Context Factors Activity Around Ingestion SMS (Alexander et al., 2017)
-Intake Quantity SMS (Poddar et al., 2024; Kreutzer et al., 2015; Moses et al., 2023; Pankajavalli et al., 2017; Wijanarko et al., 2019), app/screen message (Faltaous et al., 2021; Chaudhry et al., 2016; Chiu et al., 2009; Alshurafa et al., 2015; Maramis et al., 2014; Varun et al., 2023; Tommy et al., 2017), LED message (Ravindran et al., 2022) app/screen graph (Faltaous et al., 2021; Chaudhry et al., 2016; Scisco et al., 2011; Maramis et al., 2014; Varun et al., 2023; Chiu et al., 2009; Ko et al., 2007; Wijanarko et al., 2019; Tommy et al., 2017), projector (Yıldız and Coşkun, 2019) pneumatic membrane (Xie et al., 2023), thermo -chromic (Kaner et al., 2018), LED (Bobin et al., 2018; Lessel et al., 2016; Ravindran et al., 2022), living-media (Zhou et al., 2021) microphone speaker (Xie et al., 2023; Farooq et al., 2017), alarm (Varun et al., 2023; Tommy et al., 2017), symbolic auditory (Zhou et al., 2021) piezoelectric vibration (Faltaous et al., 2021), motor vibration (Li and Bailenson, 2018; Chiu et al., 2009; Varun et al., 2023), pneumatic Deformation (Chen et al., 2022) essential oil (Li and Bailenson, 2018)
-Ingestion Pace and Oral Motor Behavior SMS (Lee et al., 2019; Mendi et al., 2013; Krishna et al., 2025; Moses et al., 2023), app/screen message (Bi et al., 2016; Kim and Bae, 2018b; Maramis et al., 2014; Kamachi et al., 2023; Turner-McGrievy et al., 2017; Hori et al., 2023; Lo et al., 2007; Ioakimidis et al., 2009) app/screen graph (Nakamura et al., 2023; Scisco et al., 2011; Kim and Bae, 2018b; Maramis et al., 2014; Turner-McGrievy et al., 2017; Hori et al., 2023; Lo et al., 2007), AR/MR graph (Kim and Bae, 2018a; Sugita et al., 2018) LED (de Vries et al., 2020; Kim et al., 2016b, a; Hermans et al., 2017; Kim et al., 2016c) chewing sound feedback (Kleinberger et al., 2023; Chen and Yen, 2022), microphone speaker (Ioakimidis et al., 2009; Sugita et al., 2018; Mendi et al., 2013; Koizumi et al., 2011; Kadomura et al., 2013b), physical symbolic feedback (Mitchell et al., 2015) piezoelectric vibration (Faltaous et al., 2021; Zhang et al., 2019), motor vibration (Kim et al., 2016b, a; Kim and Bae, 2018b; Kamachi et al., 2023; Hermans et al., 2017) Intraoral electrical stimulation (Biyani et al., 2019; Zheng et al., 2025)
-Food Awareness or Emotional State SMS(Carroll et al., 2013), app/screen message (Mayumi et al., 2023; Sun et al., 2020; Parra et al., 2023; Wang et al., 2025) app/screen graph(Sun et al., 2020; Parra et al., 2023; Zhang and Parker, 2020), AR/MR graph (Narumi et al., 2011) LED (Han and Kang, 2017; Khot et al., 2020) microphone speaker (Kadomura et al., 2013b; Koizumi et al., 2011; Han and Kang, 2017; Mayumi et al., 2023; Parra et al., 2023), chewing sound feedback (Kleinberger et al., 2023; Zhang and Parker, 2020), physical symbolic feedback (Mitchell et al., 2015) motor vibration (Khot et al., 2020) odor filter (Narumi et al., 2011)
-Target Factors and Behaviors Process Behaviors Safe Eating Practices motor vibration (Wang and Zhang, 2021; Singhal et al., 2024)
+Most existing intervention channels still rely on smartphones or static screens rather than leveraging interactive or perceptual sensors for embodied or context-aware interventions.
+This reflects a substantial unexploited design space, which we will discuss further in Chapter 7, including specific hypotheses about effective sensing–intervention combinations generated from the design space.
 6. Effectiveness and User Acceptance of Current Ingestion Health Intervention
 In this section, we examine papers that include user studies to assess the effectiveness and user acceptance of their work.
-To systematically understand their results, we have categorized these papers into four types: behavioral feedback/intervention, educational feedback/intervention, social feedback/intervention, and other specific feedback/intervention.
-Furthermore, we have divided each type into three paradigms: those that include feedback only, those that include intervention only, and those that include both feedback and intervention.
+To systematically understand their results, we have categorized these papers into four types: behavioral intervention, educational intervention, social intervention, and other specific intervention.
+Furthermore, we have divided each type into three paradigms: passive intervention, active intervention, and combined intervention which includes both passive and active intervention.
 In total, 106 papers are included.
-6.1. Behavioral Feedback/Intervention
-Behavioral feedback/intervention refers to approaches that primarily focus on altering specific ingestive behaviors of users, such as eating speed, chewing counts, and drinking frequency, ultimately assisting them in achieving healthier eating habits.
-The technical approaches in this category primarily rely on sensor-based technologies (e.g., IMU, piezoelectric sensors, RFID) and real-time feedback mechanisms (e.g., visual, haptic, and auditory cues).
-6.1.1. Behavioral Feedback Only
-Behavioral feedback refers to the presentation of real-time or retrospective data that directly reflects a user’s actions, enabling self-awareness and self-regulation without explicit instructions or forced behavioral changes.
-It objectively mirrors user behavior, allowing individuals to draw their own conclusions and adjust their habits autonomously.
-Systems that provide behavioral feedback only — such as continuous glucose monitoring (Liang, 2024) and smart hydration bracelets (Fortmann et al., 2014) — offer unobtrusive data visualization without explicit instructions, relying on users’ self-reflection to drive behavioral changes.
-Similarly, ambient logging devices (Tang et al., 2015) and smart nutrition trackers (Alshurafa et al., 2015) collect and display consumption patterns through minimal interfaces, facilitating autonomous behavior adjustment.
+6.1. Behavioral Intervention
+Behavioral intervention refers to approaches that primarily focus on altering specific ingestive behaviors of users, such as eating speed, chewing counts, and drinking frequency.
+The technical approaches in this category primarily rely on sensor-based technologies and real-time feedback mechanisms.
+6.1.1. Passive Behavioral Intervention
+Passive behavioral intervention presents real-time or retrospective data reflecting a user’s actions, fostering self-awareness and autonomous behavior change without direct instruction.
+Systems that provide passive behavioral intervention — such as continuous glucose monitoring (Liang, 2024) and smart hydration bracelets (Fortmann et al., 2014) — offer unobtrusive data visualization without explicit instructions.
+Also, ambient logging devices (Tang et al., 2015) and smart nutrition trackers (Alshurafa et al., 2015) collect and display consumption patterns through minimal interfaces, facilitating autonomous behavior adjustment.
 A total of 20 papers are included in this category.
 Feedback effectiveness is evidenced by measurable behavioral changes.
-For instance, real-time glucose feedback (Liang, 2024) reduced average glucose levels in 50% of participants and prompted adjustments in meal timing, aligning with the ”action” stage of the transtheoretical model.
-Bite-rate feedback (Kim et al., 2016b; Scisco et al., 2011) led to a 57.7% reduction in energy intake among fast eaters, although the effects varied based on baseline consumption habits.
-Isomorphic feedback, such as the waist inflation from FatBelt (Pels et al., 2014) and ambient displays (Kaner et al., 2018), utilized physical or visual cues to reduce calorie intake and increase water consumption by 20–30%, respectively.
-Notably, the modality of feedback matters: younger users (ages 20s) preferred vibration feedback, while older adults (ages 50–70) favored light-based feedback (Kim et al., 2016b).
+For instance, real-time glucose feedback (Liang, 2024) reduced average glucose levels in 50% of participants and prompted adjustments in meal timing.
+Bite-rate feedback (Kim et al., 2016b; Scisco et al., 2011) led to a 57.7% reduction in energy intake among fast eaters.
+Isomorphic feedback, such as the waist inflation from FatBelt (Pels et al., 2014) and ambient displays (Kaner et al., 2018), utilized physical or visual cues to reduce calorie intake and increase water consumption by 20–30% respectively.
+Notably, the modality of passive intervention matters: younger users preferred vibration feedback, while older adults favored light-based feedback (Kim et al., 2016b).
 Additionally, manual input requirements, such as the time-interval logging in the SAL logger (Tang et al., 2015), can reduce adherence, particularly among older adults.
-The success of behavioral feedback presumably hinges on balancing autonomy (minimal manual input) and personalization (age-appropriate feedback and aesthetic adaptability).
-Devices that integrate seamlessly into daily routines typically outperform those that require active engagement.
-Behavioral feedback demonstrates consistent patterns in user acceptance and effectiveness across multiple studies.
+Passive behavioral intervention demonstrates consistent patterns in user acceptance across multiple studies.
+Passive behavioral intervention succeeds by balancing autonomy and personalization, with its seamless daily integration outperforming those requiring active engagement.
 For instance, wrist-worn devices, such as the WaterJewel bracelet (Fortmann et al., 2014) and the IDEA wristband (Lee et al., 2019), achieve a long-term usage willingness of 70–83% due to their non-intrusive design and plug-and-play functionality.
@@
 Additionally, aesthetic customization, such as WaterJewel’s design catering to both male and female users (Fortmann et al., 2014) and the GROW bottle’s thermo-chromic display (Kaner et al., 2018), further enhances acceptance by aligning with personal preferences.
-6.1.2. Behavioral Intervention Only
-Behavioral interventions actively modify ingestive behaviors through direct mechanisms, such as physical constraints or real-time prompts.
+6.1.2. Active Behavioral Intervention
+Active behavioral intervention actively modifies ingestive behaviors through direct mechanisms, such as physical constraints or real-time prompts.
 Physical constraint systems, like the pneumatic shape-changing spoon (Chen et al., 2022), mechanically regulate portion sizes and chewing rates by modifying the tangible interface.
@@
 More immersive interventions, including the sensory interactive table (de Vries et al., 2020) and the smart tray system (Jung et al., 2017), combine environmental sensing with real-time behavioral prompts to actively reshape dining interactions.
-Unlike passive feedback that merely presents data, these interventions enforce immediate behavioral changes by altering how users interact with food.
 A total of 15 papers are included in this category.
-Behavioral intervention systems actively reshape user actions through physical, computational, or gamified means.
-The mandatory approach proves particularly effective for users needing structured support, as it bypasses reliance on self-motivation to drive habitual change.
+Active behavioral intervention systems actively reshape user actions through physical, computational, or gamified means.
 Interventions that physically alter user behavior demonstrate particularly strong efficacy.
@@
 Additionally, intraoral closed-loop systems increased mastication force by 26% through real-time electrical stimulation (Biyani et al., 2019) and improved swallowing duration via neuromuscular triggers (Zheng et al., 2025).
-Structured guidance approaches have also proven effective; for example, the DIMA app enhanced portion estimation accuracy across various literacy levels (Chaudhry et al., 2016), and computer-tailored interventions achieved a twofold greater reduction in fat intake compared to improvements in physical activity through enforced rules (Vandelanotte et al., 2005).
+Mandatory approaches effectively provide structured support by circumventing reliance on self-motivation, just as structured guidance methods do; for example, the DIMA app enhanced portion estimation accuracy across various literacy levels (Chaudhry et al., 2016), and computer-tailored interventions achieved a twofold greater reduction in fat intake compared to improvements in physical activity through enforced rules (Vandelanotte et al., 2005).
 Furthermore, WaterCoaster increased drinking frequency from 2–3 to 4–6 times daily by combining personalized hydration goals with virtual rewards (Lessel et al., 2016).
-Regarding user acceptance of behavioral interventions, physical manipulation systems like the SSpoon (Chen et al., 2022) achieved high acceptance by subtly altering ingestive behavior without disrupting appetite.
+Regarding user acceptance of active behavioral interventions, physical manipulation systems like the SSpoon (Chen et al., 2022) achieved high acceptance by subtly altering ingestive behavior without disrupting appetite.
 Structured guidance approaches, such as computer-tailored interventions, revealed user segmentation; simultaneous interventions were more effective for non-compliant users, while sequential approaches suited compliant individuals (Vandelanotte et al., 2005).
@@
 For instance, WaterCoaster’s gamified approach engaged 70.6% users (Lessel et al., 2016), whereas the complex gamification of the Sensory Interactive Table likely led to user fatigue (de Vries et al., 2020).
-6.1.3. Behavioral Feedback with Intervention
-Behavioral feedback with intervention combines the features of real-time data representation and mandatory guidance.
-It typically takes the form of comprehensive health system applications, such as the smart home-based IoT nutritional intake monitoring system (Faltaous et al., 2021), which provides unobtrusive health intervention advice through a smartphone application along with symbolic feedback mechanisms.
-This dual approach simultaneously offers users visualized insights while enforcing behavioral modifications through structured guidance.
+6.1.3. Combined Behavioral Intervention
+Combined behavioral intervention integrates the features of real-time data representation and mandatory guidance.
+It typically takes the form of comprehensive health system applications, such as the smart home-based IoT nutritional intake monitoring system (Faltaous et al., 2021), which simultaneously combines symbolic feedback with visualized insights and structured behavioral modification.
 A total of 7 papers are included in this category.
-Integrated systems that combine behavioral feedback with active interventions demonstrate enhanced effectiveness and user acceptance when they balance automation with personalization.
+Combined systems that integrate both passive and active interventions demonstrate enhanced effectiveness and user acceptance when they balance automation with personalization.
 For instance, a smart home-based nutritional monitoring system achieved over 50% habit adjustment through goal-aligned mobile reminders, although privacy concerns were noted by 66.6% users (Faltaous et al., 2021).
 Similarly, wristband-based eating speed systems reduced bite size with a detection accuracy of 96.7%, while maintaining high acceptance through haptic or visual feedback (Kim et al., 2016a).
-In clinical populations, the Parkinson’s smart spoon improved stability (X/Y-axis deviation ¡15% under simulated tremors) while integrating diagnostic alerts; however, its niche design limits broader adoption (Singhal et al., 2024).
+In clinical populations, the Parkinson’s smart spoon improved stability (X/Y-axis deviation ¡15%) while integrating diagnostic alerts; however, its niche design limits broader adoption (Singhal et al., 2024).
 Low-cost solutions, such as receipt-based nutrition tracking, proved feasible for promoting dietary awareness, particularly for monitoring calcium intake, but relied heavily on user compliance (Mankoff et al., 2002).
-6.2. Educational Feedback/Intervention
-Educational feedback/intervention refers to feedback or interventions that have educational significance or involve educational content.
+6.2. Educational Intervention
+Educational intervention refers to interventions that have educational significance or involve educational content for enhancing users’ food literacy.
 This includes food and recipe recommendations, nutrition education, educational gaming, and guidance for users to adopt healthy eating habits.
-Ultimately, these approaches aim to enhance food literacy among users through educational approach.
-6.2.1. Educational Feedback Only
-Educational feedback focuses exclusively on providing knowledge-based guidance to enhance learning outcomes, without direct instructional intervention prompts.
-It prioritizes pedagogical value over real-time behavioral modification and delivers structured information (e.g., nutritional facts, eating patterns) to cultivate cognitive awareness and self-reflection, empowering users to make informed dietary decisions.
-For example, a puzzle game can be utilized to enhance children’s food literacy, encouraging self-driven exploration of healthy eating patterns without explicit behavioral directives (Cai et al., 2024).
-Additionally, an acoustic sensor can detect ingestion and deliver real-time nutritional insights via a mobile application, thereby raising awareness of everyday dietary habits through unobtrusive monitoring (Bi et al., 2016).
-This type of feedback emphasizes knowledge dissemination over actual modification, leveraging playful interaction (Cai et al., 2024) or ambient data capture (Bi et al., 2016) to implicitly influence ingestive behaviors.
+6.2.1. Passive Educational Intervention
+Passive educational intervention delivers pedagogical and knowledge-based guidance to enhance learning, prioritizing cognitive awareness over behavioral modification to inform dietary decisions.
+For example, a puzzle game can be utilized to encouraging children’s self-driven exploration of healthy eating patterns without explicit behavioral directives (Cai et al., 2024).
+Additionally, an acoustic sensor can detect ingestion and deliver real-time nutritional insights via a mobile application (Bi et al., 2016).
+This type of intervention emphasizes knowledge dissemination over actual modification, leveraging playful interaction (Cai et al., 2024) or ambient data capture (Bi et al., 2016) to implicitly influence ingestive behaviors.
 As such, it positions itself as an educational tool for long-term nutrition self-regulation.
 A total of 11 papers are included in this category.
-Educational feedback demonstrates significant potential in improving food literacy and dietary behaviors, particularly when leveraging engaging, multi-sensory interfaces.
+Passive educational intervention demonstrates significant potential in improving food literacy and dietary behaviors, particularly when leveraging engaging interfaces.
 For instance, multi-sensory puzzle games increased children’s food literacy by 59.16% through tangible interactions that fostered cognitive connections with food (Cai et al., 2024).
@@
 In clinical populations, the SaltSwitch application (Eyles et al., 2017) successfully reduced salt purchases among patients with cardiovascular disease by providing real-time, barcode-based traffic light nutrition labels.
-Educational feedback excels when it reduces cognitive load (e.g., speech input in FoodScrap (Luo et al., 2021) compared to manual logging) and leverages intuitive interfaces (e.g., mixed reality labels versus traditional charts (Fuchs et al., 2019)).
+Passive educational intervention excels when it reduces cognitive load (e.g., speech input in FoodScrap (Luo et al., 2021) compared to manual logging) and leverages intuitive interfaces (e.g., mixed reality labels versus traditional charts (Fuchs et al., 2019)).
 It achieves higher user acceptance by prioritizing intuitive interactions and minimizing user burden.
-6.2.2. Educational Intervention Only
-Educational interventions emphasize the delivery of structured knowledge and instructional guidance through direct conversational interactions, without incorporating data-driven feedback mechanisms.
-Their design aligns with didactic principles, leveraging conversational interfaces to scaffold learning.
+6.2.2. Active Educational Intervention
+Active educational interventions emphasize the delivery of structured knowledge and instructional guidance through direct conversational interactions.
 This type of intervention is typically exemplified by dialogue-based chatbots, such as Foodbot (Prasetyo et al., 2020), and humanoid social robots (Robinson et al., 2021), which provide goal-oriented educational content (e.g., nutrition advice and dietary management) through scripted or AI-generated conversations.
@@
 Similarly, social robots outperformed chatbots in promoting healthier eating intentions among children, leveraging their physical presence to enhance persuasiveness (De Carolis et al., 2024).
-Interactive tableware, such as FunEat’s gamified plate (Zhao et al., 2021), improved children’s food choices through projected animations, although parents noted potential distractions during meals.
+Interactive tableware, such as FunEat’s gamified plate (Zhao et al., 2021), improved children’s food choices through projected animations.
 Other ubiquitous computing approaches, like sensor-embedded forks, achieved 89% child engagement by replacing abstract lessons with playful interactions (Kadomura et al., 2014).
 Personalized educational dietary interventions demonstrate strong effectiveness when incorporating adaptive algorithms.
-The SmartDiet system (Hsiao and Chang, 2010) generates nutritionally balanced meal plans within seconds while maintaining high user acceptance through interactive optimization.
-Linear optimization frameworks have achieved less than 1.2% calorie recommendation errors by dynamically adjusting to user deviations (Arefeen et al., 2022).
-Intelligent nutrition services further enhance adoption by combining image recognition with flexible service selection, although ecosystem factors did not significantly impact perceived educational-targeted behavioral control (Hsiao et al., 2011).
+SmartDiet (Hsiao and Chang, 2010) generates nutritionally-optimized meal plans in seconds via interactive optimization, and similar approaches achieves ¡1.2% calorie errors through dynamic user adaptation (Arefeen et al., 2022).
+Subsequent enhancements with image recognition also improved adoption, though ecosystem factors didn’t affect educational behavior control (Hsiao et al., 2011).
 Therefore, future systems should balance guidance intensity (Liu et al., 2024) with user autonomy (Ritschel et al., 2018), while leveraging AI optimization for sustainable personalization (Arefeen et al., 2022; Hsiao and Chang, 2010).
-6.2.3. Educational Feedback with Intervention
-Educational feedback and intervention combine knowledge-based data insights with structured guidance to create a comprehensive learning support system.
-This approach delivers real-time educational feedback (e.g., nutrition analytics) while providing direct instructional interventions (e.g., personalized study plans and goal-setting prompts) to reinforce desired behaviors through a standardized framework (Kadomura et al., 2013b).
-Gamification is also widely applied in such scenarios (Nakamura et al., 2023).
-This dual mechanism ensures that learners not only understand their performance metrics but also receive actionable, curriculum-aligned strategies to bridge knowledge-practice gaps, distinguishing it from standalone feedback or intervention approaches.
-Related studies demonstrate a convergence of real-time monitoring and pedagogical guidance in dietary feedback and interventions.
-These implementations collectively operate within the ”observe-educate-act” framework through embedded sensors and algorithmic personalization, although they vary in modality (tangible vs.
-conversational) and intervention immediacy (real-time vs.
-retrospective).
+6.2.3. Combined Educational Intervention
+Combined educational intervention merges data-driven insights with structured guidance, delivering real-time feedback alongside instructional support through standardized frameworks (Kadomura et al., 2013b), and are often enhanced by gamification (Nakamura et al., 2023).
+Related studies demonstrate a convergence of real-time monitoring and pedagogical guidance in dietary interventions.
+This dual approach bridges knowledge-practice gaps with actionable strategies, operating via sensor-based ’observe-educate-act’ cycles across varying modalities and intervention timings.
 A total of 13 papers are included in this category.
-Integrated systems that combine real-time feedback with structured educational interventions achieve behavioral change through multimodal engagement, expert-guided personalization, and interactive reinforcement.
+Integrated systems that add real-time feature into structured educational interventions achieve behavioral change through multimodal engagement, expert-guided personalization, and interactive reinforcement.
 Multimodal engagement proves particularly effective in pediatric contexts.
 For instance, systems like EducaTableware’s sound-emitting tableware (Kadomura et al., 2013b) and FoodWorks’ AR meal augmentations (Ganesh et al., 2014) increased children’s willingness to try new foods by 45-60% while reducing mealtime distractions.
-Meanwhile, eat2pic’s color-based food visualization significantly slowed eating speed (meal time/intervals) (Nakamura et al., 2023), although effectiveness varied with users’ artistic interests.
+Meanwhile, eat2pic’s color-based food visualization significantly slowed eating speed (meal time/intervals) (Nakamura et al., 2023).
 Nutrifit’s sensor-tracked meals improved the consumption of nutrient-dense foods by 3-8% through real-time weight monitoring and app feedback (Krishna et al., 2025).
-Expert-guided systems also demonstrate strong clinical validity.
-Therefore, if such educational systems combine both embedded sensing and adaptive guidance, they will maximize user acceptance and effective outcomes.
+Expert-guided systems also demonstrate strong clinical validity (Achilleos et al., 2017).
 Regarding user acceptance of this combined educational approach, the OCARIoT platform reduced childhood obesity prevalence by 75.5% through IoT-collected nutrition and activity data, along with coordination from healthcare professionals, achieving over 80% user acceptance (Bastida et al., 2023).
-Web-based health monitoring platforms maintained 95% daily user engagement through real-time expert feedback, although 85% of nutritionists preferred weekly reviews over daily ones (Achilleos et al., 2017).
-Additionally, non-invasive bioelectric devices like ”Keep and Eat” demonstrated 98.45% accuracy (±1.55% error) in body composition measurements, enabling precise dietary recommendations (Balbin et al., 2020).
-6.3. Social Feedback/Intervention
-Social feedback/intervention refers to feedback or interventions that involve more than one person, incorporating social support or the eating environment.
-This type of feedback/intervention leverages group dynamics and shared experiences to reinforce closed-loop paradigms through social norms.
-They often incorporate interactive elements, such as family meal tracking or community network challenges, to foster collective accountability.
-By embedding feedback and interventions in natural social contexts, they tap into intrinsic motivations for conformity and belonging that individual ingestion-related approaches cannot address.
-6.3.1. Social Feedback Only
-Social feedback creates shared accountability through real-time group comparisons (e.g., displaying community members’ vegetable intake rates) or peer endorsements.
-It leverages social norms and observational learning—showing how similar individuals achieve balanced diets to motivate self-adjustment without prescriptive rules.
-This type of feedback makes healthy choices feel collectively validated rather than individually enforced.
+Web-based health monitoring platforms maintained 95% daily user engagement through real-time expert feedback (Achilleos et al., 2017).
+Additionally, non-invasive bioelectric devices like ”Keep and Eat” demonstrated 98.45% accuracy in body composition measurements, enabling precise dietary recommendations (Balbin et al., 2020).
+6.3. Social Intervention
+Social intervention refers to behavioral strategies that leverage group dynamics and social contexts (e.g., family meals or community networks) to drive change and reinforce the closed-loop paradigms.
+By incorporating shared experiences and social norms, this approach fosters collective accountability and taps into intrinsic motivations for belonging — aspects that individual-focused interventions cannot address.
+6.3.1. Passive Social Intervention
+Passive social intervention fosters accountability through real-time peer comparisons, leveraging social norms and observational learning to motivate self-regulation while making healthy choices feel socially validated.
 A total of 2 papers are included in this category.
-HealthyStadium (Eigen et al., 2018) and Crumbs (Epstein et al., 2016) utilize social feedback—through meal photo rankings and shared challenges—to encourage healthier eating habits more effectively than solo tracking.
-Both systems demonstrate how passive peer interaction fosters dietary mindfulness and self-adjustment.
-Their designs intentionally avoid direct intervention, instead relying on social dynamics such as normative comparison and gentle competition to organically shift ingestive behaviors.
-Both systems demonstrate the effectiveness of social feedback in promoting healthier ingestive habits through lightweight, non-intrusive engagement rather than active intervention.
-Critically, this lightweight, feedback-driven approach has proven to be more acceptable than traditional interventions, as users tend to resist rigid rules but embrace socially reinforced habits.
-Users in HealthyStadium exhibited higher engagement by striving for better health ratings, with participants consciously opting for healthier meals due to peer evaluations.
-One user noted, “When I was wondering whether to have a delicious meal or a healthy meal, I chose a healthy one” (Eigen et al., 2018).
+HealthyStadium (Eigen et al., 2018) and Crumbs (Epstein et al., 2016) utilize passive social intervention — through meal photo rankings and shared challenges — to encourage healthier eating habits more effectively than solo tracking.
+Both systems validate that passive social intervention cultivates dietary self-regulation through non-intrusive peer engagement, rather than active interventions.
+Critically, this lightweight, feedback-driven approach has proven to be more acceptable than traditional interventions, as users tend to embrace socially reinforced habits.
+Users in HealthyStadium exhibited higher engagement by striving for better health ratings, with participants consciously opting for healthier meals due to peer evaluations (Eigen et al., 2018).
 Meanwhile, Crumbs leveraged daily challenges and social interaction, with 72% of users enjoying the challenges and 46% expressing intent to continue using the platform.
-Additionally, the social group reported motivational effects, such as ”not wanting to fall behind others” (Epstein et al., 2016).
-6.3.2. Social Intervention Only
-Social intervention for healthy ingestion actively leverages interpersonal dynamics to reshape dietary behaviors through actual engagement.
-This type of intervention employs real-time social triggers—such as parent-child tactile feedback—that directly mediate ingestive actions within relational contexts.
+Moreover, the social group reported motivational effects, such as ”not wanting to fall behind others(Epstein et al., 2016).
+6.3.2. Active Social Intervention
+Active social intervention for healthy ingestion leverages interpersonal dynamics to reshape dietary behaviors through actual engagement.
+This type of intervention employs real-time social triggers that directly mediate ingestive actions within relational contexts.
 By embedding behavioral guidance within social interactions (e.g., family meals and shared environments), these interventions transform abstract nutritional awareness into contextually anchored behavior change.
@@
 For older groups, auto-generated nutritional postcards link eating patterns with emotional associations (Sun et al., 2020), offering a subtle behavioral nudge.
-The interactive tableware system (Joi et al., 2016) successfully encouraged vegetable consumption in children aged 3–4, with observed behavioral shifts, such as a 4-year-old transitioning from being spoon-fed to self-feeding.
-Additionally, it improved parent-child interaction through gamified learning.
-Similarly, a middle-school healthy eating intervention demonstrated gender-dependent effects, notably reducing fat intake in girls—particularly when parents were involved—though it had a limited impact on fruit and water intake as well as soft drink consumption (Haerens et al., 2007).
-Meanwhile, the auto-generated postcard system facilitated self-reflection, leading to expressed intent to modify behaviors (Sun et al., 2020).
+The interactive tableware system (Joi et al., 2016) increased vegetable consumption in 3-4-year-olds, demonstrating behavioral shifts like transitioning from spoon-feeding to self-feeding while enhancing parent-child interactions through gamification.
+Similarly, a school-based intervention (Haerens et al., 2007) showed gender-specific effects, significantly reducing girls’ fat intake.
+Complementing these, the automated postcard system (Sun et al., 2020) also effectively promoted self-reflection and behavior change intentions.
 A key strength of these interventions is their embedded social context: family dynamics (Joi et al., 2016), peer environments (Haerens et al., 2007), and personal nostalgia (Sun et al., 2020) all serve as external motivators that extend beyond individual willpower, highlighting the importance of the ”social” element.
@@
 In contrast, adults engaged more with emotionally resonant self-reflection tools (Sun et al., 2020).
-6.3.3. Social Feedback with Intervention
-Social feedback with intervention uniquely hybridizes data-driven insights with social dynamics to synergistically promote healthy ingestive behaviors.
+6.3.3. Combined Social Intervention
+Combined social intervention uniquely hybridizes data-driven insights with social dynamics to synergistically promote healthy ingestive behaviors.
 These approaches integrate real-time nutritional tracking and visualization with socially embedded prompts — such as dietary monitoring applications for parents (Alloghani et al., 2016) or office hydration competitions (Chiu et al., 2009) — creating a dual mechanism for behavior change.
@@
 A total of 2 papers are included in this category.
-Regarding the outcomes of this approach, the mobile health monitoring application successfully enhanced user trust and acceptance by combining QR-code food tracking, geo-location, and wearable sensors.
-Users prioritized ease of use and security as key determinants of perceived usefulness (Alloghani et al., 2016).
-Similarly, Playful Bottle leveraged social reminders alongside system-based hydration games, demonstrating that social prompts were more effective than automated reminders.
-This approach not only increased water intake volume but also improved regularity and response speed among office workers (Chiu et al., 2009).
-Notably, social elements consistently outperformed system-only interventions.
-Users of Playful Bottle responded faster and more reliably to peer reminders than to impersonal alerts (Chiu et al., 2009), echoing findings from the obesity app study that indicated trust and usability are heightened by socially embedded design choices (Alloghani et al., 2016).
-6.4. Other Specific Feedback/Intervention
-Other specific approaches not covered above are categorized into three key themes: Perceptual Augmentation (i.e., enhancing sensory engagement with food), Mindful and Psychological (i.e., fostering awareness and emotional connections to ingestion), and Smart Monitoring (i.e., automated tracking systems).
+Regarding the outcomes of this approach, the mobile health app (Alloghani et al., 2016) boosted user trust and acceptance by integrating QR-code tracking, geo-location, and wearable sensors, with ease of use and security being key determinants of perceived usefulness.
+Similarly, Playful Bottle (Chiu et al., 2009) demonstrated that social reminders were more effective than automated alerts, improving both the volume and regularity of water intake among office workers.
+These studies collectively show that socially-embedded interventions consistently outperform system-only approaches, with peer-based prompts generating faster responses and socially-aware designs enhancing both usability and trust (Chiu et al., 2009; Alloghani et al., 2016).
+6.4. Other Specific Interventions
+Other specific approaches not covered above are categorized into three key themes: Perceptual Augmentation, Mindful and Psychological, and Smart Monitoring.
 These frameworks illustrate a range of technological strategies for promoting changes in ingestive behavior.
 6.4.1. Perceptual Augmentation
-Perceptual augmentation actively modifies ingestion experiences by technologically enhancing sensory feedback during consumption.
-These approaches employ real-time cross-modal stimulation—such as altering chewing sounds to influence flavor perception (Kleinberger et al., 2023) or using augmented reality (AR) to visually enhance food texture and taste (Koizumi et al., 2011; Narumi et al., 2011)—to directly reshape how food is sensorially experienced.
-By dynamically manipulating auditory, visual, or haptic cues during the act of ingestion, such interventions create immersive sensory modifications rather than relying on traditional feedback or interventions.
+Perceptual augmentation technologies modify ingestion experiences through real-time cross-modal stimulation, such as altering chewing sounds to enhance flavor perception (Kleinberger et al., 2023) or using AR to visually manipulate texture and taste (Koizumi et al., 2011; Narumi et al., 2011).
+These non-invasive sensory enhancements effectively alter ingestive behaviors—auditory feedback improves texture ratings and reduces hunger, while AR-driven cues (e.g., Metacookie+) achieve a 79.3% success rate in modifying familiar flavors, albeit with lower recognition for unfamiliar tastes (42.9%) (Narumi et al., 2011).
 A total of 3 papers are included in this category.
-Emerging perceptual augmentation technologies demonstrate how non-invasive sensory enhancement can effectively alter ingestive behaviors and perceptions.
-Auditory feedback systems leverage cross-modal effects to enhance flavor perception—significantly improving texture ratings such as crispiness and reducing hunger (Kleinberger et al., 2023; Koizumi et al., 2011).
-Meanwhile, Metacookie+ utilizes AR-driven visual and olfactory cues to manipulate taste perception, achieving a 79.3% success rate in altering familiar flavors, although unfamiliar tastes, such as mushroom, saw lower recognition at 42.9% (Narumi et al., 2011).
 6.4.2. Mindful & Psychological
-Mindful and psychological approaches reshape ingestive behaviors by targeting cognitive-emotional processes and fostering real-time awareness during consumption.
-These interventions employ evidence-based techniques, such as future-self visualization to motivate long-term behavior change (Rho et al., 2017), emotion-aware wearables to detect and interrupt emotional eating (Carroll et al., 2013), and conversational agents to guide mindful food preparation (Parra et al., 2023).
-These strategies directly address the psychological roots of maladaptive eating patterns.
-Ultimately, these interventions cultivate mindful eating by transforming abstract nutritional knowledge into embodied psychological awareness—where every bite becomes an opportunity to nourish both body and mind, and where present-moment consciousness bridges our current selves with the healthier future selves we aspire to become.
+Mindful and psychological approaches reshape ingestive behaviors by targeting cognitive-emotional processes through techniques like future-self visualization (Rho et al., 2017), emotion-aware wearables (Carroll et al., 2013), and conversational agents (Parra et al., 2023).
+These interventions foster real-time awareness, with FutureSelf increasing goal consciousness and EmoTree improving emotional eating awareness by 37.5%.
+While effective in promoting introspection and short-term behavioral change, they require refinement in adaptive personalization and long-term habit formation (Carroll et al., 2013; Rho et al., 2017).
 A total of 4 papers are included in this category.
-Mindful and psychological approaches demonstrate strong potential in promoting healthier ingestive behaviors.
-FutureSelf effectively motivated users by forecasting weight outcomes, with participants reporting increased goal awareness and behavioral conscientiousness (Rho et al., 2017).
-EmoTree’s Just-In-Time interventions improved emotional eating awareness by 37.5% and received an averaged rating of 3 out of 5 for effectiveness in stress reduction, although its impact on preventing eating was limited, with an averaged rating of 2.2 out of 5 (Carroll et al., 2013).
-These interventions succeed by leveraging introspection and emotional engagement but require refinement in adaptive personalization and long-term habit formation (Carroll et al., 2013; Rho et al., 2017).
-Future work should aim to balance immediate support with sustained behavioral reinforcement.
 6.4.3. Smart Monitoring
-Smart monitoring leverages wearable technologies and ambient sensors—from alcohol-detecting ice cubes (Matsui et al., 2018) to eyeglass-mounted detectors (Bedri et al., 2022) and wrist-worn motion trackers (Ye et al., 2016)—to autonomously capture ingestive behaviors (e.g., drinking and chewing) and environmental interactions (e.g., food selection (Volkova et al., 2016)).
-This approach transforms passive observation into actionable behavioral datasets.
+Smart monitoring utilizes wearable and ambient sensors — including alcohol-detecting ice cubes (Matsui et al., 2018), eyeglass-mounted detectors (Bedri et al., 2022), and wrist-worn trackers (Ye et al., 2016) — to autonomously capture ingestive behaviors and environmental interactions (Volkova et al., 2016).
+While demonstrating effectiveness (FitNibble reduced missed logs by 19.6% and improved journaling ease (Bedri et al., 2022), while Pebble scored 2.7 on a scale from -3 to 3 for usability (Ye et al., 2016)), challenges persist (50% of FitNibble users received more than 6 false alerts (Bedri et al., 2022)), highlighting wearability and accuracy tradeoffs.
 A total of 4 papers are included in this category.
-Key strengths in this category include reduced cognitive load (via automated detection) and enhanced self-awareness, particularly regarding snack consumption (Bedri et al., 2022; Ye et al., 2016).
-However, challenges remain, such as false positives—evidenced by 50% of users receiving 6 or more unnecessary alerts in FitNibble (Bedri et al., 2022)—and issues related to device wearability.
-Smart monitoring systems have demonstrated effectiveness in improving dietary tracking compliance.
-FitNibble reduced missed logs by 19.6% and lowered journaling difficulty (Bedri et al., 2022).
-Similarly, the wrist-worn Pebble system received positive usability scores (2.7 on a scale from -3 to 3), particularly benefiting irregular eaters through visual logging prompts (Ye et al., 2016).
 7. Discussion
-We synthesize insights from the reviewed systems to reflect on broader limitations and opportunities in the design of behavioral sensing and feedback/intervention in ingestion health.
-By revisiting the conceptual and technical landscape through the lens of system completeness, behavioral scope, and interdisciplinary integration, we offer a summary analysis that responds to our fourth research question.
-The discussion is organized around three key challenges: (1) the difficulty of accurately measuring and intervening in real-world ingestive behavior, (2) the disconnect between sensing-focused and intervention-driven research efforts, and (3) the methodological and epistemological gap between HCI and clinical health domains.
-For each, we outline design directions aimed at informing future development of more effective and ecologically embedded closed-loop systems.
+We synthesize insights from the reviewed systems to identify challenges and opportunities in designing a closed-loop paradigm for ingestion health, drawing on system completeness, behavioral scope, and interdisciplinary integration to address our fourth research question.
+The discussion centers on three key challenges: (1) the difficulty of accurately measuring and intervening in real-world ingestive behavior, (2) the disconnect between sensing technologies and intervention practices, and (3) the methodological and epistemological gap between HCI and clinical health domains.
+For each, we summarize the underlying issues and propose design principles to support the development of more effective, ecologically embedded closed-loop systems.
 7.1. Challenges in Measuring and Intervening in Real-World Ingestive Behavior
-Despite recent advances in sensing technologies and intervention strategies, significant challenges remain in accurately measuring and applying ingestion behavior interventions in real-world settings.
-These reflections are informed by our previous analysis of intervention effectiveness and user acceptance (Chapter  6), where we observed overall evaluation methods and user feedback in ecological validity, evaluation consistency, and behavioral coverage.
-In particular, the absence of robust and context-sensitive evaluation frameworks continues to hinder the translation of promising closed-loop sensing–intervention systems into everyday practice.
-These challenges can be broadly categorized into three areas:
+Despite rapid advances in sensing technologies and intervention strategies, accurately measuring and intervening in real-world ingestion behavior remains challenging.
+Our analysis of intervention effectiveness and user acceptance in Chapter  6 revealed the absence of ecological validity, consistency, and behavioral coverage.
+The lack of robust and context-sensitive evaluation frameworks makes it difficult for promising closed-loop systems to integrate into daily life.
+This challenge can be further unpacked into the following aspects.
 7.1.1. Validity and Consistency of Ingestive Behavior Measurement Indicators
-As observed in our analysis of feedback and intervention effectiveness (Chapter 6), the lack of consistency in how ingestive behaviors are defined and evaluated poses a major barrier to assessing and comparing intervention outcomes.
-Although many studies focus on similar behavioral constructs, the measurement indicators they employ vary widely, resulting in inconsistencies across interpretation, validation, and generalization.
-Traditional instruments including questionnaires and scales are prevalent in the field of ingestion health intervention, such as the Three-Factor Eating Questionnaire (TFEQ) (Stunkard and Messick, 1985), which assesses cognitive restraint, disinhibition and hunger, and the Eating Behavior Inventory (EBI) (O’Neil et al., 1979), which targets weight-control behaviors.
-Despite their widespread use, their applicability across diverse populations and contexts has still been questioned, remaining non-universal; for emerging eating behavior interventions that rely on sensing and computational methods, the challenge is further compounded by diverse evaluation metrics and the lack of standardized assessment frameworks.
-Even when interventions address the same behavioral target—such as eating speed or intake pace—the metrics used to quantify outcomes are highly heterogeneous.
-To date, medical literature offers no standardized clinical definition for many of these behavioral constructs.
-For example, although rapid eating is commonly linked to adverse health outcomes, there is no standardized definition or measurement approach; current systems rely on varied metrics such as chewing rate, swallowing frequency, bite intervals, or total meal duration to approximate eating speed.
-This diversity highlights not only methodological flexibility but also the absence of a shared ground truth in computationally mediated behavior evaluation.
-Developing valid, standardized, and context-aware behavioral metrics is therefore a critical direction for ensuring scientific rigor and comparability in ingestion health research.
+Our analysis of intervention effectiveness revealed that studies frequently apply inconsistent operational definitions and behavioral metrics, even when addressing the same ingestive behavior construct-hindering interpretation, validation, and cross-study comparability.
+In traditional research of ingestion health, validated questionnaire tools such as the Three-Factor Eating Questionnaire (TFEQ) (Stunkard and Messick, 1985) and the Eating Behavior Inventory (EBI) (O’Neil et al., 1979) are commonly used to assess psychological and behavioral traits.
+While these instruments are well established, they are not well aligned with emerging context-aware approaches.
+Studies using computational methods often adopt their own evaluation criteria, lacking a unified ”gold standard”.
+For instance, in measuring a common behavioral target such as eating speed, no standardized definition currently exists.
+Current systems approximate eating speed using a range of metrics, such as chewing rate, swallowing frequency, bite intervals, or total meal duration.
+This diversity, while reflecting methodological flexibility, also underscores a lack of consistency and validation, which limits cross-study comparability and introduces potential bias in evaluating intervention outcomes.
+Developing valid, standardized, and context-sensitive behavioral indicators is therefore essential to ensuring scientific rigor and cross-study comparability in ingestion health research.
 7.1.2. Laboratory-Based Research vs. Real-World Ingestive Behaviors
-A substantial portion of ingestive behavior research has been conducted in controlled laboratory environments, which, though offering experimental rigor, often fail to reflect the complexity of real-world contexts.
-Lab protocols typically exclude environmental, social, and emotional factors that influence natural eating, leading to potential over- or underestimation of intervention effectiveness when deployed outside the lab (De Castro, 2000; Bell et al., 2020; Hiraguchi et al., 2023).
-For instance, bite and chewing detection systems are often validated in constrained conditions—such as fixed posture, single food types, and minimized distraction—within laboratory settings [e.g., (Scisco et al., 2011; Kim et al., 2016b; Hori et al., 2023; Zhang et al., 2019)].
-However, in real-world environments, behaviors are shaped by multitasking, social interaction, and contextual variability, which can significantly impact system performance and user adherence (Bell et al., 2020; Hiraguchi et al., 2023; Jansen et al., 2022).
-Moreover, some systems using head-mounted cameras, chest-worn sensors, or bulky wearables often perform well in short-term lab trials, but can feel intrusive or uncomfortable in daily life, potentially disrupting natural behavior and lowering compliance (Bedri et al., 2022; Bell et al., 2020).
+A substantial portion of researches on ingestion health is still conducted in controlled laboratory settings.
+While offering rigor, these settings simplify reality by excluding environmental, social, and emotional factors-facilitating statistical detection but risking over- or underestimation of effectiveness when interventions are applied in real-world contexts (De Castro, 2000; Bell et al., 2020; Hiraguchi et al., 2023).
+Bite and chewing detection systems, for instance, are typically validated under constrained lab conditions — fixed posture, limited food types, and minimal distraction [e.g., (Scisco et al., 2011; Kim et al., 2016b; Hori et al., 2023; Zhang et al., 2019)].
+However, in real-world settings, multitasking, social interaction, and context variability can significantly affect system performance and adherence (Bell et al., 2020; Hiraguchi et al., 2023; Jansen et al., 2022).
+Wearable systems, such as head-mounted cameras or chest sensors, may perform well in short lab trials, but often feel intrusive in daily life, potentially disrupting natural behavior and lowering compliance (Bedri et al., 2022; Bell et al., 2020).
 These real-world deployment issues, such as behavioral reactivity and long-term disengagement, are frequently underestimated in lab-based evaluations (Jansen et al., 2022).
 Bridging this gap calls for validation in situation that accounts for behavioral variability, contextual complexity, and sustained user acceptance (Hiraguchi et al., 2023).
-7.1.3. Potential Adverse Effects of Intervention Deployments
-While designed to promote healthier behaviors, eating interventions may produce unintended negative effects.
-A key concern is ”food noise”, which means persistent and intrusive thoughts about eating triggered by constant monitoring.
-It can elevate anxiety and contribute to disordered eating patterns (Hayashi et al., 2023).
-Meanwhile, interventions requiring user interaction during meals, such as smart utensils or app prompts, may also disrupt social interactions or conflict with cultural norms, especially in shared settings (Robinson et al., 2020; Kadomura et al., 2014).
-Moreover, long-term use of intrusive devices like head-mounted sensors or chewing monitors has been linked to discomfort and user dropout (Bedri et al., 2022).
-These risks underscore the need for balancing behavioral influence with psychological comfort and contextual fit.
+7.1.3. Potential Adverse Effects and Privacy/Security Concerns of Intervention Deployments
+While designed to promote healthier behaviors, ingestion interventions may lead to unintended side effects.
+A key concern is ”food noise” — persistent eating-related thoughts triggered by continuous monitoring, which may increase anxiety and contribute to disordered patterns (Hayashi et al., 2023).
+Meanwhile, interventions involving in-meal interaction (e.g., smart utensils, app prompts) may disrupt social dynamics or clash with cultural norms, especially during shared meals (Robinson et al., 2020; Kadomura et al., 2014).
+Moreover, intrusive devices (e.g., head-mounted sensors, chewing monitors) may cause discomfort and user dropout over time (Bedri et al., 2022).
+These risks underscore the need to balance behavioral influence with psychological comfort and contextual fit.
+Finally, privacy and security pose critical ethical concerns for real-world intervention deployment.
+Continuous sensing systems collect highly sensitive health and behavioral data, raising ethical concerns over encryption, anonymization, and transparency in data use (Radanliev, 2025).
+However, with little access to concrete context-aware computing methods in the background, users often struggle to establish trust during usability evaluations (Faltaous et al., 2021).
+Various data-collection modalities with tough methods may provoke user discomfort or resistance (Islam and Bae, 2024; Gao et al., 2023; Mollyn et al., 2022).
+Thus, ensuring data privacy and security is not only ethically imperative but essential for sustained user engagement.
+Together, these concerns call for a user-centered approach to intervention design — one that maximizes behavioral impact without compromising user comfort, privacy, or dignity.
 7.2. Gaps between Sensing-Based and Intervention-Based Studies on Ingestive Behavior
-Recent reviews on sensor-based research in ingestive behavior [e.g., (Hiraguchi et al., 2023; Bell et al., 2020)] have highlighted that while the target behaviors addressed in sensing studies overlap with those identified in intervention-focused research, significant gaps remain between the two approaches.
-These gaps primarily stem from differences in modality alignment and the application of behavior change strategies.
+Recent reviews on sensor-based research in ingestive behavior [e.g., (Hiraguchi et al., 2023; Bell et al., 2020)] reveal a clear gap between sensing-driven and intervention-oriented studies.
+While both address overlapping behavioral targets, they differ significantly in modality integration and the use of behavior change strategies.
 7.2.1. Modality Misalignment between Sensing and Intervention Phases
-Sensing studies have made significant strides in utilizing a wide array of modalities, including wearable devices, ambient sensors, and smartphone-integrated tools, to detect ingestive behaviors with increasing granularity and accuracy [e.g., (Bedri et al., 2022; Zhang et al., 2020)].
-These tools offer real-time monitoring of behaviors like food intake, bite frequency, and meal duration.
-In contrast, interventions rely on a narrower set of modalities, mostly using visual cues like text prompts, graphical feedback, or auditory signals delivered via mobile devices.
-Although some studies [e.g., (Kim et al., 2016b)] have used wearable devices for both sensing and delivering interventions, such integrated applications remain relatively scarce, as shown in Chapter 5.3.
-This disconnect between detection and intervention modalities creates a lack of continuity in the intervention process, limiting the potential for closed-loop systems where detection and feedback are fully integrated.
-For instance, while detection technologies capture behavioral data, the feedback often remains confined to basic prompts, which may not effectively engage users or encourage sustained behavior change.
+Sensing studies have employed diverse modalities, including wearable devices, ambient sensors, and smartphone-integrated tools, to detect ingestive behaviors with increasing granularity and accuracy [e.g., (Bedri et al., 2022; Zhang et al., 2020)].
+These systems offer real-time monitoring of behaviors like food intake, bite frequency, and meal duration.
+In contrast, most intervention applications rely on a narrow set of output channels, typically visual cues like text prompts, graphical feedback, or auditory signals via mobile devices.
+Although some efforts [e.g., (Kim et al., 2016b)] have explored using the same device for both sensing and intervention, such integrated applications remain uncommon, as shown in Chapter 5.3.
+This misalignment between sensing and intervention modalities disrupts continuity in the closed-loop process — while sensors capture detailed, real-time behavioral data, interventions are often reduced to basic prompts that struggle to engage users or support long-term behavior change.
+This gap calls for tighter integration between sensing and intervention modalities through compatible devices, enabling a unified user experience.
 7.2.2. Disconnect Between Detection-Centric Studies and Actual Intervention Implementation
-A second significant gap exists in the large body of research only focus on improving detection accuracy ,but without progressing to actionable interventions.
-Many sensor-based studies mention behavior change as a long-term goal but fall short of translating their findings into real-world applications.
-This disconnection is largely due to the differing complexities of sensing and intervention research.
-Detection-centric studies typically focus on technical performance metrics, such as the accuracy of behavior detection and system usability, which are easier to evaluate.
-On contrary, intervention studies must deal with a broader range of challenges, including incorporating behavior change theories, designing meaningful user interactions, and establishing long-term engagement.
-The higher threshold for evaluating interventions has contributed to a bottleneck, where many studies stop after validating detection methods, while fewer studies advance to actual intervention implementation.
-As a result, there is an imbalance in the literature, with a higher volume of studies dedicated to detection accuracy, and fewer addressing the complexity of behavior modification and sustained interventions.
+Many sensor-based studies of ingestive behavior emphasize detecting and monitoring intake-related events but stop short of delivering actionable interventions.
+While behavior change is often cited as a long-term goal and occasionally guided by behavioral frameworks (Hekler et al., 2013), most studies ultimately focus on improving detection accuracy, as noted in Section 5.
+This imbalance stems from differing evaluation paradigms: detection studies prioritize quantifiable metrics (e.g., recognition accuracy), which are easier to validate in controlled settings, whereas interventions are often treated as downstream demonstrations—supporting sensing pipelines rather than contributing novel design knowledge (Rapp and Boldi, 2023), as discussed in Chapter 6.
+Designing interventions, however, involves more complex challenges, including applying behavior change theories, supporting meaningful user interaction, and sustaining long-term engagement—factors difficult to capture through standard evaluation methods.
+These demands contribute to a bottleneck where sensing advances outpace the development of effective intervention strategies.
+Beyond methodological barriers, a deeper issue lies in the limited integration of behavioral theory.
+Although behavior theories are often cited, they rarely inform system design in a structured way (Hekler et al., 2013).
+For instance, the Fogg Behavior Model (Fogg, 2009) defines behavior as a convergence of motivation, ability, and trigger.
+Though some interventions loosely map sensing accuracy to ”ability”, context to ”motivation”, and reminders to ”triggers”, often retrofitted to justify design choices.
+In our review, few systems explicitly used such frameworks to shape intervention logic or personalize support.
+This shallow application reflects a broader reliance on intuition over theory, with key constructs, like habit, emotion, and social support, frequently overlooked.
+As a result, interventions may be theoretically underpowered and less effective in driving sustained behavior change.
 7.3. Human-Computer Interaction vs. Medical Perspectives on Ingestive Behavior
-Although both Human-Computer Interaction (HCI) and medical research aim to improve ingestive behaviors, such as dietary patterns and meal composition, they differ in their assessment frameworks and metrics.
-HCI typically focuses on real-time, sensor-based interventions and user experience, whereas medical research emphasizes clinical assessments and physiological measures.
-These differences pose challenges in aligning the methodologies and measuring the effectiveness of intervention.
-Additionally, while both fields acknowledge the importance of subjective experiences and physiological states, their integration into behavior metrics differs.
-HCI research tends to incorporate subjective feedback via interactive devices, while medical research relies more heavily on objective physiological data.
-This gap highlights the limitations of current HCI technologies in capturing the complex physiological aspects of ingestive behavior.
+Although both Human-Computer Interaction (HCI) and medical research aim to promote healthier ingestive behaviors, they differ significantly in how they define, measure, and interpret such behaviors.
+HCI approaches tend to focus on context-aware computing and interactive methods (Nahum-Shani et al., 2016; Hekler et al., 2016), while medical research emphasizes clinically validated assessments and physiological indicators.
+These differences pose challenges in aligning methodologies, interpreting efficacy, and integrating subjective experiences and biological states.
+In the following sections, we explore two key dimensions of this divergence — the assessment frameworks and treatment of subjective and physiological aspects of ingestive behavior.
 7.3.1. Divergent Assessment Frameworks
@@
 Medical practitioners use comprehensive nutritional assessments that combine food history, biochemical data, medical tests, and clinical procedures (Kesari and Noel, 2022).
-Tools like the Food Frequency Questionnaire (FFQ) (Cade et al., 2004), Three-Factor Eating Questionnaire (TFEQ) (Karmali et al., 2013), and Adult Eating Behavior Questionnaire (AEBQ) (Hunot et al., 2016) offer a validated, multi-domain approach for assessing ingestive behaviors and their relationship to health outcomes.
-In contrast, HCI research often relies on isolated validation methods, lacking the cross-validation mechanisms inherent in medical assessments.
-While HCI interventions focus on sensor-based technologies and user behavior tracking, they do not integrate the multi-domain assessments seen in medical protocols.
-This difference limits HCI-based interventions in capturing the complex interactions between physiological states and behavioral patterns.
-HCI tools primarily focus on real-time data acquisition (e.g., food intake, eating speed), but they often overlook broader health contexts, such as digestive symptoms or stress levels.
-This results in an incomplete understanding of the long-term impact of HCI interventions on health outcomes.
+For instance, validated instruments such as the Food Frequency Questionnaire (FFQ) (Cade et al., 2004), Three-Factor Eating Questionnaire (TFEQ) (Karmali et al., 2013), and Adult Eating Behavior Questionnaire (AEBQ) (Hunot et al., 2016) can capture multiple aspects, like appetite traits, restraint, and emotional eating, and help relate ingestion patterns to risks.
+In contrast, HCI research often relies on isolated validation methods, lacking the multi-domain cross-validation seen in clinical practice.
+For example, HCI studies rarely combine physiological indicators (e.g., glucose, hormones) with psychological measures to evaluate intervention effects.
+This limits the ability to capture interactions between physiological states and behavioral patterns.
+HCI tools focus on real-time data (e.g., food intake, eating speed), but often ignore broader health factors like stress or digestion.
+As a result, their long-term health impact remains unclear.
+Medical approaches, in contrast, emphasize multi-dimensional and longitudinal assessments, offering a deeper view of intervention effectiveness.
 7.3.2. Limited Perception of Subjective Experience and Physiological States in Target Behavior Metrics
-Medical assessments integrate subjective experiences (e.g., emotional eating, satiety cues) with physiological states, offering a comprehensive evaluation of ingestive behaviors essential for disease diagnosis (Kesari and Noel, 2022).
-These tools benefit from (1) long-term validation, (2) structured clinical environments that enhance patient compliance, and (3) professional interpretation by healthcare specialists.
-These assessments are particularly effective in documenting longitudinal changes in variables such as pain, function, and psychosocial distress.
-In contrast, early HCI research relied on questionnaires that (1) had limited long-term validation, (2) faced low accuracy and compliance due to experimental constraints, and (3) lacked professional interpretation.
-While HCI interventions track behaviors like chewing or swallowing, they fail to capture subjective experiences (e.g., satiety or emotional states), essential for understanding ingestive behaviors comprehensively.
-This gap highlights a key limitation: sensors can tell us how we eat, but they do not address why we eat or how we feel during eating, which are crucial aspects for understanding ingestive behaviors in the context of health outcomes.
-Thus, while HCI tools focus on quantifiable metrics (e.g., food intake or eating speed), they overlook the subjective aspects emphasized by medical assessments.
-To address this gap, HCI interventions should integrate both subjective and physiological data to offer a more holistic understanding of ingestive behaviors, particularly in real-world contexts.
+Medical assessments emphasize the integration of subjective experiences (e.g., emotional eating, satiety, post-meal discomfort) with physiological states to form a comprehensive evaluation of ingestive behavior, particularly for diagnosing and managing conditions like functional dyspepsia or eating disorders (Kesari and Noel, 2022).
+These assessments benefit from validated instruments with long-term reliability, structured clinical environments that support patient compliance, and professional interpretation by healthcare practitioners.
+For example, longitudinal evaluation of chronic conditions often relies on self-reported changes in symptoms (e.g., pain, bloating, appetite) combined with physiological data to guide treatment.
+By contrast, most HCI studies rely on short-term self-report tools with limited validation and low compliance in non-clinical settings.
+While sensor-based systems excel at quantifying behaviors (e.g., eating rate, chewing duration), they often fail to capture the subjective dimensions essential to understanding why users eat the way they do.
+Early HCI studies seldom included validated measures of satiety or emotional states, and even with the rise of ubiquitous computing, the capacity to sense nuanced experiential factors remains limited.
+This overemphasis on observable behavior, without accounting for users’ internal states, restricts the interpretability and impact of HCI interventions.
 7.4. Design Recommendations
+To address the core challenges identified above, we propose design recommendations and future directions to support more effective and ecologically embedded closed-loop interventions for ingestive behavior.
+Advancing this domain requires coordinated efforts across researchers, designers, and practitioners, spanning theory, technical implementation, and real-world integration.
+The following subsections align our suggestions with the key challenges, offering concrete guidelines for each domain.
 7.4.1. Optimizing Experimental Design and Intervention Implementation
-To bridge the gap between laboratory-based research and real-world interventions, clear and standardized measurement metrics should be established, drawing from validated constructs in the existing literature.
-Combining qualitative and quantitative methods provides a comprehensive view of intervention outcomes, helping reduce bias and ensuring a more accurate evaluation.
-Additionally, integrating self-reported data with objective physiological measures offers a balanced approach to long-term tracking, avoiding participant fatigue.
-For example, targeted micro-assessments, triggered by unusual behavior patterns, can provide timely and context-specific insights without overwhelming participants.
-When designing experimental protocols, it is essential to consider real-world conditions, user experience, and long-term feasibility.
-Instead of excluding common unhealthy behaviors, such as continuous eating without proper pauses, interventions should directly address them as part of the behavior change process.
-This ensures that interventions are more relevant to natural eating contexts, promoting sustainable behavior change.
-For long-term engagement, the technical design should prioritize usability and integration with daily life.
-Embedding sensing capabilities into familiar, unobtrusive devices—such as smartwatches with inertial sensors or bone conduction earphones—can enhance natural interaction and reduce disruption.
-This approach will ensure continuous monitoring while preserving comfort and user adherence.
-Furthermore, ethical safeguards must be incorporated into the intervention design.
-Participants should be clearly informed that they can withdraw at any time if they feel discomfort or emotional distress, ensuring their psychological well-being is not compromised by the intervention.
-Finally, given the inconsistency in measurement tools across studies, future interventions should consider context-sensitive evaluation methods.
-These should account for cultural differences and ensure interventions align with both natural ingestive behaviors and social norms, improving applicability across diverse populations.
+To move beyond proof-of-concept prototypes toward impactful interventions, researchers should adopt standardized, theory-driven, and context-sensitive evaluation frameworks.
+Researchers should combine validated behavioral constructs with multimodal sensing data (e.g., self-report diaries and sensors) to contextualize outcomes and reduce measurement bias.
+Methodologically, designs such as factorial experiments, SMART, and micro-randomized trials are well-suited to evaluating multi-component systems (Nahum-Shani et al., 2022).
+Such designs enable precise testing of when and how interventions are most effective, ultimately informing evidence-based guidelines.
+To reduce burden while preserving ecological validity, researchers may adopt micro-assessments triggered by behavioral anomalies to gather timely and lightweight feedback.
+Grounding evaluations in behavioral theories helps bridge the disconnect between technological innovation and behavior change science.
+Intervention tools should be intuitive, low-burden, and compatible with daily routines, such as embedding this loop into smart utensils or wearables.
+Co-locating sensing and intervention delivery within a single ecosystem improves usability and streamlines interaction.
+Designers should provide timely, actionable interventions rather than passive data to support real-time engagement.
+For example, FitNibble (Bedri et al., 2022) and Slowee (Kim et al., 2016b) demonstrate how subtle prompts can improve adherence and engagement in natural settings.
+Rather than filtering out common unhealthy behaviors in study design, interventions should address them explicitly as change targets to better reflect real-world eating challenges.
+Privacy and trust should be core design priorities in always-on, closed-loop systems.
+Designers should implement privacy-preserving strategies, including context-aware data collection (Nepal et al., 2024; Saha et al., 2021) and layered encryption (Kumar et al., 2020).
+Practitioners should ensure transparent communication about data use and obtain sustained informed consent (Haghighi et al., 2023; Adler et al., 2022; Islam and Bae, 2024).
+Systems should enable user autonomy, allowing for flexible participation and empathetic adjustments when engagement decreases or discomfort arises.
+To enhance generalizability, practitioners should iteratively adapt interventions to diverse physiological and cultural contexts.
+Integrating behavioral coaching with tech-driven systems ensures more personalized and effective support.
+By combining scientific rigor, user-centered design, and ethical awareness, interventions can better support sustained ingestive behavior change in the real world.
 7.4.2. Enhancing Integration and Adaptability in Sensing-Driven Interventions for Ingestion Health
-To effectively bridge the gap between sensing and intervention in ingestion health, future interventions should focus on the seamless integration of real-time data from sensing technologies into personalized, actionable, and adaptive interventions.
-To enhance the continuity of the closed-loop system, interventions should increasingly leverage the same or compatible devices for both sensing and feedback delivery.
-Such consistency in device use not only improves the intuitiveness and coherence of the intervention process but also fosters user familiarity and acceptance over time.
-More personalized and actionable intervention requires moving beyond basic detection capabilities, not only by gathering more granular physiological and behavioral data but also by ensuring that feedback mechanisms are tailored to the specific needs, preferences, and circumstances of each user.
-For instance, interventions must consider factors such as individual eating patterns, meal timing, emotional states, and physiological markers like glucose levels, enabling a level of personalization that ensures more effective engagement and results (Liang, 2024; Reynolds et al., 2019).
-Moreover, interventions should be designed with adaptability in mind.
-Success in behavior change depends not only on tracking short-term ingestive behaviors but also on continually adapting the intervention based on long-term data.
-This means creating systems that can evolve their feedback strategies in response to behavioral trends and progress over time.
-Such adaptive interventions maintain their relevance and motivational power, preventing user fatigue and promoting long-term engagement.
-A crucial aspect of this adaptability involves understanding the dynamic nature of ingestive behaviors.
-For example, certain behaviors may be triggered by emotional states or specific environmental factors, which require intervention systems that can quickly detect and respond to these variables.
-This requires the continuous collection and analysis of both physiological and behavioral data, ensuring that interventions are timely and appropriate.
-Finally, the long-term success of such systems hinges on the ability to sustain behavior change.
-The intervention should encourage not just short-term compliance but foster lasting habits by gradually increasing the complexity or frequency of the feedback, thereby encouraging the user to internalize healthier behaviors and develop autonomous control over their eating habits.
-This continuous, evolving support through adaptive interventions will ensure that users are not only supported in their current behavior change but are also equipped with the tools to sustain these changes in the long run.
+To close the gap between sensing and intervention in ingestion health, future systems must better integrate behavioral theory into personalized, adaptive strategies.
+Many recent interventions emphasize surface-level mechanisms, like rational reflection or automated nudges, without deeply engaging with why those mechanisms work across individuals.
+This limits their long-term impact, as users may revert to old habits once external support ends.
+Researchers should embed theory-driven constructs into system logic.
+Behavior change theories such as the Fogg Behavior Model (Fogg, 2009) or Social Cognitive Theory offer constructs like motivation, ability, and self-efficacy that should guide not just evaluation but delivery decisions — when to intervene, what content to present, and how to adapt dynamically.
+Rather than cherry-picking isolated elements, researchers should operationalize full frameworks to understand both success and failure.
+They should also personalize adaptively.
+Building on the JITAI paradigm (Nahum-Shani et al., 2016), systems can adjust message type and timing based on real-time behavioral or physiological trends.
+For instance, if late-night snacking consistently correlates with stress spikes, systems may prioritize emotion regulation prompts over dietary tracking.
+Designers should tightly couple sensing and intervention in a closed loop.
+Physiological signals like chewing rate or glucose dips should directly trigger timely, context-aware interventions.
+Integrated systems, such as SenseSupport (Liang, 2024), reduce user friction and enhance moment-to-moment relevance.
+They should also scaffold long-term autonomy.
+Systems should evolve with the user—starting with frequent, specific prompts, then gradually tapering or increasing complexity to foster independence.
+Practitioners should ground behavior change in lived experience.
+Traditional models often abstract behavior from context, but effective change requires alignment with users’ identities, values, and goals (Lee et al., 2020).
+To foster sustained motivation, interventions should facilitate meaning-making by helping users relate new eating habits to their personal values, such as family responsibility or personal development (Rapp and Boldi, 2023).
 7.4.3. Aligning with Medical Goals and Practices
-To bridge the gap between HCI-based interventions and medical objectives, personalized interventions should account for the diversity in ingestive behaviors, physiological baselines, and cultural backgrounds among participants (Mirmiran et al., 2021).
-Incorporating both subjective and physiological data, especially baseline parameters like blood glucose and heart rate variability, will improve the precision and effectiveness of interventions (Liang, 2024; Harous et al., 2017).
-Future studies should adopt dynamic intervention models and extend trial durations to better capture long-term effects.
-Exploring dose-response relationships, including varying intervention frequency, duration, and intensity, will help identify the optimal parameters for sustained health benefits (Pankajavalli et al., 2017; Kim et al., 2016b).
-Emerging behavioral and physiological targets such as postprandial heart rate variability, blood glucose response curves, and digestive symptoms (e.g., bloating, nausea) offer valuable insights into metabolic health.
-Continuous glucose monitoring (CGM) has shown promise in real-time glucose tracking, and integrating it with wearable technologies, including smartwatches and fitness trackers, can offer multi-dimensional feedback on ingestive behavior and metabolic response (Liang, 2024; Bedri et al., 2022; Zhao et al., 2021).
-Multimodal sensing approaches, integrating wearable and minimally invasive technologies, could enable more comprehensive tracking of physiological and behavioral signals, enhancing intervention outcomes (Zhang and Parker, 2020).
-Future sensor fusion techniques, combining CGM with emotion detection (EDA) and activity tracking, will offer a complete view of ingestive behavior and its impact on health (Carroll et al., 2013).
-As sensor technology progresses, real-time systems could detect subtle digestive processes, like esophageal peristalsis during swallowing, to assess intervention effects on digestive function (Tirasirichai et al., 2018; Biyani et al., 2019).
-These advancements will improve our understanding of nutrition-physiology relationships and provide actionable insights for chronic disease prevention (Reynolds et al., 2019; Bi et al., 2016).
+For sensing-driven interventions to achieve lasting health benefits and clinical relevance, they must align with medically meaningful outcomes and practices.
+Researchers should work with health professionals to define targets beyond behavioral proxies—such as improvements in glycemic control, cholesterol, gastrointestinal symptoms, or long-term weight trends.
+Incorporating baseline physiological parameters (e.g., glucose levels, heart rate variability) alongside behavioral data can enhance personalization and clinical interpretability (Liang, 2024; Harous et al., 2017; Mirmiran et al., 2021).
+Future studies should adopt longer-term and dynamic intervention models to assess sustained physiological effects.
+Exploring dose-response relationships and optimizing frequency, intensity, and timing can help identify scalable strategies (Pankajavalli et al., 2017; Kim et al., 2016b).
+Drawing on techniques from clinical behavior change approaches—such as motivational interviewing or cognitive-behavioral therapy—can further ground interventions in established medical frameworks.
+Multimodal sensing and sensor fusion offer promising directions.
+Combining continuous glucose monitoring (CGM), emotion sensing (e.g., EDA), activity tracking, and digestive symptom monitoring enables more comprehensive understanding of ingestive behavior and metabolic response (Bedri et al., 2022; Zhang and Parker, 2020; Zhao et al., 2021).
+These data streams can support real-time personalization, such as prompting slower eating when postprandial glucose rises or initiating mindfulness exercises during periods of physiological stress.
+Finally, aligning with medical goals invites broader thinking about precision health.
+Though this work centers on eating behavior, the underlying strategies can extend to domains like sleep, medication adherence, and stress management.
+Framing these systems within the larger precision health movement will improve scalability, adoption, and long-term public health impact.
+7.4.4. Extending the Paradigm Beyond Ingestion Health
+The proposed behavioral closed-loop paradigm, while developed in the context of ingestion health, offers a generalizable framework for a wide range of behavior change interventions.
+As demonstrated by prior reviews (Lara and Labrador, 2012; Thomas Craig et al., 2021; Mair et al., 2023), wearable and mobile sensing technologies have been successfully employed in domains such as physical activity promotion, dietary behavior optimization, chronic disease management, substance use reduction, and mental health support.
+These systems often combine continuous sensing (e.g., step counts, glucose levels, stress markers), reasoning components (e.g., activity inference, risk prediction), and real-time interventions (e.g., feedback messages, prompts), aligning closely with the structure of our proposed closed-loop model.
+Continuous, context-aware sensing should serve as the foundation for detecting actionable behavior patterns across domains.
+For instance, physical activity interventions detect sedentary behavior and deliver timely prompts to encourage movement (Boerema et al., 2019; Van Dantzig et al., 2018); chronic disease management systems combine physiological sensing with persuasive feedback to improve self-management (Chandler et al., 2019; Coorey et al., 2019).
+Real-time, personalized intervention strategies should adapt to users’ physiological, psychological, and contextual states.
+These examples highlight the potential to extend our framework beyond eating-related behaviors while retaining its core responsiveness and personalization.
+The closed-loop structure provides a modular blueprint that can unify design across domains, enabling cross-disciplinary translation and theory-driven development.
+By abstracting core design elements — continuous sensing, adaptive reasoning, and context-aware intervention — the paradigm facilitates broader scalability and long-term health impact in ubiquitous computing.
 8. Conclusion and Limitations
 In this review, we systematically analyzed 136 studies focused on behavioral sensing and intervention paradigms in ingestion health, with an emphasis on sensor-based systems and interactive interventions.
-These studies were categorized into various types based on the modality of feedback and the behavioral targets of the interventions.
-A key contribution of this review is the development of a behavioral closed-loop framework, which highlights the importance of seamlessly integrating sensing, reasoning, and feedback mechanisms for effective behavior change.
-This framework offers a detailed structure for the design of adaptive, context-aware interventions, targeting both physiological and behavioral aspects of ingestive health.
+These studies were categorized into various types based on the modality and the behavioral targets of the interventions.
+A key contribution of this review is the development of a behavioral closed-loop framework, which highlights the importance of seamlessly integrating sensing, reasoning, and intervention mechanisms for effective behavior change.
+This framework offers a detailed structure for the design of adaptive, context-aware interventions, targeting both physiological and behavioral aspects of ingestion health.
+It is theoretically grounded in context-aware computing and HCI behavior change frameworks, offering a modular foundation for system design across domains.
 However, despite the comprehensive nature of this review, there are several limitations that should be acknowledged.
-Our keyword search strategy, centered around the term “ingestion health,” may have excluded relevant studies that do not explicitly focus on ingestion but contribute to broader research in behavior modification or health interventions.
+Our keyword search strategy, centered around the term ”ingestion health”, may have excluded relevant studies that do not explicitly focus on ingestion but contribute to broader research in behavior modification or health interventions.
 Moreover, the majority of the studies in HCI domain reviewed were short-term, laboratory-based experiments, which limits the generalizability of our findings to real-world settings.
 The dynamic nature of real-life eating behaviors requires more longitudinal studies to better understand the long-term effectiveness and challenges of these interventions.
+Finally, the modular structure of the proposed paradigm — grounded in continuous sensing, adaptive reasoning, and real-time intervention — holds promise for application beyond ingestion health, such as in sleep, stress, or physical activity interventions.
+We hope this review helps lay the foundation for more generalizable, stakeholder-sensitive, and theoretically informed systems in ubiquitous health computing.
 Acknowledgements.
-Appendix A RESULT COUNTS OF PAPER VENUES
+Appendix A SEARCH DETAILS CORRESPONDING TO EACH SOURCE
+ACM Digital Library query target: Title, Keywords, Abstract filters: No keynote, tutorial, invited talk, panel inquiry result: 679 IEEE Explore query target: Document Title (eating), Abstract (eating, intervention), All Metadata (technology) filters: Conferences, Journals inquiry result: 911 Google Scholar filters: No books inquiry result: 100
+Appendix B RESULT COUNTS OF PAPER VENUES
 Table 1. Venues and number of publications found after applying the SLR search queries.
@@
 Total 136
-Appendix B DISTRIBUTION OF PUBLICATIONS ACROSS TARGET FACTORS AND BEHAVIORS
+Appendix C DISTRIBUTION OF PUBLICATIONS ACROSS TARGET FACTORS AND BEHAVIORS
 Table 2. Taxonomy of Target Factors and Behaviors with Involved Publications
@@
   Safe Eating Practices (Wang and Zhang, 2021; Singhal et al., 2024)
-Appendix C DISTRIBUTION OF PUBLICATIONS IN SENSING MODALITY
+Appendix D DISTRIBUTION OF PUBLICATIONS IN SENSING MODALITY
 Table 3. Taxonomy of Sensing Modality with Involved Publications
@@
    Magnetic (Vazquez et al., 2015; Han and Kang, 2017; Huang et al., 2010; Jung et al., 2017; Yıldız and Coşkun, 2019)
-Appendix D DISTRIBUTION OF PUBLICATIONS IN FEEDBACK/INTERVENTION MODALITY
-Table 8. Taxonomy of Feedback/Intervention Modality with Involved Publications
-Feedback/Intervention Modality References
+Appendix E DISTRIBUTION OF PUBLICATIONS IN INTERVENTION MODALITY
+Table 4. Taxonomy of Intervention Modality with Involved Publications
+Intervention Modality References
 Visual
@@
   Electrodermal (Zheng et al., 2025; Biyani et al., 2019)
+Appendix F MODALITY CATEGORIES IN SENSING PROCESS
+Table 5. Descriptions and category of modality terms in sensing process
+Modality Submodality Description Paper count
+Visual Human-based visual captures human-derived visual cues to infer the user’s attention, emotional state, or physical actions during ingestion-related behaviors 8
+Gaze Gaze monitors the user’s eye movements, fixation points, and visual attention patterns to infer focus, engagement, or intention during ingestion episodes 2
+Facial expression Facial expression analyzes dynamic changes in facial muscle movements to detect emotional responses, physiological states, or behavioral reactions associated with eating activities 4
+Body posture Body posture captures the orientation, alignment, and movement of the user’s body or head to interpret ingestion-related behaviors via camera, such as eating gestures, drinking actions, or body positioning during meals 3
+Auditory Auditory captures human-generated acoustic signals, including speech and non-speech sounds, to infer behavioral states, contextual cues, or ingestion-related activities 14
+Speech Speech detects and analyzes verbal expressions, vocalizations, or conversational cues to infer user intentions, social interactions, or eating-related contexts 7
+Sound Sound captures non-verbal acoustic signals generated during ingestion activities, such as chewing sounds, swallowing noises, or utensil interactions, to infer ingestive behaviors or consumption events 8
+Haptic Haptic captures tactile and kinesthetic information through physical interactions with the body, providing insights into touch-based behaviors during ingestion activities 14
+Motion Motion use inertial measurement unit (IMU), bioelectrical sensors, or spectral sensors to capture the movement and dynamic postures of the human body or its parts, enabling the recognition of behaviors, gestures, or ingestion-related activities 42
+Gesture Gesture detects voluntary movements of the hands, arms, or other body parts using motion sensors to infer user actions, commands, or contextual cues during ingestion-related behaviors 37
+Chewing and Micro-Head Motion Chewing and micro-head motion monitors rhythmic jaw movements and subtle head or neck dynamics associated with eating, drinking, or swallowing 5
+Physiological Physiological captures internal bodily signals related to brain activity, cardiovascular function, and autonomic system responses, enabling the inference of users’ physiological states during ingestion-related activities 16
+Cardiac Cardiac enables continuous monitoring of cardiovascular dynamics, including heart rate (HR) and heart rate variability (HRV), typically through electrocardiogram (ECG) or photoplethysmography (PPG) sensors, to assess physiological responses during eating episodes 7
+Metabolic Metabolic captures nutritional and biochemical markers such as blood glucose, cholesterol, and body fat rate, typically using biosensors to assess metabolic responses and long-term ingestion-related health status 6
+Human-based Autonomic Autonomic captures involuntary signals, including electrodermal activity (EDA) and skin temperature, using biosensors to monitor autonomic activity during ingestion 7
+Visual Environment-based visual captures information about food items, environments, or objects through image-based techniques such as classification, segmentation, or object detection 29
+Acoustic Acoustic detects environmental or object-related features by analyzing sound wave propagation, reflection, or emission patterns 7
+Force Force measures mechanical pressure or weight changes to infer interactions with food containers, plates, or utensils during ingestion activities 13
+Spectral Spectral analyzes the absorption, reflection, or emission of electromagnetic waves to non-invasively detect material properties such as food composition or temperature 10
+Electrical Electrical captures environmental conditions or object characteristics by measuring electrical properties like resistivity, conductivity, or electromagnetic responses 15
+Conductivity Conductivity measures the electrical conductance of liquids or materials to infer properties such as food consistency, liquid presence, or liquid quantity 10
+Environment-based Magnetic Magnetic sensing detects the presence, identity, or proximity of ingestion-related objects using techniques such as Radio-Frequency Identification (RFID) or Near-Field Communication (NFC) technologies 5
+Appendix G MODALITY CATEGORIES IN INTERVENTION PROCESS
+Table 6. Descriptions and category of modality terms in intervention process
+Modality Submodality Description Paper count
+Visual Textual Delivering feedback through text-based information, such as messages, notifications, or instructions 85
+Graphical Providing feedback using graphical elements like charts, progress bars, icons, or visualizations to convey behavioral states or intervention cues 41
+Symbolic Giving visual feedback through abstract or simplified representations, such as symbols, flashing lights, or color changes, to convey states or prompts with minimal cognitive load or linguistic interpretation 21
+[1pt/1pt] Auditory Delivering feedback via sound signals, including speech output, alerts, tones, or auditory cues 32
+[1pt/1pt] Haptic Vibration Providing feedback through mechanical oscillations delivered by actuators in devices, enabling users to perceive alerts, prompts, or behavioral cues 10
+Deformation Delivering feedback by altering the physical shape, structure, or volume of an interface, allowing users to perceive behavioral cues through tangible disruptions embedded within their natural interaction actuators 7
+[1pt/1pt] Olfactory Providing feedback through controlled release or modulation of odors to stimulate the sense of smell, potentially affecting appetite, cravings, or ingestive behavior 3
+[1pt/1pt] Physiological Electrodermal Delivering feedback by modulating skin conductance through electrical stimulation, aiming to influence autonomic nervous system activity and behavioral states during ingestion processes 2
+Appendix H DESIGN METRIC BETWEEN SENSING MODALITIES AND TARGET FACTORS
+Table 7. Mapping of Sensing Modalities to Target Factors and Behaviors in Ingestion Health
+Sensing Modality Human-based Environment-based Physiological Electrical Visual Auditory Haptic Motion Cardiac Metabolic Autonomic Visual Acoustic Force Spectral Conductivity Magnetic Food Choices facial expression (Robinson et al., 2020; Triantafyllidis et al., 2019), gaze (VR) (Karkar et al., 2018) speech (Prasetyo et al., 2020; Arefeen et al., 2022; Luo et al., 2021; Robinson et al., 2020), intake sound (Bi et al., 2016; Maramis et al., 2014; Kalantarian et al., 2014) touch (Cai et al., 2024; Robinson et al., 2020), oral behavior (Abisha and Rajalakshmy, 2017), body weight (Harous et al., 2017; Ludwig et al., 2011; Bastida et al., 2023; Nedungadi et al., 2018) wrist trajectory (Faltaous et al., 2021; Dragoni et al., 2017; Maramis et al., 2014), Object trajectory (Kadomura et al., 2013b; Nakamura et al., 2023; Joi et al., 2016; Kadomura et al., 2014, 2013a; Bedri et al., 2022), exercise (Kumbhare et al., 2023; Mishra et al., 2024; Harous et al., 2017; Dragoni et al., 2017; Maramis et al., 2014; Bastida et al., 2023; Chukwu, 2011; Badawi et al., 2012a; Fuchs et al., 2019; Tirasirichai et al., 2018; Jung et al., 2007; Vazquez et al., 2015) blood pressure (Huang et al., 2018; Jung et al., 2007; Ludwig et al., 2011; Park and Min, 2007; Nedungadi et al., 2018), heart rate (ECG) (Jung et al., 2007; Park and Min, 2007; Badawi et al., 2012b; Nedungadi et al., 2018), blood oxygen (Park and Min, 2007; Nedungadi et al., 2018) CGM (Liang, 2024; Arefeen et al., 2024; Nedungadi et al., 2018), BIA (Balbin et al., 2020; Chukwu, 2011), cholesterol (Harous et al., 2017) temperature (Park and Min, 2007; Nedungadi et al., 2018; Jung et al., 2007) food picture (Faltaous et al., 2021; Nakamura et al., 2023; Cai et al., 2024; Kadomura et al., 2014, 2013a; Ganesh et al., 2014; Arefeen et al., 2022; YR et al., 2024; Mishra et al., 2024; Alloghani et al., 2016; Hsiao et al., 2011, 2011; Achilleos et al., 2017; Kanjalkar et al., 2024; Kwon et al., 2016; Eigen et al., 2018; Cordeiro et al., 2015; Waltner et al., 2015; Epstein et al., 2016; Akter et al., 2021), Nutrition Label (Hsiao et al., 2011; Fernandez et al., 2013; Gutiérrez et al., 2018; Fuchs et al., 2019; Volkova et al., 2016; Eyles et al., 2017; Mankoff et al., 2002; Vazquez et al., 2015) food removal (Liu et al., 2024) Meals and plate weight (de Vries et al., 2020; Joi et al., 2016; Jung et al., 2017; Zhao et al., 2021; Ritschel et al., 2018; Krishna et al., 2025; Maramis et al., 2014; Kim et al., 2016c) alcohol (Matsui et al., 2018), temperature (Bastida et al., 2023; Kwon et al., 2016; Badawi et al., 2012b), object state (Ritschel et al., 2018), food distance (Bedri et al., 2022) humidity (Bastida et al., 2023; Kwon et al., 2016; Badawi et al., 2012b) Intake Timing intake sound (Bi et al., 2016) button click (Fortmann et al., 2014) wrist gesture (Ye et al., 2016) CGM (Liang, 2024) heart rate (Varun et al., 2023) liquid level (Varun et al., 2023) food weight (Lessel et al., 2016) temperature (Varun et al., 2023) humidity (Varun et al., 2023) Environment indoor position (Faltaous et al., 2021) Context Factors Activity Around Ingestion postprandial exercise (Alexander et al., 2017) Intake Quantity body gesture (Xie et al., 2023) oral behavior (Farooq et al., 2017) bite count (Scisco et al., 2011), object trajectory (Zhou et al., 2021; Li and Bailenson, 2018) food picture (Chiu et al., 2009) leftovers (Faltaous et al., 2021), liquid level (Varun et al., 2023; Tommy et al., 2017; Ravindran et al., 2022) liquid level (Lessel et al., 2016) liquid level (Poddar et al., 2024) liquid level (Kaner et al., 2018; Kreutzer et al., 2015; Ko et al., 2007; Pankajavalli et al., 2017; Wijanarko et al., 2019), alcohol (Bobin et al., 2018) liquid level (Yıldız and Coşkun, 2019) Ingestion Pace and Oral Motor Behavior facial expression (Kim and Bae, 2018a) intake sound (Kleinberger et al., 2023; Koizumi et al., 2011; Bi et al., 2016; Maramis et al., 2014; Kamachi et al., 2023; Chen and Yen, 2022) oral behavior (Kim et al., 2016b; Biyani et al., 2019; Zheng et al., 2025) wrist trajectory (Mendi et al., 2013; Lee et al., 2019; Scisco et al., 2011; Kim et al., 2016a; Kim and Bae, 2018b; Maramis et al., 2014; Turner-McGrievy et al., 2017), object trajectory (Kadomura et al., 2013b; Nakamura et al., 2023; Biyani et al., 2019; Hermans et al., 2017; Zhang et al., 2019), skin distance (Hori et al., 2023) wrist action (EMG) (Lee et al., 2019; Kim et al., 2016b; Sugita et al., 2018) food weight (de Vries et al., 2020; Krishna et al., 2025; Lo et al., 2007; Kim et al., 2016c; Ioakimidis et al., 2009; Mitchell et al., 2015; Hermans et al., 2017) Food Awareness or Emotional State intake sound (Kleinberger et al., 2023; Koizumi et al., 2011; Zhang and Parker, 2020), speech (Mayumi et al., 2023; Parra et al., 2023; Zhang and Parker, 2020) wrist trajectory (Wang et al., 2025), object trajectory (Khot et al., 2020) emotion state (EDA) (Carroll et al., 2013) food picture (AR) (Narumi et al., 2011) food weight (Mitchell et al., 2015) physical contact (Han and Kang, 2017) Target Factors and Behaviors Process Behaviors Safe Eating Practices hygiene (Wang and Zhang, 2021) temperature (Singhal et al., 2024)
+Appendix I DESIGN METRIC BETWEEN INTERVENTION MODALITIES AND TARGET FACTORS
+Table 8. Mapping of intervention Modalities to Target Factors and Behaviors in Ingestion Health
+Intervention Modality
+Visual
+Textual Graphical Symbolic Auditory Haptic Olfactory Physiological
+Food Choices (total 62 pubs) SMS, app/screen message, AR/MR message (Gutiérrez et al., 2018; Fuchs et al., 2019) (total 26 pubs) app/screen graph, AR/MR graph (Ganesh et al., 2014; Gutiérrez et al., 2018; Fuchs et al., 2019) LED (de Vries et al., 2020; Matsui et al., 2018; Ganesh et al., 2014; Kim et al., 2016c; Ritschel et al., 2018; Triantafyllidis et al., 2019) microphone speaker (Kadomura et al., 2013b; Cai et al., 2024; Kadomura et al., 2014; Robinson et al., 2021; Kadomura et al., 2013a; Zhao et al., 2021; De Carolis et al., 2024; Liu et al., 2024; Ritschel et al., 2018; Chukwu, 2011; Terzioğlu et al., 2023; Triantafyllidis et al., 2019; Agarwal et al., 2024; Luo et al., 2021; Robinson et al., 2020) piezoelectric vibration (Faltaous et al., 2021), Aerodynamic deformation (Pels et al., 2014), friction (Cai et al., 2024) Scented liquid (Cai et al., 2024)
+Intake Timing SMS (Bi et al., 2016; Varun et al., 2023), app/screen message (Liang, 2024; Faltaous et al., 2021; Ye et al., 2016) app/screen graph (Liang, 2024; Faltaous et al., 2021) pneumatic membrane (Xie et al., 2023), LED (Fortmann et al., 2014) microphone speaker (Xie et al., 2023), alarm (Varun et al., 2023) piezoelectric vibration (Faltaous et al., 2021), motor vibration (Varun et al., 2023; Ye et al., 2016) LED (Lessel et al., 2016; Ravindran et al., 2022)
+Environment microphone speaker (Faltaous et al., 2021)
+Context Factors
+Activity Around Ingestion SMS (Alexander et al., 2017)
+Intake Quantity SMS (Poddar et al., 2024; Kreutzer et al., 2015; Moses et al., 2023; Pankajavalli et al., 2017; Wijanarko et al., 2019), app/screen message (Faltaous et al., 2021; Chaudhry et al., 2016; Chiu et al., 2009; Alshurafa et al., 2015; Maramis et al., 2014; Varun et al., 2023; Tommy et al., 2017), LED message (Ravindran et al., 2022) app/screen graph (Faltaous et al., 2021; Chaudhry et al., 2016; Scisco et al., 2011; Maramis et al., 2014; Varun et al., 2023; Chiu et al., 2009; Ko et al., 2007; Wijanarko et al., 2019; Tommy et al., 2017), projector (Yıldız and Coşkun, 2019) pneumatic membrane (Xie et al., 2023), thermo -chromic (Kaner et al., 2018), LED (Bobin et al., 2018; Lessel et al., 2016; Ravindran et al., 2022), living-media (Zhou et al., 2021) microphone speaker (Xie et al., 2023; Farooq et al., 2017), alarm (Varun et al., 2023; Tommy et al., 2017), symbolic auditory (Zhou et al., 2021) piezoelectric vibration (Faltaous et al., 2021), motor vibration (Li and Bailenson, 2018; Chiu et al., 2009; Varun et al., 2023), pneumatic Deformation (Chen et al., 2022) essential oil (Li and Bailenson, 2018)
+Ingestion Pace and Oral Motor Behavior SMS (Lee et al., 2019; Mendi et al., 2013; Krishna et al., 2025; Moses et al., 2023), app/screen message (Bi et al., 2016; Kim and Bae, 2018b; Maramis et al., 2014; Kamachi et al., 2023; Turner-McGrievy et al., 2017; Hori et al., 2023; Lo et al., 2007; Ioakimidis et al., 2009) app/screen graph (Nakamura et al., 2023; Scisco et al., 2011; Kim and Bae, 2018b; Maramis et al., 2014; Turner-McGrievy et al., 2017; Hori et al., 2023; Lo et al., 2007), AR/MR graph (Kim and Bae, 2018a; Sugita et al., 2018) LED (de Vries et al., 2020; Kim et al., 2016b, a; Hermans et al., 2017; Kim et al., 2016c) chewing sound feedback (Kleinberger et al., 2023; Chen and Yen, 2022), microphone speaker (Ioakimidis et al., 2009; Sugita et al., 2018; Mendi et al., 2013; Koizumi et al., 2011; Kadomura et al., 2013b), physical symbolic feedback (Mitchell et al., 2015) piezoelectric vibration (Faltaous et al., 2021; Zhang et al., 2019), motor vibration (Kim et al., 2016b, a; Kim and Bae, 2018b; Kamachi et al., 2023; Hermans et al., 2017) Intraoral electrical stimulation (Biyani et al., 2019; Zheng et al., 2025)
+Food Awareness or Emotional State SMS(Carroll et al., 2013), app/screen message (Mayumi et al., 2023; Sun et al., 2020; Parra et al., 2023; Wang et al., 2025) app/screen graph(Sun et al., 2020; Parra et al., 2023; Zhang and Parker, 2020), AR/MR graph (Narumi et al., 2011) LED (Han and Kang, 2017; Khot et al., 2020) microphone speaker (Kadomura et al., 2013b; Koizumi et al., 2011; Han and Kang, 2017; Mayumi et al., 2023; Parra et al., 2023), chewing sound feedback (Kleinberger et al., 2023; Zhang and Parker, 2020), physical symbolic feedback (Mitchell et al., 2015) motor vibration (Khot et al., 2020) odor filter (Narumi et al., 2011)
+Target Factors and Behaviors
+Process Behaviors
+Safe Eating Practices motor vibration (Wang and Zhang, 2021; Singhal et al., 2024)